### v6

# CUBIC DIOPHANTINE EQUATION SATISFIABILITY IS $\Sigma^{0}_{1}$-COMPLETE

Abstract
Following a mechanization in Rocq, we prove that satisfiability of single Diophantine equations of total degree $\leq 3$ over $\mathbb{N}$ (with a computably bounded number of variables depending on the instance) is $\Sigma^{0}_{1}$-complete, and hence undecidable, in the sense of Hilbert’s Tenth Problem. The core construction is a uniform primitive recursive compiler mapping arithmetic sentence codes to finite structured cubic constraint systems, such that provability in $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ is equivalent to solvability of the emitted system. Conjunction is eliminated without squaring: all emitted constraints are merged into a single cubic Diophantine equation via a two-channel equality that preserves the cubic bound without sum-of-squares. All syntactic and inference checks compile to quadratic constraints; the only cubic terms come from linear selector variables multiplying those quadratic obligations. The mechanized core is constructive and requires no classical axioms: uniform computability, checker–constraint equivalence, and the degree-$3$ bound are verified by structural induction. From this base, undecidability is obtained by a classical Diagonal Argument: any total correct decider for cubic satisfiability would compose with the compiler to decide arithmetic theoremhood.
1 Introduction
1.1 Register Semantics
Intuition.
CircuitRegisterALUMemoryInputOutput I/O realization in Register Arithmetic ($\mathsf{RA}$) Mechanization yields mathematical programs. $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ ISA–logic duality witnesses effectivity of weak arithmetics.
Figure 1: Execution traces as finite sequences of configurations, understood arithmetically as coded finite objects, within a Von Neumann Architecture. Each step requires only Register Arithmetic $\mathsf{RA}=(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}$.
That certain forms of Diophantine satisfiability are $\Sigma_{1}$-complete is “old news.” One obtains this by composing MRDP [Matiyasevich, 1993] with reduction techniques such as Shielding [Jones, 1980]. However, these constructions typically achieve degree bounds through auxiliary padding and simulation gadgets; the resulting systems encode computation faithfully, but the structural complexity reflects the mechanics of simulation rather than any intrinsic proof-theoretic invariant.
On the other hand, mathematical practice and computer science have accumulated a vast repertoire of methods, ranging from ad hoc constructions to fully abstract algorithms. Yet whenever any such method is carried out, it is ultimately compiled into machine code (Figure 1) and executed as a finite sequence of operations over a fixed Instruction Set Architecture (ISA).
This observation is decisive for a problem as intrinsically tied to computability as Hilbert’s 10th. Every computationally enumerable language is recognized by some finite program over a fixed instruction set; hence any reduction establishing $\Sigma_{1}$-completeness must, implicitly or explicitly, encode such a finite execution trace. In this sense, Diophantine representations inherit a structural bottleneck common to all computably enumerable sets.
Accordingly, the appropriate framework should serve simultaneously as a language of machines and of recursion, making explicit the finite operational trace underlying existential search. At its core, it should require no more than bounded arithmetic over addition—computationally no stronger than an abacus.
Definition 1.1 ($\mathsf{RA}$).
We define Register Arithmetic as a specific fragment of first-order arithmetic. The language comprises of
$\mathcal{L}_{\mathsf{RA}}\;=\;\{0,\,S,\,+,\,=\},$ (1.1)
with constant 0, unary successor $S$, binary operation ++, and equality. The axioms of $\mathsf{RA}$ are:
  1. (i)
    Successor axioms reminiscent of Robinson Arithmetics,
    $S(x)\neq 0,$ (1.2)
    $S(x)=S(y)\to x=y,$ (1.3)
    $x\neq 0\to\exists y\;x=S(y),$ (1.4)
    $x+0=x,$ (1.5)
    $x+S(y)=S(x+y).$ (1.6)
  2. (ii)
    A bounded induction schema for $\Delta_{0}$-formulae $\varphi(x,\vec{y})$,
    $\Big(\varphi(0,\vec{y})\;\wedge\;\forall x\,\big(\varphi(x,\vec{y})\to\varphi(S(x),\vec{y})\big)\Big)\;\to\;\forall x\,\varphi(x,\vec{y}).$ (1.7)
  3. (iii)
    A collection schema $\mathrm{B}\Sigma_{1}$, $\varphi(x,y,\vec{z})$,
    $\forall u\;\big(\,\forall x\leq u\;\exists y\;\varphi(x,y,\vec{z})\;\to\;\exists v\;\forall x\leq u\;\exists y\leq v\;\varphi(x,y,\vec{z})\,\big).$ (1.8)
Thus $\mathsf{RA}$ is the restriction of $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ to the multiplication-free language $\mathcal{L}_{\mathsf{RA}}$:
$\mathsf{RA}\;=\;(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}.$ (1.9)
Multiplication ($\mathtt{Mult}$) must not be a primitive of $\mathsf{RA}$ but it must be representable as a $\Sigma_{1}$-relation via iterated addition (Theorem 1.2). Via this representation, $\mathsf{RA}$ interprets $\mathsf{Q}$ and is interpreted in $\mathsf{EFA}$:
$\mathsf{Q}\;\leq_{\mathrm{int}}\;\mathsf{RA}+\mathtt{Mult}\;\leq_{\mathrm{int}}\;\mathsf{EFA}.$ (1.10)
The set must be also recursive, in that membership of any sentence in the axiom set is decidable. Three meta-properties of $\mathsf{RA}$ are used in the sequel:
  1. (i)
    Representability. Each primitive recursive function or relation used in this development is numeralwise representable in $\mathsf{RA}$ (Proposition 1.10).
  2. (ii)
    Arithmetization. $\mathsf{RA}$ supports standard arithmetization of syntax: proof codes, formula codes, and the proof predicate $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ are all primitive recursive and hence representable.
  3. (iii)
    Derivability. The chosen proof predicate
    $\operatorname{Prov}_{\mathsf{RA}}(u)\equiv\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u)$ (1.11)
    must satisfy the Hilbert–Bernays–Löb conditions (see Theorem 3.13).
Theorem 1.2 (Representability of multiplication).
One of our first theorems is multiplication, defined as
$\mathtt{Mult}(x,y,z).$ (1.12)
This is the $\Sigma_{1}$-formula expressing that there exists a bounded register computation witnessing $z$ as the result of iterated addition of $x$, repeated $y$ times. Concretely, $\mathtt{Mult}(x,y,z)$ asserts the existence of a finite sequence
$t_{0},\;\ldots,\;t_{y}$ (1.13)
such that
$t_{0}=0,\quad t_{i+1}=t_{i}+x\;\text{ for }\;i<y,\quad t_{y}=z.$ (1.14)
This relation is primitive recursive and therefore representable in $\mathsf{RA}$ (Proposition 1.10). Verification of a proposed witness trace $(t_{0},\ldots,t_{y})$ requires only $\Delta_{0}(\mathcal{L}_{\mathsf{RA}})$ checks on sequence-encoded entries, so $\mathtt{Mult}$ fits the verification floor of $\mathsf{RA}$: the theory can verify any multiplication witness but does not have multiplication as a function symbol. We treat expressions of the form $x\cdot y=z$ in $\mathsf{RA}$-internal reasoning as abbreviations for $\mathtt{Mult}(x,y,z)$. In polynomial constraints over $\mathbb{N}$ (the Diophantine object language), the symbol $\cdot$ retains its standard meaning as ring multiplication.
Intuition.
The above are implicit in the foundational theory of bounded arithmetic [Hájek and Pudlák, 1993; Buss, 1997; Cook and Nguyen, 2010]. We make their role explicit because the theory is the calibration point of this work: the weakest arithmetic in which bounded verification of finite objects—proof traces, polynomial assignments, encoded sequences—is formalizable. All internal reasoning proceeds by inspecting given encoded witnesses; no uniform totality claims are available. The theory verifies correctness of given finite witnesses; it does not certify that witnesses exist for all instances.
This distinction is visible already at the polynomial level. For any fixed assignment $\vec{x}^{*}$ and $P$ of bounded degree, the equality $P(\vec{x}^{*})=0$ is a bounded check formalizable in $\mathsf{RA}$. The existential statement $\exists\,\vec{x}\;P(\vec{x})=0$ is $\Sigma_{1}$ and lies strictly beyond such instance verification.
The cubic threshold arises naturally from this gap. All structural and syntactic verification—axiom matching, formula decoding, target equality—compiles to quadratic constraints (degree $\leq 2$), while escalation arises from enforcing bounded existential witnesses. One standard schematic guard uses
$u=1+v^{2},\quad\exists w,(\mathtt{Mult}(u,E,w)\wedge w=0),$ (1.15)
where $\mathtt{Mult}$ is the $\Sigma_{1}$-representation of multiplication by iterated addition (Theorem 1.2). This guard does not invoke multiplication as a primitive; rather, it demands a witness trace of iterated addition that $\mathtt{RA}$ can verify via $\Delta_{0}$ checks on the encoded sequence. The guard enforces that either $E=0$ or the square term forces strict positivity. Eliminating the bounded existential therefore introduces exactly one multiplicative interaction and raises the degree from $2$ to $3$. In the implemented constraint system, the actual cubic terms come from degree-$1$ selector variables multiplying quadratic obligations; the guard template is kept only as a conceptual schematic and is not multiplied by any quadratic constraint.
Lemma 1.3 (Guard correctness and degree).
Let $E(\vec{x})$ be a polynomial of degree $\leq 1$ and define $u=1+v^{2}$. Then
$u\cdot E(\vec{x})=0\;\Longleftrightarrow\;E(\vec{x})=0,$
because $u\geq 1$ for all $v$. The guarded equation has degree at most $3$. In this development, quadratic constraints are not multiplied by quadratic guards; selectors (degree $1$) are used for activation of quadratic constraints.
We recall the degree hierarchy of Diophantine equations over $\mathbb{N}$:
  1. (i)
    Integer roots of a single polynomial can be effectively determined.
  2. (ii)
    Degree $2$: Certain structured quadratic forms are decidable via classical number-theoretic methods (Hasse–Minkowski [Siegel, 1972; Grunewald and Segal, 1981, 2004] for quadratic forms over $\mathbb{Q}$, Lagrange [Lagrange, 1770] for sums of squares). Crucially, degree-$2$ constraints lack sufficient expressive power to encode arbitrary proof checking—verifying a multiplication witness against a verification constraint requires degree $\geq 3$.
  3. (iii)
    Degree $4$: Jones [1980, 1982] showed that every recursively enumerable set is the solution set of a degree-44 Diophantine equation with 5858 variables. Undecidability follows classically from MRDP [Robinson et al., 1961; Matiyasevich, 1970; Davis, 1973; Matiyasevich, 1993].
  4. (iv)
    Degree $3$: Black-box reductions from MRDP composed with Jones Shielding [Jones, 1980] and a quartic-to-cubic-system closure yield $\Sigma_{1}$-completeness of cubic systems (with unbounded variables), but furnish no degree-local construction and no single-equation result at degree $3$.
  5. (v)
    The existence of a single universal cubic Diophantine equation (in a fixed number of variables) remains open; the present work establishes $\Sigma^{0}_{1}$-completeness of single cubic equations with computably bounded variable count via a direct proof-to-polynomial compiler, and attests to their constructibility.
Gasarch [2021] (cf. Definition 1.4 for notation) summarizes the state of the problem:
“Work on showing $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ or $\mathrm{H}{\mathbb{Z}}(d,n)={\rm U}$ seems to have stalled. Perhaps the problems left are too hard. Perhaps the problems left could be resolved but it would be very messy. Perhaps computer-work could help […]” [Gasarch, 2021]
The existence of a universal Diophantine equation of degree $3$ in a fixed number of variables sits in the open region between the fully understood quadratic case and the high-degree expressive universality constructions.
Jones explicitly singled out Degree $3$ (over $\mathbb{N}$):
“We know that there exist universal Diophantine equations of degree $4$ […] but none of degree $2$. […] Thus the only open problem as regards the degree is $\delta=3$.” [Jones, 1980]
1.2 Problem Semantics
Definition 1.4 ($\mathrm{H}{\mathbb{N}}(d,n)$ notation).
We fix Gasarch’s convenient notation (from the excerpt) for parameterized Diophantine solvability throughout. For $d,n\in\mathbb{N}$:
  1. (i)
    $\mathrm{H}{\mathbb{N}}(d,n)$ is the decision problem: “Given
    $P\in\mathbb{Z}[x_{1},\ldots,x_{k}]\;\text{with}\;k\leq n\;\text{and}\;\deg P\leq d,$ (1.16)
    does $P(\vec{x})=0$ have a solution in $\mathbb{N}^{k}$?” While $\mathrm{H}{\mathbb{Z}}(d,n)$ is the analogous problem with solutions in $\mathbb{Z}^{k}$.
  2. (ii)
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ (resp. $\rm U$}) means the problem is decidable (resp. undecidable). Similarly for $\rm H$$\mathbb{Z}$.
Lemma 1.5 (Transfer).
Since any $\mathbb{N}$-decider can be used to decide $\mathbb{Z}$ by sign enumeration,
$m\in\mathbb{N}\;\Longleftrightarrow\;\exists\,x_{1},x_{2},x_{3}\in\mathbb{Z}:\;m=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3}$ (1.17)
so that it embeds $\mathbb{N}$-witnesses into $\mathbb{Z}$-witnesses. Replacing each $x_{i}$ in $P$ by
$x_{i1}^{2}+x_{i2}^{2}+x_{i3}^{2}+x_{i3}$ (1.18)
yields a polynomial over $\mathbb{Z}$ of degree $2d$ in $3n$ variables, so ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ implies $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$. Fix $d,n\in\mathbb{N}$ and write $\ast\in\{\mathbb{N},\mathbb{Z}\}$. The following similar structural implications hold.
  1. (i)
    By monotonicity in parameters, if $d^{\prime}\leq d$ and $n^{\prime}\leq n$,
    $\mathrm{H}{\ast}(d,n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm D},$ (1.19)
    $\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d,n)={\rm U}.$ (1.20)
  2. (ii)
    By $\mathbb{N}\Longrightarrow\mathbb{Z}$ sign expansion,
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(d,n)={\rm D},$ (1.21)
    ${\rm H}\mathbb{Z}(d,n)={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm U}.$ (1.22)
  3. (iii)
    $\mathbb{Z}\Rightarrow\mathbb{N}$ via three–square encoding,
    ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm D},$ (1.23)
    $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(2d,3n)={\rm U}.$ (1.24)
It is easy to see that undecidability propagates upward in $(d,n)$, and decidability downward, while the only domain transfers requiring parameter growth are those encoding $\mathbb{Z}$ in $\mathbb{N}$ or $\mathbb{N}$ in $\mathbb{Z}$.
By logic. Monotonicity follows since enlarging $(d,n)$ enlarges the input class, while restricting $(d,n)$ shrinks it. Domain transfer without blow-up uses sign enumeration. Difference encoding represents $x\in\mathbb{Z}$ as uvu-v with $u,v\in\mathbb{N}$, preserving degree and doubling variables. Three–square encoding replaces $x\in\mathbb{N}$ by
$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3},$ (1.25)
doubling degree and tripling variables. The undecidability statements are contrapositives. Downward transfer fails because no general compression of degree or variables preserves solvability. ∎
Lemma 1.6 (Systems and general aggregation).
For $d,n\in\mathbb{N}$ and $\ast\in\{\mathbb{N},\mathbb{Z}\}$, let $\mathrm{H}{\ast,S}(d,n)$ denote the decision problem: “Given a finite system
$P_{1}(\vec{x})=0,\ldots,P_{m}(\vec{x})=0$ (1.26)
with $k\leq n$ variables and $\deg P_{i}\leq d$ for all $i$, determine whether there exists $\vec{x}\in\ast^{k}$ satisfying all equations.” All implications of Lemma 1.5 hold verbatim with $\mathrm{H}{\ast,S}(d,n)$ in place of $\mathrm{H}{\ast}(d,n)$. Over $\mathbb{N}$ and $\mathbb{Z}$, sum-of-squares aggregation
$P_{1}=\cdots=P_{m}=0\;\Longleftrightarrow\;\sum_{i=1}^{m}P_{i}(\vec{x})^{2}=0$ (1.27)
reduces systems to one equation trivially, but doubles the degree bound ($d\mapsto 2d$). This general collapse is therefore not suitable for preserving a cubic degree bound.
Proposition 1.7 (Aggregation for emitted systems).
Our first contribution will be to show that valid systems emitted by the compiler of Theorem 4.31 admit an effective map producing a single equation of degree $\leq 3$ with the same solution set (up to harmless slack variables for band capacities). The construction rewrites each constraint as an equality of two nonnegative polynomials and merges all equalities into a single cubic Diophantine equation using Fibonacci-banded Zeckendorf isolation (Section 4.8); no sum-of-squares blow-up is used.
1.3 Ansatz
Ansatz 1.8 (Target).
Our target is to prove that the predicate
$\mathsf{CubicEqSat}(e):\equiv\exists\vec{x}\;P_{e}(\vec{x})=0,$
where $e$ codes a single polynomial of total degree $\leq 3$ with its variable count as part of the code, is $\Sigma^{0}_{1}$-complete. Equivalently, satisfiability of single cubic equations over $\mathbb{N}$ with instance-dependent arity is undecidable. The argument proceeds in three conceptual layers:
  1. Layer 1.
    Compilation. We construct a uniform primitive recursive compiler
    $\varphi\longmapsto\mathcal{S}_{\varphi}$ (1.28)
    mapping $\mathsf{RA}$-sentence codes to finite systems of Diophantine equations over $\mathbb{N}$, each of degree $\leq 3$, such that
    $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\;\Longleftrightarrow\;\mathsf{CubicSat}(\ulcorner\mathcal{S}_{\varphi}\urcorner),$ (1.29)
    where $\mathsf{CubicSat}$ denotes satisfiability of a single degree-$\leq 3$ equation coinciding with “theoremhood” inside a Hilbert–Ackermann System.
  2. Layer 2.
    Constraining. The degree budget is controlled locally: finite proof traces are arithmetized via Carryless Pairing in the Fibonacci basis (Section 3), syntactic and inference checks compile to quadratic constraints (Section 4), and the only cubic terms arise when a linear selector activates a quadratic obligation. These remain separate through the local degree certification; only at the final step are they collapsed to the single cubic equation by two-channel aggregation (Section 4.8).
  3. Layer 3.
    Metatheory. By the classic Diagonal argument, any total correct decider for $\mathsf{CubicSat}$ would compose with the compiler to decide $\mathsf{Thm}$, (Section 5) leading to contradiction. If $\mathsf{RA}$ could internally certify correctness of such a decider on the compiler image, the provability bridge would force $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, contradicting Gödel’s Second Incompleteness Theorem (Gödel’s 2nd henceforth) [Gödel, 1931; Kleene, 1952; Feferman, 1960; Boolos et al., 2012] (Section 5.3).
Ansatz 1.9 (Metalogic).
The development harmonizes two formal methodologies: $\mathsf{RA}$ as our theory, while proof ereasure inside Calculus of Inductive Constructions of Coquand and Huet [1988] acts as a second metalogical verification tool. Layer $1$ and $2$ of the argument are fully constructive: all maps are primitive recursive, equivalences are witnessed by explicit transformations, and the entire pipeline is mechanized in Rocq without axioms beyond weak aritmetics. The construction remains constructive as far as it can: classical reasoning enters only at Layer $3$, where the Diagonal Lemma produces a fixed-point sentence whose existence relies on the final discharge via Gödel’s 2nd. This boundary is not a limitation but a structural feature: the undecidability of cubic satisfiability is inherently in need of a classical impossibility result, and the point at which constructivity is exhausted marks precisely the logical depth of the claim. Interestingly, the pipeline does not appeal to MRDP [Matiyasevich, 1993]: the Diophantine representation of proof evaluation is constructed directly, with explicit degree control at every step, rather than inherited from the historic theorem.
Proposition 1.10 (Numeralwise representability).
For every primitive recursive function $f:\mathbb{N}^{k}\to\mathbb{N}$ and relation $R\subseteq\mathbb{N}^{k}$ there is an $\mathcal{L}_{\mathsf{RA}}$-formula $\varphi_{f}(\vec{x},y)$ such that for all numerals $\overline{\vec{n}},\overline{m}$:
$f(\vec{n})=m$ $\;\Longrightarrow\;\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m}),$ (1.30)
$\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m})\wedge\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m^{\prime}})$ $\;\Longrightarrow\;m=m^{\prime}.$ (1.31)
A similar clause holds for relations $R$ with a witnessing formula $\varphi_{R}(\vec{x})$. This numeralwise representability suffices to arithmetize the compiler and checker.
Proposition 1.11 (Instancewise checker–constraint equivalence).
The map $\varphi\mapsto\mathcal{S}_{\varphi}$ is primitive recursive. Moreover, for each specific pair of numerals $(\bar{p},\bar{t})$,
$\mathrm{check}(\bar{p},\bar{t})\;\Rightarrow\;\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x}),$ (1.32)
and conversely,
$\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x})\;\Rightarrow\;\mathrm{check}(\bar{p},\bar{t})\text{ holds in }\mathbb{N}.$ (1.33)
In particular, for each standard proof code $p$ of $\varphi$, $\mathsf{RA}$ proves the existence of a satisfying assignment for the emitted constraint system.
By construction. Each initial function (zero, successor, projection) is definable in $\mathcal{L}_{\mathsf{RA}}$. Closure under composition and primitive recursion follows by induction on the defining scheme, using $\mathrm{I}\Delta_{0}$ for the recursive step and $\mathrm{B}\Sigma_{1}$ to bound intermediate values. We use these facts only for numeralwise representability: for each concrete tuple of numerals $\overline{\vec{n}}$, $\mathsf{RA}$ proves the corresponding instance of $\varphi_{f}$. We do not assert that $\mathsf{RA}$ proves $\forall\vec{x}\,\exists!y\,\varphi_{f}(\vec{x},y)$ for arbitrary primitive recursive $f$. ∎
2 Constructive Environment
2.1 Mechanization
Ansatz 2.1.
This section establishes the computational infrastructure, coding of syntax, the proof predicate with derivability conditions, and some prerequisites of our fixed point. The pipeline described in the Ansatz 1.9 is implemented in Gallina, the specification language of
By Curry–Howard [Sørensen and Urzyczyn, 2006], the implicational fragment with $\bot$ admits a “simply-typed” $\lambda$-definable interpretation: each axiom (K, S, EFQ) realizes as a closed term and Modus ponens corresponds to application. The development formalizes the Hilbert-style proof evaluator in $\mathtt{Kernel\_API}$, $\mathtt{R13}$ (Section 4.1), compiles proof traces to cubic Diophantine constraints (Section 4), and extracts a certified boolean kernel,
$\mathtt{cubic\_accepts}:\mathbb{N}\longrightarrow\mathbb{B},$ (2.34)
whose acceptance set coincides with the acceptance predicate of the proof evaluator. The computational content extracts to a standalone OCaml artifact.
Repository: github.com/…/Proofcase/…/T002/
2.2 Repository Structure and Trust Base
Ansatz 2.2 (Compilation layers).
The reduction $\varphi\mapsto\mathcal{S}_{\varphi}$ is organized in four formally separated layers, each corresponding to a segment of the $\mathtt{R\_SigmaReduction}$ module chain:
  1. Module 1.
    Foundations ($\mathtt{R00}$$\mathtt{R03}$). The expression-degree framework ($\mathtt{R00}$), Fibonacci/Zeckendorf arithmetic for Carryless Pairing ($\mathtt{R01}$$\mathtt{R02}$), and carryless sequence indexing via bounded unpair recursion ($\mathtt{R03}$). All encoding primitives are quadratic.
  2. Module 2.
    Verification and constraint compilation ($\mathtt{R04}$$\mathtt{R11}$). Hilbert formula syntax and the executable proof checker ($\mathtt{R04}$$\mathtt{R05}$), formula and sequence coding via Carryless Pairing ($\mathtt{R06}$), and constraint assembly ($\mathtt{R08}$$\mathtt{R11}$) mapping proof-checking conditions to polynomial equations. Axiom and target constraints have degree $\leq 2$; cubic terms arise only when degree-1 selector variables activate quadratic obligations, keeping every emitted constraint at degree $\leq 3$. The assembly theorem $\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ ($\mathtt{R11}$) certifies the uniform bound $\delta\leq 3$.
  3. Module 3.
    Aggregation and kernel interface ($\mathtt{R12}$$\mathtt{R13}$). Two-channel equality collapses the constraint system to a single cubic equation ($\mathtt{R12}$). The collapse uses a no-carry lemma from Zeckendorf uniqueness plus explicit capacity constraints; the implementation aggregates as $\mathtt{main::capacities}$, and a second banding pass yields a single equation. The public kernel API exports the certified boolean function $\mathtt{cubic\_accepts}$ with correctness and degree-bound guarantees ($\mathtt{R13}$).
  4. Module 4.
    Reduction endpoint and proof theory ($\mathtt{R14}$$\mathtt{R19}$ and $\mathtt{P00}$$\mathtt{P06}$). The reduction core and code bridge ($\mathtt{R14}$$\mathtt{R15}$), the reduction nucleus $\mathtt{compiler\_primitive}$ with coded evaluator semantics ($\mathtt{R16}$), OCaml extraction ($\mathtt{R17}$), and stable API boundaries ($\mathtt{R18}$$\mathtt{R19}$). The proof-theory layer ($\mathtt{P00}$$\mathtt{P06}$) mechanizes concrete provability ($\mathtt{P00}$), HBL derivability conditions ($\mathtt{P01}$), the fixed-point construction ($\mathtt{P02}$), the premise-free coded undecidability theorem ($\mathtt{P03}$), the RA-certification corollary ($\mathtt{P04}$), and the toggle-witness contradiction core ($\mathtt{P05}$$\mathtt{P06}$).
Ansatz 2.3.
Rocq extraction ($\mathtt{R17}$) emits two self-contained OCaml modules, each free of standard-library dependencies:
  1. (i)
    cubic_checker.ml exports:
    $\mathtt{cubic\_accepts}:\mathsf{Proof}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\},$ (2.35)
    the proof-checking kernel. Internally the module re-derives its own bool, nat, list, and form types and contains the full verification pipeline: structural formula equality (form_eqb), axiom-schema recognizers (is_K, is_S, is_EFQ), bounded Modus ponens witness search (mp_witness), and line-by-line validation (check_lines). The entry point $\mathtt{cubic\_accepts}$ is definitionally equal to $\mathtt{check}$; the kernel API ($\mathtt{R13}$) re-exports the checker under a reduction-facing alias.
  2. (ii)
    cubic_compile.ml exports a (legacy) polynomial,
    $\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial})$ (2.36)
    the single equation,
    $\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{Polynomial}\times\mathsf{Polynomial},$ (2.37)
    together with the symbolic expression type,
    $\mathtt{Expr}(\mathtt{Const}\mid\mathtt{Var}\mid\mathtt{Add}\mid\mathtt{Sub}\mid\mathtt{Mult})$ (2.38)
    and the structural-recursion function,
    $\mathtt{degree}:\mathtt{Expr}\to\mathbb{N}$ (2.39)
    Given a proof and target, $\mathtt{emit\_cubic\_system}$ produces polynomial constraints, while the the endpoint $\mathtt{emit\_single\_cubic}$ applies two-channel aggregation and returns a single cubic equality surface $(P,Q)$ representing $P=Q$. Mapping $\mathtt{degree}\circ\mathtt{poly\_expr}$ over the emitted expressions empirically verifies the degree-$3$ bound of Proposition 4.28. The module additionally includes the constraint types ($\mathtt{C\_Axiom}$, $\mathtt{C\_MP}$, $\mathtt{C\_Justification}$, $\mathtt{C\_Target}$), the canonical assignment builder, and the Carryless Pairing and carryless sequence-indexing infrastructure required by the compilation pipeline.
The correctness of the mechanized development depends on at least three components external to the formalization: (i) the Rocq kernel (the core type checker); (ii) the extraction mechanism; (iii) OCaml compilation. No axioms, parameters, or $\mathtt{Admitted}$ obligations are introduced: $\mathtt{PrintAssumptions}$ on the endpoint theorems of both $\mathtt{SigmaReduction}$ and $\mathtt{ProofTheory}$ yields an empty context.
Intuition.
A clarification on the diagonal construction is necessary, since it appears at two distinct levels. The Rocq development contains a code-level diagonal ($\mathtt{P02}$$\mathtt{P03}$): it operates entirely on natural-number codes, producing a fixed-point code $g$ satisfying $g=\mathtt{apply\_template}\;T\;g$ by arithmetic on encodings, and derives $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$)—the impossibility of a coded decider—constructively, without classical axioms. This is a theorem about codes: it quantifies over natural numbers representing programs, not over Turing Machines as external objects. The classical meta-theoretic diagonal of Section 5.2 (Theorem 5.2) is a separate argument: it applies the Diagonal Lemma to arbitrary Turing Machines with case analysis via the Law of the Excluded Middle, and is not Rocq-verified. Likewise, Gödel’s 2nd as applied to $\mathsf{RA}$ (Corollary 5.7) relies on the arithmetization of Section 3.4 and standard classical logic, and is presented as an ordinary mathematical argument. A machine-readable cross-reference map of manuscript claims to Rocq theorems is provided in $\mathtt{TechnicalSummary.yaml}$.
Remark.
The reduction used for $\Sigma^{0}_{1}$-hardness does not fix a proof-length bound. The compiler $f(u)$ treats the entire proof witness as a single carryless list code $p$ (together with its per-line selectors), and emits one cubic equation whose free variables include $p$, the target code $u$, and all selector/auxiliary variables. The Diophantine instance is therefore
$\exists\,p,\,\vec{x}\;\Phi(u,p,\vec{x})=0$
with $\deg\Phi\leq 3$ and a fixed arity determined solely by the structure of the checker/aggregator—not by a chosen proof length. Internally, universally quantified line ranges $i<|p|$ are compiled into selector-indexed constraints over the encoded list, so the existential quantifier over $p$ subsumes all proof lengths. This yields, for each $u$,
$\operatorname{Prov}_{\mathsf{RA}}(u)\;\Longleftrightarrow\;\exists\,p,\vec{x}\;\Phi(u,p,\vec{x})=0\;\Longleftrightarrow\;\mathsf{CubicEqSat}(f(u)),$
where $f$ is primitive recursive and $\mathsf{CubicEqSat}$ ranges over a single cubic equation with fixed variable budget.
The banded aggregation uses dynamically sized Fibonacci bands: the width parameter $K$ is computed from the size of the emitted constraint system (number of constraints / channels), not from a fixed global constant. Each channel is paired with a capacity constraint of the form $B_{K}-1-A_{i}=\sum_{j=1}^{4}s_{ij}^{2}$, where $B_{K}=F_{K+1}$ and $s_{ij}$ are slack variables. Thus any assignment with $A_{i}\geq B_{K}$ is rejected. Since every $A_{i},B_{i}$ is a polynomial in the proof/selector variables of degree $\leq 2$, the capacity bound is computable from the instance size, and the added slack variables preserve equivalence: a satisfying assignment to the original system extends to one satisfying the capacity equations, and any violation of the bound is ruled out by the slack equality.
3 Arithmetic Infrastructure
3.1 Fibonacci Preliminaries
Intuition.
The encoding of finite proof traces as Diophantine witnesses requires an injective, primitive recursive pairing function over $\mathbb{N}$. Classical binary pairing functions induce structural interactions through carry propagation, which complicates Diophantine reflection. We instead construct a pairing mechanism using Zeckendorf representations in the Fibonacci basis. This return to Fibonacci arithmetic is not incidental; it follows a historical line of development in work on Hilbert’s Tenth Problem by Chudnovsky, Jones, Matiyasevich, and others, see Jones [1975]; Matiyasevich [1993].
The resulting “carryless” coding discipline and band separation refine this tradition. The additive Fibonacci recurrence yields a canonical representation with explicit inversion, while avoiding interference between adjacent components. This separation permits controlled aggregation of constraints, and the Diophantine degree cost remains quadratic throughout the encoding layer. See Rosko [2026] for a self-contained treatment of the construction and its formal properties.
Definition 3.1 (Fibonacci canonicalization).
The Fibonacci sequence is defined here recursively by
$F_{0}=0,\quad F_{1}=1,\quad F_{n+2}=F_{n+1}+F_{n}\quad\text{for all }n\geq 0.$ (3.40)
Lemma 3.2 (Fibonacci monotonicity).
For all $n\geq 1$,
$F_{n+1}\geq F_{n},\quad F_{n+2}>F_{n}.$ (3.41)
Induction coincides with the recursive definition. ∎
Theorem 3.3 (Zeckendorf’s Theorem).
Every natural number $n>0$ admits a unique representation
$n=\sum_{k\in S}F_{k},\quad S\subseteq\{2,3,4,\ldots\}$ (3.42)
where (i) $S$ is a finite set of indices and (ii) no two consecutive integers belong to $S$. We call $S$ the Zeckendorf support of $n$ and write $Z(n):=S$.
By Zeckendorf [1972]. The non-adjacency constraint is preserved by the monotonic gap property $F_{k+2}>F_{k}$. Assume two distinct representations for $n$. Let $k_{\max}$ be the largest index where they differ. Without loss of generality, suppose the first representation includes $F_{k_{\max}}$ and the second does not. The Fibonacci identity
$\sum_{i=2}^{k_{\max}-2}F_{i}=F_{k_{\max}}-1$ (3.43)
implies that any combination of non-consecutive indices below $k_{\max}$ sums to strictly less than $F_{k_{\max}}$, yielding a contradiction. ∎
3.2 Carryless Pairing
Intuition.
The Zeckendorf Representation assigns to each $n$ a canonical support set $Z(n)\subseteq\mathbb{N}$. We construct an injective pairing by embedding the supports of $x$ and $y$ into disjoint frequency bands within a combined Zeckendorf expansion. The even-indexed Fibonacci numbers encode $x$; the odd-indexed numbers, shifted beyond the range of $x$’s encoding, encode $y$. No arithmetic carries occur and composition is achieved by set-theoretic union of supports. For a presentation with consult either Rosko [2026] or the repository:
Repository: github.com/…/Proofcase/…/T001/
Theorem 3.4 (Carryless Pairing).
Define $Z(0):=\varnothing$. For $x,y\in\mathbb{N}$, set
$E(x)$ :={ $:=\{\,2k:k\in Z(x)\,\},$ (3.44)
$r(x)$ $:=\min\{\,m\geq 2:F_{m}>x\,\},$ (3.45)
$B(x)$ $:=2r(x),$ (3.46)
$O(x,y)$ $:=\{\,B(x)+(2j-1):j\in Z(y)\,\}.$ (3.47)
Define the pairing
$\mathrm{pair}(x,y)\;:=\;F_{2}\;+\;\sum_{i\in E(x)\cup O(x,y)}F_{i}.$ (3.48)
Then $\mathrm{pair}(x,y)\geq 1$ for all $x,y$, the sets $E(x)$ and $O(x,y)$ are disjoint with
$\max E(x)<B(x)\leq\min O(x,y)-3,$ (3.49)
and $\mathrm{pair}$ is injective. Moreover there are primitive recursive projections $\mathrm{hd},\mathrm{tl}$ such that
$u>0\;\to\;\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u,\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0,$ (3.50)
and $\mathrm{unpair}(u)=(\mathrm{hd}(u),\mathrm{tl}(u))$ is the inverse of $\mathrm{pair}$ on $\mathbb{N}^{2}$.
By induction over uniquness, construction of $B(x)=2r(x)$, and rank definition. Band separation ensures the union $E(x)\cup O(x,y)$ is a valid Zeckendorf support. By uniqueness (Theorem 3.3), the support is recoverable; the even and odd bands are then separated and decoded independently. ∎
Definition 3.5 (List encoding).
Define recursively:
$\mathrm{code}_{\mathrm{list}}([])$ $:=0,$ (3.51)
$\mathrm{code}_{\mathrm{list}}(x_{1},\ldots,x_{n})$ $:=\mathrm{pair}(x_{1},\mathrm{code}_{\mathrm{list}}(x_{2},\ldots,x_{n})).$ (3.52)
The function $\mathrm{code}_{\mathrm{list}}$ inherits injectivity from Theorem 3.4.
Remark.
The construction is termed “carryless” because composite codes are formed by set union of disjoint supports rather than additive mixing: decoding is local and requires no interdependent constraints. All operations are primitive recursive (bounded shifts and Fibonacci summation), and the arithmetic complexity remains quadratic.
3.3 Carryless Sequence Indexing
Intuition.
Finite proof traces are encoded directly by nested Carryless Pairing. The same pairing infrastructure used for formula coding also provides line access: the $i$-th entry is obtained by iterating the tail projection $i$ times and reading the head. This removes any dependence on modular sequence decoding while keeping extraction primitive recursive and degree-local.
Definition 3.6 (Head and tail projections).
Let $\mathrm{pair}:\mathbb{N}^{2}\to\mathbb{N}$ be the carryless pairing of Theorem 3.4. Define primitive recursive projections
$\mathrm{hd},\mathrm{tl}:\mathbb{N}\to\mathbb{N}$ (3.53)
by
$\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u\;\;(u>0),\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0.$ (3.54)
Definition 3.7 (Indexed extractor).
Define by primitive recursion:
$\mathrm{tail}^{0}(s)$ $:=s,$ (3.55)
$\mathrm{tail}^{i+1}(s)$ $:=\mathrm{tl}(\mathrm{tail}^{i}(s)),$ (3.56)
$\mathrm{line}(s,i)$ $:=\mathrm{hd}(\mathrm{tail}^{i}(s)).$ (3.57)
Theorem 3.8 (Carryless sequence representation).
For every finite sequence $a=(a_{0},\ldots,a_{n-1})$, there exists $s\in\mathbb{N}$ such that
$\mathrm{line}(s,i)=a_{i}\quad\text{for all }i<n.$ (3.58)
Take $s:=\mathrm{code}_{\mathrm{list}}(a_{0},\ldots,a_{n-1})$ from Definition 3.5. The claim follows by induction on $i$, using invertibility of $\mathrm{pair}$ (Theorem 3.4). ∎
Definition 3.9 (Bounded extraction witnesses).
For fixed index $i$, introduce witness variables
$t_{i,0},\ldots,t_{i,i+1},\quad h_{i,0},\ldots,h_{i,i}.$ (3.59)
Define the chain by
$t_{i,0}$ $=s,$ (3.60)
$t_{i,r}$ $=\mathrm{pair}(h_{i,r},t_{i,r+1})\quad(0\leq r\leq i).$ (3.61)
Then $h_{i,i}=\mathrm{line}(s,i)$ and $t_{i,i+1}=\mathrm{tail}^{\,i+1}(s)$.
Intuition.
In the sequel, proof sequences are represented by a single carryless list code $s$. Formula coding and line extraction now use the same pairing/unpairing primitives. The extraction layer contributes only quadratic constraints before selector activation of Section 4.
3.4 Coding of Syntax
Intuition.
[…]
Definition 3.10 (Syntactic and proof predicates).
Using Gödel Numbering, we define the following predicates on $\mathbb{N}$:
[…]
$\operatorname{Sent}(u)$ $\;:\equiv\;\operatorname{Form}(u)\;\wedge\;u\text{ codes a sentence},$ (3.63)
$\operatorname{Proof}_{\mathsf{RA}}(p,u)$ $\;:\equiv\;p\text{ codes a valid $\mathsf{RA}$-proof coded by }u,$ (3.64)
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;:\equiv\;\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u).$ (3.65)
Observe how $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ holds when $p$ codes a finite sequence such that $\varphi_{n-1}$ is the sentence coded by $u$, and for each $i<n$, either $\varphi_{i}$ is an instance of an axiom schema of $\mathsf{RA}$, or there exist $j,k<i$ with $\varphi_{k}\equiv\varphi_{j}\to\varphi_{i}$ by Modus ponens.
Lemma 3.11 (Primitive recursiveness).
The predicates $\operatorname{Form}$, $\operatorname{Sent}$, and $\operatorname{Proof}_{\mathsf{RA}}$ are primitive recursive, consequently, $\operatorname{Prov}_{\mathsf{RA}}$ is $\Sigma_{1}$.
By definition. Each predicate is decided by bounded inspection of the input code. $\operatorname{Form}$ and $\operatorname{Sent}$ are checked by structural induction on code depth. $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ is checked by scanning the sequence coded by $p$: at each line, either an axiom-schema match is verified by pattern matching, or a Modus ponens witness is found by search bounded by the line index. All checks are bounded computations, hence primitive recursive. The $\Sigma_{1}$-classification of $\operatorname{Prov}_{\mathsf{RA}}$ follows from the existential quantification over $p$. ∎
Definition 3.12 (Bridge translation).
Let $\operatorname{tr}:\mathcal{L}_{\mathsf{RA}}\to\mathcal{L}_{\bot,\to}$ be the primitive recursive encoding used by the kernel checker $\mathrm{check}_{\to}(p,u)$ on the implicational language. The bridge is captured by two numeralwise implications:
$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;\Longrightarrow\;\mathsf{RA}\vdash\exists p\,\mathrm{check}_{\to}(p,\operatorname{tr}(\bar{u})),$ (3.66)
[…]
That is, provable sentences have kernel proofs (formalizable in $\mathsf{RA}$), and any kernel proof witnesses truth of the corresponding $\mathsf{RA}$ sentence in the standard model. No global biconditional inside $\mathsf{RA}$ is assumed.
3.5 Fixed-Point Construction
Theorem 3.13 (Derivability Conditions).
The provability predicate $\operatorname{Prov}_{\mathsf{RA}}$ satisfies the Hilbert–Bernays–Löb derivability conditions. For all sentences $\varphi,\psi$:
$\textsf{D1}:\;$ $\text{If}\;\mathsf{RA}\vdash\varphi,\;\text{then}\;\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner).$ (3.68)
$\textsf{D2}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\urcorner\bigr).$ (3.69)
$\textsf{D3}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\to\psi\urcorner)\to\bigl(\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\psi\urcorner)\bigr).$ (3.70)
These conditions are required for Gödel’s 2nd as applied in Corollary 5.7.
$\mathsf{D1}$ is immediate: a proof of $\varphi$ in $\mathsf{RA}$ is a concrete finite sequence whose code witnesses $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)$. (D2) and (D3) are standard consequences of the primitive recursiveness of $\operatorname{Proof}_{\mathsf{RA}}$ (Lemma 3.11) and the representability of primitive recursive functions in $\mathsf{RA}$ (Proposition 1.10): $\mathsf{RA}$ can internally simulate concatenation of proof sequences and verify that the result is again a valid proof. See Boolos et al. [2012]. ∎
Theorem 3.14 (Diagonal Lemma).
For every $\mathcal{L}_{\mathsf{RA}}$-formula $\varphi(x)$ with exactly one free variable $x$, there exists a sentence $u$ such that
$\mathsf{RA}\;\vdash\;u\;\leftrightarrow\;\varphi(\ulcorner u\urcorner).$ (3.71)
Define the substitution function $\mathrm{sub}(e,n)$ to be the code of the sentence obtained by substituting the numeral $\overline{n}=S^{n}(0)$ for the free variable in the formula coded by $e$. The function $\mathrm{sub}$ is primitive recursive, it acts by bounded structural replacement in the formula tree, and is therefore representable in $\mathsf{RA}$ by Proposition 1.10. Let
$\psi(x):\equiv\varphi(\mathrm{sub}(x,x))\quad\text{with}\quad e:=\ulcorner\psi\urcorner$ (3.72)
Define $u$ to be the sentence with code $\mathrm{sub}(e,e)$. Then
$u\;\equiv\;\psi(\overline{e})\;\equiv\;\varphi\bigl(\mathrm{sub}(\overline{e},\overline{e})\bigr)\;\equiv\;\varphi(\ulcorner u\urcorner),$ (3.73)
and $\mathsf{RA}$ proves this equivalence because $\mathrm{sub}$ is representable. See Rogers [1987]; Boolos et al. [2012]. ∎
4 From Proof Checking to Polynomial Constraints
4.1 The Hilbert–Ackermann Kernel
Intuition.
We fix a minimal proof system for propositional logic in the style of Hilbert and Ackermann [1938], with implication and falsity as primitive connectives. A proof is a finite sequence of formulas, each either an axiom instance or a consequence of Modus ponens from earlier lines, concluding at the target formula. A total boolean checker decides correctness of any proposed proof; soundness (Theorem 4.5) confirms it is faithful to derivability. Being primitive recursive, the checker arithmetizes directly via the encoding mechanisms of the preceding sections. We then reflect each of its verification conditions into polynomial equalities over $\mathbb{N}$, obtaining a finite constraint system whose solvability is equivalent to checker acceptance (Theorem 4.19) with every emitted equation of degree at most $3$ (Proposition 4.20).
Definition 4.1 (Formula syntax).
The set $\mathsf{Form}$ of formulas is defined inductively by:
$\mathsf{Form}\;::=\;\bot\mid(\varphi\to\psi),$ (4.74)
where $\bot$ denotes falsity and $\varphi\to\psi$ denotes implication. The mechanized kernel ($\mathtt{R04}$) uses exactly these two constructors; propositional variables are not needed, since the axiom schemas already quantify over all formulas.
Definition 4.2 (Axiom schemas).
The proof system admits three axiom schemas [Troelstra and Schwichtenberg, 2000]. The schemas K and S are the logical counterparts of the combinators $\mathbf{K}$ and $\mathbf{S}$. For all formulas $\varphi,\psi,\chi\in\mathsf{Form}$:
  1. (i)
    Schema K:
    $\varphi\to(\psi\to\varphi).$ (4.75)
  2. (ii)
    Schema S:
    $\big(\varphi\to(\psi\to\chi))\to((\varphi\to\psi)\to(\varphi\to\chi)\big).$ (4.76)
  3. (iii)
    Schema EFQ (“ex falso quodlibet”):
    $\bot\to\varphi.$ (4.77)
Definition 4.3 (Proof sequence).
A proof of a formula $\varphi$ is a finite sequence
$p=(\varphi_{0},\ldots,\varphi_{n-1})$ (4.78)
such that:
  1. (i)
    For each $i<n$, the formula $\varphi_{i}$ satisfies one of the following:
    1. (a)
      Axiom: $\varphi_{i}$ is an instance of one of the axiom schemas (Definition 4.2).
    2. (b)
      Modus ponens: There exist indices $j,k<i$ such that
      $\varphi_{k}=(\varphi_{j}\to\varphi_{i}).$ (4.79)
  2. (ii)
    The final line matches the target:
    $\varphi_{n-1}=\varphi.$ (4.80)
Definition 4.4 (Proof checker).
Define a total boolean function
$\mathrm{check}:\mathsf{ProofSeq}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\}$ (4.81)
that returns $\mathrm{true}$ if and only if $p$ satisfies the conditions of Definition 4.3 with target $\varphi$. The search for Modus ponens witnesses at each line $i$ is bounded by $i$, so $\mathrm{check}$ is total and primitive recursive.
Theorem 4.5 (Kernel soundness).
If $\mathrm{check}(p,\varphi)=\mathrm{true}$, then $\varphi$ is derivable in the Hilbert–Ackermann System, i.e.,
$\vdash\varphi.$ (4.82)
By induction on the length of the proof sequence $p$. Assume $\mathrm{check}(p,\varphi)=\mathrm{true}$. Base case. If $n=1$, then $\varphi_{0}$ is an axiom instance with $\varphi_{0}=\varphi$, hence derivable. Inductive step. Assume the claim holds for all shorter proof sequences. For each line $i<n$: (i) If $\varphi_{i}$ is an axiom, it is derivable. (ii) If $\varphi_{i}$ follows by Modus ponens from $\varphi_{j}$ and $\varphi_{k}=(\varphi_{j}\to\varphi_{i})$ with $j,k<i$, then by the inductive hypothesis both $\varphi_{j}$ and $\varphi_{j}\to\varphi_{i}$ are derivable, so $\varphi_{i}$ is derivable. Since $\varphi_{n-1}=\varphi$ and $\varphi_{n-1}$ is derivable, we conclude $\vdash\varphi$. ∎
4.2 Arithmetization
Proposition 4.6 (Decidability structure).
The provability evaluation
$\mathrm{Provable}(\varphi)\equiv\exists p\;\mathrm{check}(p,\varphi)=\mathrm{true}$ (4.83)
is $\Sigma^{0}_{1}$. When the proof system is instantiated for arithmetic, $\mathrm{Provable}$ is $\Sigma^{0}_{1}$-complete [Rogers, 1987].
Intuition.
Formulas are coded as natural numbers via Carryless Pairing (Section 3), and proof sequences are represented by carryless list codes with bounded extraction witnesses (Section 3.3). The proof checker $\mathrm{check}$ is therefore a primitive recursive relation over $\mathbb{N}$, and its defining conditions—axiom matching, Modus ponens witness search, target equality—are expressible as polynomial constraints.
4.3 Arithmetic Representation of Proof Data
Definition 4.7 (Proof witness assignment).
An assignment for a proof of length $n$ targeting formula code $t$ is a tuple
$a=(s,n,t,j_{0},k_{0},\ldots,j_{n-1},k_{n-1},a_{0},m_{0},\ldots,a_{n-1},m_{n-1},\ldots)$ (4.84)
where:
  1. (i)
    $s\in\mathbb{N}$ is the carryless list code for the proof sequence,
  2. (ii)
    $n\in\mathbb{N}$ is the proof length,
  3. (iii)
    $t\in\mathbb{N}$ is the target formula code,
  4. (iv)
    $j_{i},k_{i}\in\mathbb{N}$ are Modus ponens witness indices for line $i$,
  5. (v)
    $a_{i},m_{i}\in\{0,1\}$ are justification selector booleans (axiom vs. Modus ponens),
  6. (vi)
    additional auxiliary variables for bounds, line-extraction chains, and intermediate computations.
The assignment bundles all numerical witnesses required to verify proof correctness arithmetically.
4.4 Polynomial Reflection of Basic Tests
Lemma 4.8 (Equality constraint, degree $2$).
For $x,y\in\mathbb{N}$, the equality $x=y$ holds if and only if
$(x-y)^{2}=0.$ (4.85)
Over $\mathbb{N}$, $(x-y)^{2}=0$ implies xy=0x-y=0, hence $x=y$. Conversely, $x=y$ implies $(x-y)^{2}=0$ immediately. ∎
Lemma 4.9 (Boolean constraint, degree $2$).
A variable $b\in\mathbb{N}$ satisfies $b\in\{0,1\}$ if and only if
b(1b)=0.b(1-b)=0. (4.86)
If $b=0$ or $b=1$, then b(1b)=0b(1-b)=0. Conversely, b(1b)=0b(1-b)=0 forces either $b=0$ or $b=1$ over $\mathbb{N}$. ∎
Lemma 4.10 (Conjunction constraint, degree $2$).
For boolean variables $b_{1},b_{2}\in\{0,1\}$, the conjunction $b=b_{1}\land b_{2}$ is reflected by
$b=b_{1}b_{2}.$ (4.87)
Definition 4.11 (Axiom matching constraint family).
For each line index $i<n$, define a finite family $P_{\mathrm{axiom}}(a,i)$ of polynomial equalities whose simultaneous satisfaction indicates that the formula code $\mathrm{line}(s,i)$ matches one of the axiom schemas. The construction proceeds by pattern matching on the structural encoding of formulas: (i) extract subformula codes via line extraction and Carryless “unpairing”; (ii) enforce structural equality constraints using Lemma 4.8; (iii) check that the pattern matches $\mathsf{K}$, $\mathsf{S}$, or $\mathsf{EFQ}$ by verifying the implication nesting structure.
Proposition 4.12 (Axiom constraint degree).
Every member of $P_{\mathrm{axiom}}(a,i)$ has degree at most $2$.
Pattern matching and structural equality tests involve only arithmetic extraction (bounded carryless line extraction, unpairing) and quadratic equality constraints (Lemma 4.8). No cubic terms are introduced. ∎
Definition 4.13 (Modus ponens constraint family).
For each line index $i<n$, let $\mathcal{P}_{\mathrm{MP}}(a,i)$ be the finite family of polynomial equalities implementing the Modus ponens check. The family asserts the existence of witness indices $j_{i},k_{i}<i$ (included in the assignment $a$) such that
$\mathrm{line}(s,k_{i})=\mathrm{enc}_{\mathrm{form}}(\mathrm{line}(s,j_{i})\to\mathrm{line}(s,i)).$ (4.88)
It comprises: (i) Bound constraints enforcing $j_{i}<i$ and $k_{i}<i$ via four-squares representation; (ii) Structural equality constraints expressing that the implication code at index $k_{i}$ equals the pairing of codes at $j_{i}$ and $i$. Each member of $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree $\leq 2$.
Proposition 4.14 (Modus ponens constraint degree).
Every polynomial in $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree at most $2$.
Bound constraints use quadratic four-square encodings. Structural equality uses quadratic equality tests and the degree-$2$ pairing encoding. No cubic terms appear within any member of the family. ∎
4.5 Justification Constraint
Definition 4.15 (Justification disjunction).
For each line $i<n$, the justification condition requires that line $i$ is verified either as an axiom or via Modus ponens. This disjunction is encoded using boolean selector variables $a_{i},m_{i}\in\{0,1\}$ (part of the assignment $a$):
$a_{i}(1-a_{i})$ $=0,$ (4.89)
$m_{i}(1-m_{i})$ $=0,$ (4.90)
$a_{i}+m_{i}-1$ $=0,$ (4.91)
$a_{i}\cdot Q$ $=0\quad\text{for each }Q\in P_{\mathrm{axiom}}(a,i),$ (4.92)
$m_{i}\cdot Q$ $=0\quad\text{for each }Q\in\mathcal{P}_{\mathrm{MP}}(a,i).$ (4.93)
Constraints (4.89)–(4.90) enforce boolean values. Constraint (4.91) enforces exclusive selection. Constraints (4.92)–(4.93) activate the corresponding verification polynomial.
Proposition 4.16 (Justification constraint degree).
Every emitted justification constraint has degree at most $3$.
Constraints (4.89)–(4.91) are degree $\leq 2$. Each axiom constraint family member has degree $\leq 2$; activation $a_{i}\cdot Q=0$ raises degree to at most $3$. Each MP constraint is drawn from the family $\mathcal{P}_{\mathrm{MP}}(a,i)$ of degree $\leq 2$ polynomials; activation $m_{i}\cdot Q=0$ again yields degree $\leq 3$. No product of a selector with a degree-33 polynomial is emitted. ∎
4.6 Target Constraint
Definition 4.17 (Target matching polynomial).
The final line of the proof must match the target formula code $t$:
$(\mathrm{line}(s,n-1)-t)^{2}=0.$ (4.94)
This constraint has degree $2$.
4.7 The Complete Semantic Constraints
Definition 4.18 (Proof constraint system).
For a proof of length $n$ targeting formula code $t$, let $\mathcal{C}_{n,t}$ be the finite set of polynomial equalities consisting of:
  1. (i)
    For each $i<n$: boolean constraints for $a_{i},m_{i}$ and exclusivity $a_{i}+m_{i}-1=0$.
  2. (ii)
    Global extraction/decoding constraints (line extraction, Zeckendorf digit constraints) that define $\mathrm{line}(s,i)$ for all $i<n$; these are always active.
  3. (iii)
    For each $i<n$: activated axiom checks $a_{i}\cdot Q=0$ for every $Q$ in the axiom constraint family $P_{\mathrm{axiom}}(a,i)$.
  4. (iv)
    For each $i<n$: activated MP checks $m_{i}\cdot Q=0$ for every $Q$ in the MP constraint family $\mathcal{P}_{\mathrm{MP}}(a,i)$.
  5. (v)
    The target constraint (Definition 4.17).
No unguarded $P_{\mathrm{axiom}}(a,i)=0$ or $P_{\mathrm{MP}}(a,i)=0$ constraints are included; axiom- and MP-specific auxiliary constraints appear only multiplied by their selectors.
Remark.
The system $\mathcal{C}_{n,t}$ is the pre-aggregation constraint list emitted by $\mathtt{emit\_cubic\_system}$. It consists of finitely many cubic equations, each of degree at most $3$. The subsequent aggregation layer (Section 4.8) collapses this finite conjunction to a single cubic equation without increasing degree.
Theorem 4.19 (Constraint equivalence).
For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$,
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists a\in\mathbb{N}^{k}\;\bigwedge_{P\in\mathcal{C}_{n,t}}P(a)=0.$ (4.95)
By directions.
  1. Direction 1.
    Forward. Assume $\mathrm{check}(p,\varphi)=\mathrm{true}$. Construct an assignment $a$ by (i) Setting $s$ to the carryless list code of $p$ (Theorem 3.8), (ii) instantiating extraction-chain witnesses so that $\mathrm{line}(s,i)$ equals the $i$-th formula code, (iii) for each line $i$: if justified by axiom, set $a_{i}=1,m_{i}=0$; if justified by Modus ponens with witnesses $j,k$, set $a_{i}=0,m_{i}=1,j_{i}=j,k_{i}=k$, and (iv) setting all remaining auxiliary variables to satisfy bound and equality constraints. By construction, all constraints in $\mathcal{C}_{n,t}$ vanish under $a$.
  2. Direction 2.
    Reverse. Assume $a$ satisfies all constraints in $\mathcal{C}_{n,t}$. Extract the proof sequence via carryless line decoding: $\varphi_{i}=\mathrm{dec}_{\mathrm{form}}(\mathrm{line}(s,i))$ for $i<n$. The justification constraints ensure each line is either an axiom instance or follows by Modus ponens. The target constraint ensures $\varphi_{n-1}=\varphi$. Hence $\mathrm{check}(p,\varphi)=\mathrm{true}$.
Constraint Expanded equation $\boldsymbol{\delta}$ Mechanism
Pre-aggregation cubic system $\mathcal{C}_{n,t}$ (no selector activation)
Line extraction $t_{i,r}-\mathrm{pair}(h_{i,r},t_{i,r+1})=0$ 2 Carryless unpair
Tail initialization $t_{i,0}-s=0$ 1 Starts extraction chain
Line readout $f_{i}-h_{i,i}=0$ 1 Exposes $\mathrm{line}(s,i)$
Boolean digits $d_{i,\kappa}-d_{i,\kappa}^{2}=0$ 2 Forces $d_{i,\kappa}\in\{0,1\}$
Non-adjacency $d_{i,\kappa}\cdot d_{i,\kappa+1}=0$ 2 Zeckendorf constraint
Zeckendorf sum $f_{i}-\sum_{\kappa=2}^{K}F_{\kappa}d_{i,\kappa}=0$ 1 Linear combination
Axiom matcher $A_{i}(\vec{x})=0$ 2 Quadratic schema match
MP structural $M_{i}(\vec{x}):=f_{k_{i}}-\mathrm{pair}(f_{j_{i}},f_{i})=0$ 2 Implication encoding
Axiom $A_{i}(\vec{x})\cdot M_{i}(\vec{x})=0$ 3 Quadratic $\times$ linear disj.
Target check $f_{n-1}-g_{\mathrm{T}}=0$ 1 Linear equality
Table 1: Degree bookkeeping for the pre-aggregation cubic constraint system $\mathcal{C}_{n,t}$ after eliminating selector-based activation. Axiom and Modus ponens are combined into a single cubic justification equation $A_{i}\cdot M_{i}=0$, where $A_{i}$ is a quadratic axiom-matching obligation and $M_{i}$ is the (linear-in-the-encoding) MP structural equality. No boolean selectors and no exactly-one constraints are used. The subsequent mixed-radix aggregation layer collapses this finite conjunction to a single cubic equation without increasing degree.
Linear equalities are written as xy=0x-y=0 throughout; squaring is not needed for degree accounting and would only raise degrees.
Proposition 4.20 (Global degree bound).
Every polynomial in the constraint system $\mathcal{C}_{n,t}$ has degree at most $3$.
The mechanization certifies the bound syntactically. Each constraint is emitted as an $\mathtt{Expr}$ whose degree is computed by structural recursion in $\mathtt{R00\_\_Degree\_Framework}$. Axiom constraints are degree $\leq 2$, MP family members are degree $\leq 2$, and activation multiplies these by linear selectors, keeping every emitted equation at degree $\leq 3$. Linear equalities (xy=0x-y=0) are included in both channels to maintain consistent aggregation. The assembly theorem
$\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ (4.96)
proves that every emitted polynomial has syntactic degree $\leq 3$, independently of any assignment. Table 1 gives the expanded concrete bookkeeping that explains why degree $3$ suffices. ∎
Lemma 4.21 (AST degree correspondence).
For every emitted expression $e$, $\mathtt{degree}(e)$ computed on the AST equals the total polynomial degree of its denotation in the ring $\mathbb{N}[\vec{x}]$.
By structural induction on $e$: leaves $\mathtt{Const}$ and $\mathtt{Var}$ have degree 0 and $1$ respectively; constructors $\mathtt{Add},\mathtt{Sub}$ take the maximum of subdegrees; $\mathtt{Mult}$ adds subdegrees. This matches the algebraic definition of total degree. ∎
Intuition.
The reduction from proof checking to cubic solvability is uniform: given a target formula $\varphi$ with code $t$, the compiled single equation $\Phi(t,p,\vec{x})=0$ existentially quantifies over an unbounded proof code $p$; there is no computable proof-length bound. Degree $3$ arises only from selector activation of quadratic checks (e.g. axiom membership). No quadratic constraint is multiplied by a quadratic guard.
4.8 Mixed-Radix Aggregation
Intuition.
We collapse the finite list of cubic constraints without squaring by using standard base-$B$ expansion uniqueness. Each constraint is rewritten as $A_{i}-B_{i}$ with nonnegative channels. Capacity bounds enforce $0\leq A_{i},B_{i}<B$. Weighting channel $i$ by the constant $B^{i}$ yields a single equality whose base-$B$ digits are exactly the channel values, so equality of the weighted sums forces $A_{i}=B_{i}$ for every $i$. Choosing $B:=F_{K}$ (or $F_{K}+1$ for a weak inequality) keeps the Fibonacci aspect but the argument is ordinary mixed radix.
Lemma 4.22 (Two-channel decomposition).
Every polynomial $P(\vec{x})\in\mathbb{Z}[\vec{x}]$ is syntactically decomposed as P=ABP=A-B, where $A$ collects the positive-coefficient monomials and $B$ collects the absolute values of the negative-coefficient monomials. Over $\mathbb{N}$, both $A(\vec{x}),B(\vec{x})\geq 0$ on every assignment, and
$P(\vec{x})=0\quad\Longleftrightarrow\quad A(\vec{x})=B(\vec{x}).$ (4.97)
Degree does not increase: $\deg A,\deg B\leq\deg P$.
Definition 4.23 (Capacity constraints).
For each equality $A_{i}(\vec{x})=B_{i}(\vec{x})$ arising from a constraint $P_{i}=0$, fix an integer base $B\geq 2$ and enforce the digit bounds
$0\leq A_{i}(\vec{x})<B,\qquad 0\leq B_{i}(\vec{x})<B.$ (4.98)
Encode the inequalities via four-square slack variables (Lagrange):
$B-1-A_{i}(\vec{x})=u_{i,1}^{2}+u_{i,2}^{2}+u_{i,3}^{2}+u_{i,4}^{2},$ (4.99)
and similarly for $B_{i}$. The right-hand side is quadratic in the slack variables, so each capacity equation has total degree $\leq 3$. In the implementation we set $B:=F_{K}$ (or $F_{K}+1$ for non-strict bounds), but any fixed $B\geq 2$ suffices.
Lemma 4.24 (Mixed-radix aggregation).
Let $(A_{i}=B_{i})_{i=1}^{m}$ be equalities of nonnegative polynomials satisfying $0\leq A_{i},B_{i}<B$ for a fixed base $B\geq 2$. Then
$\sum_{i=1}^{m}B^{i}\,A_{i}=\sum_{i=1}^{m}B^{i}\,B_{i}\quad\Longleftrightarrow\quad\forall i,\;A_{i}=B_{i}.$ (4.100)
Under the digit bounds, every natural number has a unique representation $\sum_{i=1}^{m}B^{i}d_{i}$ with $d_{i}\in\{0,\ldots,B-1\}$. Apply uniqueness with digits $d_{i}:=A_{i}$ and $d_{i}:=B_{i}$ on the two sides. ∎
Lemma 4.25 (Aggregation correctness).
Given a finite list of cubic constraints $(L_{i}(\vec{x})=0)_{i=1}^{m}$, let $C_{1},\ldots,C_{r}$ be the associated capacity equalities. For each equality (original or capacity) write $E_{k}=A_{k}-B_{k}$ with $0\leq A_{k},B_{k}<B$. Define the aggregated polynomial
$\Phi(\vec{x},\vec{y}):=\sum_{k=1}^{m+r}B^{k}\,A_{k}(\vec{x},\vec{y})\;-\;\sum_{k=1}^{m+r}B^{k}\,B_{k}(\vec{x},\vec{y}).$ (4.101)
Then the mixed-radix construction outputs the single cubic equation $\Phi(\vec{x},\vec{y})=0$ such that
$\exists\,\vec{x}\;\bigwedge_{i=1}^{m}L_{i}(\vec{x})=0\quad\Longleftrightarrow\quad\exists\,\vec{x},\vec{y}\;\;\Phi(\vec{x},\vec{y})=0,$ (4.102)
where $\vec{y}$ comprises the slack variables for capacity bounds. Moreover, $\deg\Phi\leq 3$.
Extend the list by adding all capacity equalities; enumerate all equalities as $E_{1},\ldots,E_{m+r}$. Decompose each $E_{k}=A_{k}-B_{k}$ with digit bounds <B<B and form $\Phi$ as in (4.101). If all equalities hold, each $A_{k}=B_{k}$ and $\Phi=0$ follows. Conversely, if $\Phi=0$ then by Lemma 4.24 and the digit bounds we get $A_{k}=B_{k}$ for every $k$, so both the original constraints and the capacity equalities hold. Degree is preserved because only constant weights $B^{k}$ are introduced and capacity equations are cubic. ∎
Example 4.26.
Toy system over $\mathbb{N}$ with base $B=5$ (digit bound <5<5):
$P_{0}:x-2=0,\qquad P_{1}:y-1=0,\qquad P_{2}:x^{3}-8=0.$ (4.103)
All channels are 0 at any solution, so bounds are respected. Using weights $5^{1},5^{2},5^{3}$ gives the aggregated equality
$5(x-2)+25(y-1)+125(x^{3}-8)=0,$ (4.104)
i.e.
$125x^{3}+5x+25y-1035=0.$ (4.105)
By base-$5$ digit uniqueness and the digit bounds, this single cubic holds exactly when all three constraints hold.
Intuition.
The weighted sum encodes the channels as base-$5$ digits. Vanishing forces channelwise equality. No squaring is used, so the cubic degree bound is preserved.
Theorem 4.27 (Proof evaluation via single cubic).
For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$,
$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists\,\vec{z}\in\mathbb{N}^{k^{\prime}}:\Phi_{n,t}(\vec{z})=0,$ (4.106)
where $\Phi_{n,t}$ is the single cubic equation produced by banded aggregation of $\mathcal{C}_{n,t}$, and $\vec{z}$ extends the proof witness $\vec{x}$ by the capacity slack variables.
The forward direction follows from Theorem 4.19 and Lemma 4.25.
For the reverse direction, assume $\Phi_{n,t}(\vec{z})=0$. The single mixed-radix equation aggregates every constraint in $\mathcal{C}_{n,t}$, including all capacity equalities. Mixed-radix uniqueness forces each aggregated equality to hold separately, so in particular the digit bounds $0\leq A_{i},B_{i}<B$ hold, and every channel equality $A_{i}=B_{i}$ holds. Therefore all constraints in $\mathcal{C}_{n,t}$ vanish, and validation follows by Theorem 4.19. ∎
Proposition 4.28 (Degree bound).
The single equation $\Phi_{n,t}$ has degree at most $3$, established by structural induction on the emitted expression AST, independent of any satisfying assignment.
The aggregation multiplies channel polynomials only by constant powers $B^{i}$. No squaring of cubic expressions occurs. Capacity equations are degree $\leq 3$ (Definition 4.23), and the principal mixed-radix equality is a linear combination of degree-$\leq 3$ channels. Hence the final single equation has syntactic degree $\leq 3$, as certified in $\mathtt{aggregate\_banded\_degree\_le\_3}$. ∎
4.9 Compilation and Degree Control
Intuition.
Before diagonalization, we fix the compiler interface and the degree budget. The mechanized core provides a primitive recursive compiler to cubic systems with certified correctness and degree bound.
Definition 4.29 (Cubic system compiler).
Define a primitive recursive compiler
$\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial}),$ (4.107)
mapping a proof sequence and a target formula to a finite list of polynomial constraints over $\mathbb{N}$. The coded version, operating on $\mathbb{N}$, is $\mathtt{compiler\_primitive}$ ($\mathtt{R16}$).
Definition 4.30 (Single-equation compiler).
The endpoint
$\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to(\mathsf{Polynomial}\times\mathsf{Polynomial})$ (4.108)
emits the Fibonacci-banded single equality surface corresponding to $\mathcal{C}_{n,t}$. It composes $\mathtt{emit\_cubic\_system}$ with the banded aggregation of Section 4.8, packaging the principal banded equality and all capacity constraints into a single degree-$\leq 3$ equation. This aligns with $\mathtt{R13\_\_Kernel\_API}$.
Theorem 4.31 (Certified cubic reduction).
For every proof sequence $p$ and formula $\varphi$ so that every constraint in
$\mathtt{emit\_cubic\_system}(p,\varphi)$ (4.109)
has syntactic degree at most $3$, while proof evaluation is equivalent to cubic satisfiability:
$\mathrm{check}(p,\varphi)=\mathrm{true}\Leftrightarrow\exists a.\;\mathtt{all\_zero}\bigl(\mathtt{emit\_cubic\_system}(p,\varphi),\;a\bigr).$ (4.110)
The mechanized statement ($\mathtt{R13}$) is
$\mathtt{check\_iff\_emit\_cubic\_all\_zero}.$ (4.111)
By mechanization. In Rocq: degree bounds by structural induction on emitted constraints ($\mathtt{Constraints\_Assembly}$), and correctness by the checker–constraint equivalence ($\mathtt{Constraints\_MP}$). The reduction nucleus is $\mathtt{compiler\_primitive}$ in $\mathtt{R16}$. ∎
Remark.
The classical monomial-introduction construction of Jones [1982] remains visible as an alternative route from degree-$6$ back to degree-$3$ systems, but is not used in the present argument.
5 Main Diagonal Argument
5.1 On Local Reflection
Intuition.
Every nonexistence result in this section proceeds by modus tollens: assume the object exists and derive $\bot$. This is not a deficiency in our methodology but reflects the intrinsic difficulty of proving a negative statement. Consider, for example, how under the Brouwer–Heyting–Kolmogorov interpretation [Troelstra and van Dalen, 1988], absurdity admits no canonical realization. Accordingly, there is no direct witness of impossibility. To establish that a decider or any promissory procedure does not exist, the only known route is to show that any hypothetical witness yields a contradiction under suitable assumptions. Proving impossibility therefore amounts to demonstrating that the assumption is untenable within a consistency-preserving environment.
In our case, structural constraints determine the specific form of the transformation. Once the certified compiler of Theorem 4.31 reliably produces a system of cubic constraints for each proof target, while a two-channel aggregation (Lemma 4.25) witnesses the map from theoremhood to a single cubic equation without change in degree, we argue via an induced machine that could observe itself: the compiler $f$, composed with the aggregation, is itself primitive recursive and hence computable, such that any hypothetical total correct decider $D$ for $\mathsf{CubicSat}$ would compose with $f$ to decide $\mathsf{Thm}$. It is the existence of this induced oracle machine, not a specific impossible equation, that is refuted.
Ansatz 5.1 (Argument).
The argument is separated into three tiers:
  1. Tier 1.
    Register Arithmetic. A certified uniform compiler into a single cubic equation, yielding $\Sigma^{0}_{1}$-hardness.
  2. Tier 2.
    Fixed point logic. Assuming a total correct cubic decider, fixed-point self-reference computes a mirror cubic instance on which it fails.
  3. Tier 3.
    Internal collapse. If correctness is provable inside $\mathsf{RA}$ on the compiler image, one derives $\Sigma_{1}$-reflection and contradicts Gödel’s 2nd.
The contradiction mechanism in Tier 2 is purely diagonal and does not require $\mathsf{RA}$ to decide cubic satisfiability. Tier 3 is strictly stronger: internal certification of a correct procedure would force $\mathsf{RA}$ to prove its own $\Sigma_{1}$-reflection schema, contradicting consistency.
Intuition.
A “total machine” $\mathcal{M}$ that correctly identifies all unsolvable instances—reporting $M(e)=0$ whenever $\neg\mathsf{DiophEqSat}(e)$—is, together with $\Sigma_{1}$-enumeration of solutions, a total decider for $\mathsf{DiophEqSat}$. Detecting unsolvability and deciding solvability mutually imply each other, such that internally certifying inside $\mathsf{RA}$ that $\mathcal{M}$ correctly identifies unsolvable instances enables a $\Pi_{1}$-soundness schema on the cubic predicate:
$\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\neg\mathsf{DiophEqSat}(\bar{e})\urcorner\bigr)\;\to\;\neg\mathsf{DiophEqSat}(e).$ (5.112)
This is a form of Local Reflection Principle [Feferman, 1960]. We observe how $\mathcal{M}$ escalates to $\mathrm{Con}(\mathsf{RA})$, which is false in $\mathsf{RA}$ by Gödel’s 2nd. Any internally certified total primitive procedure for detecting unsolvability forces $\mathsf{RA}$ to prove its own consistency: a contradiction.
5.2 Fixed Point Logic
Theorem 5.2 (No total correct single-cubic decider).
Assume the mechanized compiler of Section 4.9: there is a primitive recursive map $f$ on sentence codes such that
$\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicSat}(f(u))\bigr),$ (5.113)
where $\mathsf{CubicSat}$ is satisfiability of a single degree-$\leq 3$ Diophantine equation over $\mathbb{N}$ (the system emitted by the compiler is collapsed to one equation by Fibonacci-banded aggregation, Lemma 4.25).
Since $\mathsf{CubicSat}$ is already the satisfiability of a single cubic equation, write $\mathsf{CubicEqSat}:=\mathsf{CubicSat}$. Then there is no total Turing Machine $D$ such that
$\forall s\;\bigl(D(s)=1\leftrightarrow\mathsf{CubicEqSat}(s)\bigr).$ (5.114)
Equivalently: $\mathsf{CubicEqSat}$ is not decidable by any total correct Turing Machine.
By Fixed point logic in steps. We consult Boolos et al. [2012]; Kleene [1952]
  1. Step 1.
    Fix the primitive recursive map $f$ witnessing (5.113). Since $\mathsf{CubicSat}$ is itself satisfiability of a single cubic equation (the compiler-emitted system is collapsed by Fibonacci-banded aggregation, Lemma 4.25), $f$ already maps sentence codes to single cubic equation codes. Define h(u):=f(u),h(u)\;:=\;f(u), observe that $h$ is primitive recursive and satisfies
    $\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))\bigr).$ (5.115)
    No degree blow-up or shielding step is required.
  2. Step 2.
    We assume a total correct single-cubic decider. For contradiction, there exists a total Turing Machine $D$ satisfying (5.114). Let $\mathsf{Out}_{D}(s)$ be the arithmetization:
    $D$ halts on input $s$ with output $1$.
    By totality of $D$, for every input $s$ we may identify
    $\mathsf{Out}_{D}(s)\;\leftrightarrow\;(D(s)=1).$ (5.116)
  3. Step 3.
    We diagonalize against $u\mapsto h(u)$ by considering the one-free-variable formula
    $\varphi(x)\;\equiv\;\neg\,\mathsf{Out}_{D}(h(x)).$ (5.117)
    By the Diagonal Lemma (Theorem 3.14), there exists a sentence $u_{D}$ such that
    $\mathsf{RA}\;\vdash\;u_{D}\leftrightarrow\neg\,\mathsf{Out}_{D}\bigl(h(\ulcorner u_{D}\urcorner)\bigr).$ (5.118)
    Let
    $s_{D}\;:=\;h(\ulcorner u_{D}\urcorner).$ (5.119)
    Since $h$ is primitive recursive, $s_{D}$ is uniformly computable from $D$.
  4. Step 4.
    Contradiction by exhaustive case analysis on $\mathsf{Out}_{D}(s_{D})$.
    Case 1. $\mathsf{Out}_{D}(s_{D})$, then by (5.118) $\neg\,\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, but by (5.116) and correctness (5.114), $\mathsf{CubicEqSat}(s_{D})$ holds, (5.115) yields $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, contradiction.
    Case 2. $\neg\,\mathsf{Out}_{D}(s_{D})$, then by (5.118), $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, hence (5.115) gives $\mathsf{CubicEqSat}(s_{D})$, and by correctness (5.114) plus totality (5.116) we obtain $\mathsf{Out}_{D}(s_{D})$, contradiction.
Both cases contradict, so the assumption (5.114) is false. Classically, no correct Turing Machine decides $\mathsf{CubicEqSat}$. ∎
Intuition.
The only $\mathsf{RA}$ argument required for the above is the fixed-point statement (5.118), obtained from the Diagonal Lemma for $\mathsf{RA}$ and the representability of the primitive recursive map $h$. The contradiction itself is external: it uses the totality/correctness assumption about $D$ to identify $\mathsf{Out}_{D}(s)$ with $D(s)=1$, and uses the reduction
$\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))$ (5.120)
to transfer arithmetic truth between arithmetized theoremhood and single-cubic satisfiability.
The hardness must therefore lie, as always, in the $\Pi\to\Sigma$ gap between verification and existence: $\mathsf{RA}$ can check any proposed witness by bounded polynomial evaluation, but the predicate $\exists\vec{x},P(\vec{x})=0$ escapes $\mathsf{RA}$’s deductive closure.
Corollary 5.3 (Undecidability for single Diophantine equations).
The predicate
$\mathsf{DiophEqSat}(e)\;\equiv\;\exists\,\vec{x}\in\mathbb{N}^{k}\;P_{e}(\vec{x})=0,$ (5.121)
where $e$ codes a single polynomial $P_{e}\in\mathbb{Z}[x_{1},\dots,x_{k}]$ (with $k$ part of the code), is $\Sigma^{0}_{1}$-complete and undecidable. In particular, the set of satisfiable single cubic equations over $\mathbb{N}$ (with arity encoded in the instance) is undecidable; over $\mathbb{Z}$ the corresponding degree bound is $6$ by difference encoding.
Since $\mathsf{DiophEqSat}\in\Sigma^{0}_{1}$, hardness and membership together yield $\Sigma^{0}_{1}$-completeness, and undecidability follows. ∎
Remark.
Representability of the compiler graph yields, for each numeral $u$, the internal clause $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner)$. This is the specific bridge used in the reflection tier; no global uniform biconditional inside $\mathsf{RA}$ is assumed.
5.3 Certification and Reflection
Definition 5.4 (Internal predicates).
We use the following arithmetized predicates (all are primitive recursive and representable in $\mathsf{RA}$):
  • $\operatorname{Sigma1Sent}(u)$: $u$ codes a $\Sigma_{1}$ sentence.
  • $\operatorname{Sat}_{\Sigma_{1}}(u)$: the $\Sigma_{1}$ sentence coded by $u$ is true (expanded as a bounded existential).
  • $\operatorname{Dioph}_{3}(e)$: the cubic polynomial coded by $e$ has a zero over $\mathbb{N}$ (same as $\mathsf{CubicEqSat}$).
  • $\operatorname{HaltOut}_{\mathcal{M}}(s,b)$: machine $M$ halts on input $s$ with output bit $b$.
  • $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$: $\forall u\,(\operatorname{Sigma1Sent}(u)\rightarrow(\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)))$.
Intuition.
$\operatorname{Prov}_{\mathsf{RA}}(u)$$u$ (code)Compiler fff(u)f(u)Decider $\mathcal{M}$$\operatorname{tr}$ (bridge) Internal reasoning in $\mathsf{RA}$ $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}\ \Longrightarrow\ \mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$ $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})\ \Longrightarrow\ \mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})\Longrightarrow\bot$
Figure 2: Internal certification loop. If $\mathsf{RA}$ could witness a machine $\mathcal{M}$ that decides the compiler image ($\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$), then (via the bridge $\operatorname{tr}$ and representability of $f$) provability transports to $\Sigma_{1}$-truth on that image, yielding $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathrm{Con}(\mathsf{RA})$, contradicting Gödel’s 2nd.
As said, certification of a cubic decider inside $\mathsf{RA}$ would not merely refute the decider—it would equip $\mathsf{RA}$ with the means to prove its own consistency by a Reflection Principle. The provability bridge transports $\mathsf{RA}$-proofs through the compiler $f$; the certified decision specification $\mathrm{Dec}_{\mathcal{M}}$ then closes the reflection loop (Figure 2).
Definition 5.5 (Internal decision specification).
Let $\operatorname{Sat}_{\Sigma_{1}}(u)$ be the standard $\Sigma_{1}$-satisfaction predicate and $f$ the primitive recursive compiler of Theorem 1.8. For a Turing Machine $\mathcal{M}$, define $\mathrm{Dec}_{\mathcal{M}}$ as
$\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{HaltOut}_{\mathcal{M}}(f(u),1)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.122)
The internal certification condition is $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$.
Theorem 5.6 (Provable decision on the compiler image yields $\Sigma_{1}$-reflection).
Assume $\mathsf{RA}$ proves the $\Sigma_{1}$-completeness of $f$:
$\mathsf{RA}\vdash\;\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{Sat}_{\Sigma_{1}}(u)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.123)
If $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ for some Turing Machine $\mathcal{M}$, then $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$.
Since $f$ is primitive recursive, its graph is $\mathsf{RA}$-representable (Proposition 1.10). Internalizing the compilation step yields a provability bridge: for each $\Sigma_{1}$-sentence code $u$,
$\mathsf{RA}\vdash\;\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner\bigr).$ (5.124)
Combining with $\mathrm{Dec}_{\mathcal{M}}$ and (5.123) yields $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)$, i.e. $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$. ∎
Corollary 5.7 (Incompleteness Discharge).
Assume $\mathsf{RA}$ is consistent with $\operatorname{Prov}_{\mathsf{RA}}$ satisfying the Hilbert–Bernays–Löb conditions (D1)–(D3) of Theorem 3.13. Then no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$: Theorem 5.6 would give $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$—contradicting Gödel’s 2nd Gödel [1931]; Feferman [1960]; Boolos et al. [2012]. If $\mathrm{Con}(\mathsf{RA})$ then $\nexists\mathcal{M}$, if $\lnot\mathrm{Con}(\mathsf{RA})$ then also $\nexists\mathcal{M}$.
The argument has three strictly separated tiers:
  1. Tier 1.
    Mechanized. Cubic compilation, degree bounds, and checker–constraint equivalence are verified in Rocq ($\mathtt{R00}$$\mathtt{R19}$). This tier also contains a code-level diagonal: the HBL conditions ($\mathtt{P01}$), fixed-point construction ($\mathtt{P02}$), and the coded undecidability theorem $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) are verified without classical logic ($\mathtt{P00}$$\mathtt{P06}$).
  2. Tier 2.
    Classical. Section 5.2 extends the result to arbitrary Turing Machines via metatheoretic LEM. This is the classical counterpart of $\mathtt{no\_total\_correct\_code\_CubicSat}$.
  3. Tier 3.
    Internal. $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ implies $\Sigma_{1}$-reflection (Theorem 5.6), contradicting Gödel’s 2nd (Corollary 5.7). This corresponds to $\mathtt{RA\_Certification}$ ($\mathtt{P04}$).
No tier conflates constructive compiler correctness with classical self-reference. ∎
6 Cubic Thresholds
6.1 On the Universal Cubic
Intuition.
The preceding construction already furnishes a uniform primitive recursive compiler from arithmetic sentences to single cubic equations, suggesting that universality at degree $3$ is not obstructed. Achieving a fixed universal cubic would require internalizing this compiler into one static cubic template while preserving strict degree control via band isolation and guard activation.
Conjecture 6.1 (Universal Cubic Conjecture).
There must be a finite equation,
$U(\vec{x},n)\in\mathbb{Z}[\vec{x},n]$ (6.125)
of total degree $\leq 3$ such that for every recursively enumerable set $W\subseteq\mathbb{N}$ there exists a parameter $n_{W}\in\mathbb{N}$ with
$k\in W\;\Longleftrightarrow\;\exists\vec{y}\in\mathbb{N}^{m}\;U(k,\vec{y},n_{W})=0.$ (6.126)
6.2 Cubic Upper Bound
Intuition.
The result obtained below is a quick and dense worst-case closure bound. It is not intrinsic to degree $3$, and one could certainly lower it by introducing only the multiplicative wiring actually used by the chosen quartic universal polynomial, as shown in Corollary 6.4.
Lemma 6.2 (Quartic-to-cubic closure).
For every $v\in\mathbb{N}$ there exists a uniform many–one reduction from quartic to cubic Diophantine satisfiability given by
$\mathrm{H}{\mathbb{N}}(4,v)\;\leq_{m}\;\mathrm{H}{\mathbb{N}}(3,\,v+\binom{v+1}{2}\,)$. (6.127)
Introduce auxiliary variables $p_{ij}(1\leq i\leq j\leq v)$ subject to $p_{ij}=x_{i}x_{j}.$ Each such constraint has degree $2$ and every quartic monomial $x_{a}x_{b}x_{c}x_{d}$ occurring in $P$ is rewritten as $x_{a}x_{b}x_{c}x_{d}\;\longmapsto\;p_{ab}\,p_{cd},$ applying a fixed pairing rule (e.g., $a\leq b$, $c\leq d$, $ab\leq cd$ lexicographically), let $\widetilde{P}(x,p)$ denote the resulting expression. (i) Each constraint $p_{ij}=x_{i}x_{j}$ has degree $2$ in $(x,p)$. (ii) Each rewritten quartic monomial $x_{a}x_{b}x_{c}x_{d}$ is replaced by the new monomial $p_{ab}\,p_{cd}$, which is quadratic in the auxiliary variables and therefore has total syntactic degree $2$ in the enlarged variable set $(x,p)$. (iii) All cubic and lower-degree terms of $P$ are unchanged, hence have degree $\leq 3$ in $(x,p)$. Therefore, every equation in the resulting system has total degree $\leq 3$. Solvability is preserved for $\exists x\in\mathbb{N}^{v}\;P(x)=0$ if and only if,
$\exists(x,p)\in\mathbb{N}^{v+\binom{v+1}{2}}\;\Bigl(\widetilde{P}(x,p)=0\;\wedge\;\forall\,i\leq j:\;p_{ij}=x_{i}x_{j}\Bigr).$ (6.128)
Corollary 6.3 (Trivial cubic threshold).
By Jones [1980], we have a “quartic universal” as a starting point:
$\mathrm{H}{\mathbb{N}}(4,58)={\rm U}.$ (6.129)
Hence, by Lemma 1.5 and Corollary 5.3 it can be said,
$\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}.$ (6.130)
Follows from Lemma 6.2. The number of auxiliary variables is
$\binom{59}{2}=\frac{59\cdot 58}{2}=1711.$ (6.131)
The total variable count is therefore $58+1711=1769$. Since $\mathrm{H}{\mathbb{N}}(4,58)={\rm U}$, Lemma 6.2 yields $\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}$. ∎
Corollary 6.4 (Sparsity).
Let $P$ be a fixed quartic universal polynomial over $58$ variables. Define $m_{2}(P)$ as:
Number of distinct pair-products $x_{i}x_{j}$ ($1\leq i\leq j\leq 58$) actually needed when rewriting every quartic monomial of $P$ as a product of two quadratic monomials under a fixed pairing rule.
Then,
$\mathrm{H}{\mathbb{N}}(3,58+m_{2}(P))={\rm U}.$ (6.132)
In particular,
$m_{2}(P)\leq\binom{59}{2}$, (6.133)
so this always improves or matches the dense bound $1769$.
Repeat the construction of Lemma 6.2, but introduce an auxiliary variable $p_{ij}$ only for pairs $(i,j)$ that actually appear in the chosen quartic-monomial decomposition of $P$. No other product variables are used in the rewritten system, so the number of added variables is exactly $m_{2}(P)$. The same back-substitution argument gives equivalence of solvability. ∎
6.3 Non-Enumerability of the Region
Theorem 6.5 (Region Impredicativity).
Define
$\mathcal{U}={(d,n)\mid\mathrm{H}{\mathbb{N}}(d,n)=\mathsf{U}}.$ (6.134)
Fix a standard Gödel Numbering $e\mapsto\Phi_{e}$ of Turing Machines and a uniform coding of polynomial instances of type $(d,n)$. Let $\mathrm{Sol}_{d,n}(x)$ mean
Instance $x$ has a solution in $\mathbb{N}$
Then, (i) the Membership $(d,n)\in\mathcal{U}$ is a $\Pi^{0}_{2}$ predicate, (ii) for each $d$ there exists a minimal threshold
$n^{*}(d)=\min{n\mid(d,n)\in\mathcal{U}}\in\mathbb{N}\cup{\infty},$ (6.135)
such that $n^{*}$ is not computable, (iii) hence $\mathcal{U}$ is not decidable.
By Reductio ad absurdum in steps. We consult Kleene [1952],
  1. Step 1.
    Fix Undecidability of $\mathrm{H}{\mathbb{N}}(d,n)$ as the statement where $\mathrm{Tot}_{d,n}(e)$ asserts totality of $\Phi_{e}$
    $\forall e;\neg\Big(\mathrm{Tot}*{d,n}(e)\wedge\forall x,\big(\Phi_{e}(x)=1\leftrightarrow\mathrm{Sol}*{d,n}(x)\big)\Big),$ (6.136)
    on instances of type $(d,n)$. Totality is $\Pi^{0}_{2}$, and correctness is arithmetic; hence the whole statement is $\Pi^{0}_{2}$ in $(d,n)$. This proves (i).
  2. Step 2.
    For fixed $d$, if $\mathrm{H}{\mathbb{N}}(d,n)$ is undecidable, then so is $\mathrm{H}{\mathbb{N}}(d,n^{\prime})$ for all $n^{\prime}\geq n$. Hence each degree-slice of $\mathcal{U}$ is upward closed in (n). Therefore the minimal threshold $n^{*}(d)$ is well-defined in $\mathbb{N}\cup{\infty}$.
  3. Step 3.
    Suppose toward contradiction that $n^{*}$ is computable, observe how the predicate $(d,n)\in\mathcal{U}\Longleftrightarrow n\geq n^{*}(d)$ becomes decidable. Hence the entire undecidability region would admit a computable boundary, in that $d\longmapsto n^{*}(d)$ encodes the minimal variable count at which equations admit arbitrary $\Sigma^{0}_{1}$ predicates. Determining $n^{*}(d)$ would admit whether a fixed universal simulation can be compressed to $(d,n)$. Contradiction. More concretely: from an index $e$, one could effectively construct a degree $d_{e}$ encoding of $\Phi_{e}$ whose minimal variable count reflects whether $\Phi_{e}$ is total. If $n^{*}$ were computable, totality of arbitrary machines would be decidable. Hence $n^{*}$ cannot be computable. This establishes (ii), (iii) follows immediately.
Intuition.
The present results formalize the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative. The degree-$3$ threshold is not a combinatorial accident: it is precisely where the logical resources required to define undecidability coincide with those required to cause it. The guard gadget introduces the minimal impredicative coupling—asserting existence of a positive witness without explicit construction—and it is this coupling that both raises the degree from $2$ to $3$ and enables the encoding of self-referential constraints.
This has a structural consequence. Any formal specification of the “first” degree $k$ at which undecidability arises, if carried out inside a theory strong enough to internalize the reduction of Theorem 1.8, would interact nontrivially with the same reflection mechanism. The act of precisely defining the threshold is, by the reflection route of Theorem 5.6 and Corollary 5.7, the act of crossing it. Whether $\mathcal{U}$ is itself recursively enumerable remains ope. What the present construction does establish is that the boundary resists specification from within any theory strong enough to formalize the reduction. The question,
At what point does the undecidability of undecidability begin?
is by the constructions shown, an instance of the very phenomenon it tries to pin down. The self-referential structure of a cubic closure principle suggests a stronger phenomenon. The reduction of Section 5 encodes itself at degree $3$: the very mechanism that witnesses $\Sigma^{0}_{1}$-completeness—the guard gadget, the proof-to-polynomial compiler, the diagonal fixpoint—is itself a cubic construction. One cannot “see” the cubic threshold without building the cubic encoding that establishes it. This motivates the following conjecture, understood less as a claim than as a structural bet, the act of precisely locating the undecidability boundary already require crossing it. We close with a bet that we cannot win, only lose by falsification.
Conjecture 6.6 (Diophantine Gap Conjecture).
Let $\operatorname{T}$ be a consistent recursively axiomatized extension of weak arithmetic capable of representing primitive recursive functions and its own provability predicate.
We suspect that $\operatorname{T}$ will not emit a $\rm D$} nor $\rm U$} for $\varphi$ in some arithmetized sentence expressing $\mathrm{H}{\mathbb{N}}(3,2)=\varphi$.
Acknowledgments

### v7

@@
 Abstract
-Following a mechanization in Rocq, we prove that satisfiability of single Diophantine equations of total degree $\leq 3$ over $\mathbb{N}$ (with a computably bounded number of variables depending on the instance) is $\Sigma^{0}_{1}$-complete, and hence undecidable, in the sense of Hilbert’s Tenth Problem.
-The core construction is a uniform primitive recursive compiler mapping arithmetic sentence codes to finite structured cubic constraint systems, such that provability in $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ is equivalent to solvability of the emitted system.
-Conjunction is eliminated without squaring: all emitted constraints are merged into a single cubic Diophantine equation via a two-channel equality that preserves the cubic bound without sum-of-squares.
-All syntactic and inference checks compile to quadratic constraints; the only cubic terms come from linear selector variables multiplying those quadratic obligations.
-The mechanized core is constructive and requires no classical axioms: uniform computability, checker–constraint equivalence, and the degree-$3$ bound are verified by structural induction.
-From this base, undecidability is obtained by a classical Diagonal Argument: any total correct decider for cubic satisfiability would compose with the compiler to decide arithmetic theoremhood.
+We prove that satisfiability of a single Diophantine equation of total degree $\leq 3$ over $\mathbb{N}$, with the number of variables encoded in the input instance, is $\Sigma^{0}_{1}$-complete and hence undecidable.
+Concretely, we give a uniform primitive recursive compiler mapping arithmetic sentence codes to cubic Diophantine constraint systems such that provability in $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ (restricted to $\mathcal{L}_{\mathsf{RA}}$) is equivalent to solvability of the emitted constraints.
+Degree control is local: all syntactic checks and inference verification compile to equations of degree $\leq 2$, and the only degree-$3$ terms occur when a linear selector activates a quadratic obligation.
+The resulting finite system is then collapsed to a single cubic equation by a degree-preserving aggregation construction.
+We further internalize the compiler into a fixed four-instruction register kernel with bounded deterministic trace unrolling and a two-level aggregation layout, yielding a single explicit universal cubic polynomial in a fixed number of variables.
+The concrete coefficient table and a kernel-checked degree certificate are mechanized in Rocq; the semantic bridge relating arithmetic theoremhood to solvability of the pinned polynomial is proved in a separate layer without normalizing the coefficient table.
+Undecidability follows by diagonalization.
 1 Introduction
 1.1 Register Semantics
-Intuition.
-CircuitRegisterALUMemoryInputOutput I/O realization in Register Arithmetic ($\mathsf{RA}$) Mechanization yields mathematical programs.
-$\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ ISA–logic duality witnesses effectivity of weak arithmetics.
-Figure 1: Execution traces as finite sequences of configurations, understood arithmetically as coded finite objects, within a Von Neumann Architecture.
+Remark 1.1.1.
+[…]
+Figure 1: Execution traces as finite sequences of configurations, understood arithmetically as coded finite objects, e.g.
+within a Von Neumann Architecture.
 Each step requires only Register Arithmetic $\mathsf{RA}=(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}$.
-That certain forms of Diophantine satisfiability are $\Sigma_{1}$-complete is “old news.” One obtains this by composing MRDP [Matiyasevich, 1993] with reduction techniques such as Shielding [Jones, 1980].
-However, these constructions typically achieve degree bounds through auxiliary padding and simulation gadgets; the resulting systems encode computation faithfully, but the structural complexity reflects the mechanics of simulation rather than any intrinsic proof-theoretic invariant.
-On the other hand, mathematical practice and computer science have accumulated a vast repertoire of methods, ranging from ad hoc constructions to fully abstract algorithms.
-Yet whenever any such method is carried out, it is ultimately compiled into machine code (Figure 1) and executed as a finite sequence of operations over a fixed Instruction Set Architecture (ISA).
-This observation is decisive for a problem as intrinsically tied to computability as Hilbert’s 10th.
-Every computationally enumerable language is recognized by some finite program over a fixed instruction set; hence any reduction establishing $\Sigma_{1}$-completeness must, implicitly or explicitly, encode such a finite execution trace.
-In this sense, Diophantine representations inherit a structural bottleneck common to all computably enumerable sets.
-Accordingly, the appropriate framework should serve simultaneously as a language of machines and of recursion, making explicit the finite operational trace underlying existential search.
-At its core, it should require no more than bounded arithmetic over addition—computationally no stronger than an abacus.
-Definition 1.1 ($\mathsf{RA}$).
-We define Register Arithmetic as a specific fragment of first-order arithmetic. The language comprises of
-$\mathcal{L}_{\mathsf{RA}}\;=\;\{0,\,S,\,+,\,=\},$ (1.1)
+That certain forms of Diophantine satisfiability are $\Sigma_{1}$-complete is “old news,” trivially obtained by composing MRDP [Matiyasevich, 1993] with reduction techniques such as Shielding [Jones, 1980].
+However, these constructions achieve a complexity that reflects the mechanics of simulation rather than of a proof-theoretic invariant.
+Any reduction establishing $\Sigma_{1}$-completeness must encode a finite execution trace (Figure 1).
+The appropriate framework should therefore serve simultaneously as a language of machines and of recursion, requiring no more than bounded arithmetic over addition.
+Definition 1.1.2 ($\mathsf{RA}$).
+We define Register Arithmetic as a specific fragment of first-order arithmetic. The language comprises
+$\mathcal{L}_{\mathsf{RA}}\;=\;\{0,\,S,\,+,\,=\},$ (1)
 with constant 0, unary successor $S$, binary operation ++, and equality. The axioms of $\mathsf{RA}$ are:
   1. (i)
-    Successor axioms reminiscent of Robinson Arithmetics,
-    $S(x)\neq 0,$ (1.2)
-    $S(x)=S(y)\to x=y,$ (1.3)
-    $x\neq 0\to\exists y\;x=S(y),$ (1.4)
-    $x+0=x,$ (1.5)
-    $x+S(y)=S(x+y).$ (1.6)
+    Successor axioms reminiscent of Robinson Arithmetic,
+    $S(x)\neq 0,$ (2)
+    $S(x)=S(y)\to x=y,$ (3)
+    $x\neq 0\to\exists y\;x=S(y),$ (4)
+    $x+0=x,$ (5)
+    $x+S(y)=S(x+y).$ (6)
   2. (ii)
     A bounded induction schema for $\Delta_{0}$-formulae $\varphi(x,\vec{y})$,
-    $\Big(\varphi(0,\vec{y})\;\wedge\;\forall x\,\big(\varphi(x,\vec{y})\to\varphi(S(x),\vec{y})\big)\Big)\;\to\;\forall x\,\varphi(x,\vec{y}).$ (1.7)
+    $\Big(\varphi(0,\vec{y})\;\wedge\;\forall x\,\big(\varphi(x,\vec{y})\to\varphi(S(x),\vec{y})\big)\Big)\;\to\;\forall x\,\varphi(x,\vec{y}).$ (7)
   3. (iii)
     A collection schema $\mathrm{B}\Sigma_{1}$, $\varphi(x,y,\vec{z})$,
-    $\forall u\;\big(\,\forall x\leq u\;\exists y\;\varphi(x,y,\vec{z})\;\to\;\exists v\;\forall x\leq u\;\exists y\leq v\;\varphi(x,y,\vec{z})\,\big).$ (1.8)
+    $\forall u\;\Big(\,\forall x\leq u\;\exists y\;\varphi(x,y,\vec{z})\;\to\;\exists v\;\forall x\leq u\;\exists y\leq v\;\varphi(x,y,\vec{z})\,\Big).$ (8)
 Thus $\mathsf{RA}$ is the restriction of $\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1}$ to the multiplication-free language $\mathcal{L}_{\mathsf{RA}}$:
-$\mathsf{RA}\;=\;(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}.$ (1.9)
-Multiplication ($\mathtt{Mult}$) must not be a primitive of $\mathsf{RA}$ but it must be representable as a $\Sigma_{1}$-relation via iterated addition (Theorem 1.2).
+$\mathsf{RA}\;=\;(\mathrm{I}\Delta_{0}+\mathrm{B}\Sigma_{1})\upharpoonright\mathcal{L}_{\mathsf{RA}}.$ (9)
+Multiplication ($\mathtt{Mult}$) is not a primitive of $\mathsf{RA}$ but is representable as a $\Sigma_{1}$-relation via iterated addition (Theorem 1.1.3).
 Via this representation, $\mathsf{RA}$ interprets $\mathsf{Q}$ and is interpreted in $\mathsf{EFA}$:
-$\mathsf{Q}\;\leq_{\mathrm{int}}\;\mathsf{RA}+\mathtt{Mult}\;\leq_{\mathrm{int}}\;\mathsf{EFA}.$ (1.10)
-The set must be also recursive, in that membership of any sentence in the axiom set is decidable. Three meta-properties of $\mathsf{RA}$ are used in the sequel:
+$\mathsf{Q}\;\leq_{\mathrm{int}}\;\mathsf{RA}+\mathtt{Mult}\;\leq_{\mathrm{int}}\;\mathsf{EFA}\;<\;\mathsf{PA}$ (10)
+The axiom set must also be recursive, in that membership of any sentence is decidable. Three meta-properties of $\mathsf{RA}$ are used in the sequel:
   1. (i)
-    Representability. Each primitive recursive function or relation used in this development is numeralwise representable in $\mathsf{RA}$ (Proposition 1.10).
+    Representability. Each primitive recursive function or relation used in this development is numeralwise representable in $\mathsf{RA}$ (Proposition 1.3.3).
   2. (ii)
@@
     Derivability. The chosen proof predicate
-    $\operatorname{Prov}_{\mathsf{RA}}(u)\equiv\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u)$ (1.11)
-    must satisfy the Hilbert–Bernays–Löb conditions (see Theorem 3.13).
-Theorem 1.2 (Representability of multiplication).
-One of our first theorems is multiplication, defined as
-$\mathtt{Mult}(x,y,z).$ (1.12)
+    $\operatorname{Prov}_{\mathsf{RA}}(u)\equiv\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u)$ (11)
+    must satisfy the Hilbert–Bernays–Löb conditions (see Theorem 3.5.1).
+Theorem 1.1.3 (Representability of multiplication).
+Our first theorem establishes the representability of multiplication, defined as
+$\mathtt{Mult}(x,y,z).$ (12)
 This is the $\Sigma_{1}$-formula expressing that there exists a bounded register computation witnessing $z$ as the result of iterated addition of $x$, repeated $y$ times.
 Concretely, $\mathtt{Mult}(x,y,z)$ asserts the existence of a finite sequence
-$t_{0},\;\ldots,\;t_{y}$ (1.13)
+$t_{0},\;\ldots,\;t_{y}$ (13)
 such that
-$t_{0}=0,\quad t_{i+1}=t_{i}+x\;\text{ for }\;i<y,\quad t_{y}=z.$ (1.14)
-This relation is primitive recursive and therefore representable in $\mathsf{RA}$ (Proposition 1.10).
+$t_{0}=0,\quad t_{i+1}=t_{i}+x\;\text{ for }\;i<y,\quad t_{y}=z.$ (14)
+By construction.
+The relation is primitive recursive and therefore representable in $\mathsf{RA}$ (Proposition 1.3.3).
 Verification of a proposed witness trace $(t_{0},\ldots,t_{y})$ requires only $\Delta_{0}(\mathcal{L}_{\mathsf{RA}})$ checks on sequence-encoded entries, so $\mathtt{Mult}$ fits the verification floor of $\mathsf{RA}$: the theory can verify any multiplication witness but does not have multiplication as a function symbol.
@@
 In polynomial constraints over $\mathbb{N}$ (the Diophantine object language), the symbol $\cdot$ retains its standard meaning as ring multiplication.
-Intuition.
-The above are implicit in the foundational theory of bounded arithmetic [Hájek and Pudlák, 1993; Buss, 1997; Cook and Nguyen, 2010].
-We make their role explicit because the theory is the calibration point of this work: the weakest arithmetic in which bounded verification of finite objects—proof traces, polynomial assignments, encoded sequences—is formalizable.
-All internal reasoning proceeds by inspecting given encoded witnesses; no uniform totality claims are available.
-The theory verifies correctness of given finite witnesses; it does not certify that witnesses exist for all instances.
-This distinction is visible already at the polynomial level.
-For any fixed assignment $\vec{x}^{*}$ and $P$ of bounded degree, the equality $P(\vec{x}^{*})=0$ is a bounded check formalizable in $\mathsf{RA}$.
-The existential statement $\exists\,\vec{x}\;P(\vec{x})=0$ is $\Sigma_{1}$ and lies strictly beyond such instance verification.
-The cubic threshold arises naturally from this gap.
-All structural and syntactic verification—axiom matching, formula decoding, target equality—compiles to quadratic constraints (degree $\leq 2$), while escalation arises from enforcing bounded existential witnesses.
-One standard schematic guard uses
-$u=1+v^{2},\quad\exists w,(\mathtt{Mult}(u,E,w)\wedge w=0),$ (1.15)
-where $\mathtt{Mult}$ is the $\Sigma_{1}$-representation of multiplication by iterated addition (Theorem 1.2).
-This guard does not invoke multiplication as a primitive; rather, it demands a witness trace of iterated addition that $\mathtt{RA}$ can verify via $\Delta_{0}$ checks on the encoded sequence.
-The guard enforces that either $E=0$ or the square term forces strict positivity.
-Eliminating the bounded existential therefore introduces exactly one multiplicative interaction and raises the degree from $2$ to $3$.
-In the implemented constraint system, the actual cubic terms come from degree-$1$ selector variables multiplying quadratic obligations; the guard template is kept only as a conceptual schematic and is not multiplied by any quadratic constraint.
-Lemma 1.3 (Guard correctness and degree).
+∎
+Remark 1.1.4.
+These facts are implicit in bounded arithmetic [Hájek and Pudlák, 1993; Buss, 1997; Cook and Nguyen, 2010].
+We make them explicit via the calibration point of this work: $\mathsf{RA}$ is certainly an extremely weak arithmetic in which bounded verification of finite objects is barely formalizable: it can verify correctness of given finite witnesses, yet it cannot certify witness existence for all instances.
+The cubic threshold arises from this gap.
+Syntactic verification—axiom matching, formula decoding, target equality—compiles to quadratic constraints (degree $\leq 2$).
+Escalation to degree $3$ occurs when a degree-$1$ selector variable activates a quadratic obligation.
+In the implemented system, these selector products are the only source of cubic terms.
+Lemma 1.1.5 (Guard correctness and degree).
 Let $E(\vec{x})$ be a polynomial of degree $\leq 1$ and define $u=1+v^{2}$. Then
-$u\cdot E(\vec{x})=0\;\Longleftrightarrow\;E(\vec{x})=0,$
+$u\cdot E(\vec{x})=0\;\Longleftrightarrow\;E(\vec{x})=0,$ (15)
 because $u\geq 1$ for all $v$.
@@
   3. (iii)
-    Degree $4$: Jones [1980, 1982] showed that every recursively enumerable set is the solution set of a degree-44 Diophantine equation with 5858 variables.
+    Degree $4$: Jones [1980, 1982] showed that every recursively enumerable set is the solution set of a degree-$4$ Diophantine equation with $58$ variables.
 Undecidability follows classically from MRDP [Robinson et al., 1961; Matiyasevich, 1970; Davis, 1973; Matiyasevich, 1993].
@@
   5. (v)
-    The existence of a single universal cubic Diophantine equation (in a fixed number of variables) remains open; the present work establishes $\Sigma^{0}_{1}$-completeness of single cubic equations with computably bounded variable count via a direct proof-to-polynomial compiler, and attests to their constructibility.
-Gasarch [2021] (cf. Definition 1.4 for notation) summarizes the state of the problem:
+    The present work constructs a single universal cubic Diophantine equation in 9,6929{,}692 variables (Definition 6.1.1), establishes $\Sigma^{0}_{1}$-completeness of single cubic equations via a direct proof-to-polynomial compiler, and verifies the degree bound and universality bridge in Rocq.
+Gasarch [2021] (cf. Definition 1.2.1 for notation) summarizes the state of the problem:
 “Work on showing $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ or $\mathrm{H}{\mathbb{Z}}(d,n)={\rm U}$ seems to have stalled.
@@
 Perhaps computer-work could help […]” [Gasarch, 2021]
-The existence of a universal Diophantine equation of degree $3$ in a fixed number of variables sits in the open region between the fully understood quadratic case and the high-degree expressive universality constructions.
+The existence of a universal Diophantine equation of degree $3$ in a fixed number of variables was an open problem sitting between the fully understood quadratic case and the high-degree universality constructions; the present work resolves it (Section 7).
 Jones explicitly singled out Degree $3$ (over $\mathbb{N}$):
@@
 1.2 Problem Semantics
-Definition 1.4 ($\mathrm{H}{\mathbb{N}}(d,n)$ notation).
+Definition 1.2.1 ($\mathrm{H}{\mathbb{N}}(d,n)$ notation).
 We fix Gasarch’s convenient notation (from the excerpt) for parameterized Diophantine solvability throughout. For $d,n\in\mathbb{N}$:
@@
     $\mathrm{H}{\mathbb{N}}(d,n)$ is the decision problem: “Given
-    $P\in\mathbb{Z}[x_{1},\ldots,x_{k}]\;\text{with}\;k\leq n\;\text{and}\;\deg P\leq d,$ (1.16)
-    does $P(\vec{x})=0$ have a solution in $\mathbb{N}^{k}$?” While $\mathrm{H}{\mathbb{Z}}(d,n)$ is the analogous problem with solutions in $\mathbb{Z}^{k}$.
+    $P\in\mathbb{Z}[x_{1},\ldots,x_{k}]\;\text{with}\;k\leq n\;\text{and}\;\deg P\leq d,$ (16)
+    does $P(\vec{x})=0$ have a solution in $\mathbb{N}^{k}$?” The analogous problem over $\mathbb{Z}^{k}$ is denoted $\mathrm{H}{\mathbb{Z}}(d,n)$.
   2. (ii)
     $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ (resp. $\rm U$}) means the problem is decidable (resp. undecidable). Similarly for $\rm H$$\mathbb{Z}$.
-Lemma 1.5 (Transfer).
+Lemma 1.2.2 (Transfer).
 Since any $\mathbb{N}$-decider can be used to decide $\mathbb{Z}$ by sign enumeration,
-$m\in\mathbb{N}\;\Longleftrightarrow\;\exists\,x_{1},x_{2},x_{3}\in\mathbb{Z}:\;m=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3}$ (1.17)
+$m\in\mathbb{N}\;\Longleftrightarrow\;\exists\,x_{1},x_{2},x_{3}\in\mathbb{Z}:\;m=x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3}$ (17)
 so that it embeds $\mathbb{N}$-witnesses into $\mathbb{Z}$-witnesses. Replacing each $x_{i}$ in $P$ by
-$x_{i1}^{2}+x_{i2}^{2}+x_{i3}^{2}+x_{i3}$ (1.18)
+$x_{i1}^{2}+x_{i2}^{2}+x_{i3}^{2}+x_{i3}$ (18)
 yields a polynomial over $\mathbb{Z}$ of degree $2d$ in $3n$ variables, so ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ implies $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$.
@@
     By monotonicity in parameters, if $d^{\prime}\leq d$ and $n^{\prime}\leq n$,
-    $\mathrm{H}{\ast}(d,n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm D},$ (1.19)
-    $\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d,n)={\rm U}.$ (1.20)
+    $\mathrm{H}{\ast}(d,n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm D},$ (19)
+    $\mathrm{H}{\ast}(d^{\prime},n^{\prime})={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\ast}(d,n)={\rm U}.$ (20)
   2. (ii)
     By $\mathbb{N}\Longrightarrow\mathbb{Z}$ sign expansion,
-    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(d,n)={\rm D},$ (1.21)
-    ${\rm H}\mathbb{Z}(d,n)={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm U}.$ (1.22)
+    $\mathrm{H}{\mathbb{N}}(d,n)={\rm D}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(d,n)={\rm D},$ (21)
+    ${\rm H}\mathbb{Z}(d,n)={\rm U}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm U}.$ (22)
   3. (iii)
     $\mathbb{Z}\Rightarrow\mathbb{N}$ via three–square encoding,
-    ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm D},$ (1.23)
-    $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(2d,3n)={\rm U}.$ (1.24)
+    ${\rm H}\mathbb{Z}(2d,3n)={\rm D}$ $\;\Longrightarrow\;\mathrm{H}{\mathbb{N}}(d,n)={\rm D},$ (23)
+    $\mathrm{H}{\mathbb{N}}(d,n)={\rm U}$ $\;\Longrightarrow\;{\rm H}\mathbb{Z}(2d,3n)={\rm U}.$ (24)
 It is easy to see that undecidability propagates upward in $(d,n)$, and decidability downward, while the only domain transfers requiring parameter growth are those encoding $\mathbb{Z}$ in $\mathbb{N}$ or $\mathbb{N}$ in $\mathbb{Z}$.
-By logic.
+Immediate.
 Monotonicity follows since enlarging $(d,n)$ enlarges the input class, while restricting $(d,n)$ shrinks it.
@@
 Three–square encoding replaces $x\in\mathbb{N}$ by
-$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3},$ (1.25)
+$x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{3},$ (25)
 doubling degree and tripling variables. The undecidability statements are contrapositives. Downward transfer fails because no general compression of degree or variables preserves solvability. ∎
-Lemma 1.6 (Systems and general aggregation).
-For $d,n\in\mathbb{N}$ and $\ast\in\{\mathbb{N},\mathbb{Z}\}$, let $\mathrm{H}{\ast,S}(d,n)$ denote the decision problem: “Given a finite system
-$P_{1}(\vec{x})=0,\ldots,P_{m}(\vec{x})=0$ (1.26)
-with $k\leq n$ variables and $\deg P_{i}\leq d$ for all $i$, determine whether there exists $\vec{x}\in\ast^{k}$ satisfying all equations.” All implications of Lemma 1.5 hold verbatim with $\mathrm{H}{\ast,S}(d,n)$ in place of $\mathrm{H}{\ast}(d,n)$.
-Over $\mathbb{N}$ and $\mathbb{Z}$, sum-of-squares aggregation
-$P_{1}=\cdots=P_{m}=0\;\Longleftrightarrow\;\sum_{i=1}^{m}P_{i}(\vec{x})^{2}=0$ (1.27)
+Lemma 1.2.3 (Systems and general aggregation).
+For $d,n\in\mathbb{N}$ and $\ast\in\{\mathbb{N},\mathbb{Z}\}$, let $\mathrm{H}{\ast,S}(d,n)$ denote the decision problem:
+“Given a finite system
+$P_{1}(\vec{x})=0,\ldots,P_{m}(\vec{x})=0$ (26)
+with $k\leq n$ variables and $\deg P_{i}\leq d$ for all $i$, determine whether there exists $\vec{x}\in\ast^{k}$ satisfying all equations.”
+All implications of Lemma 1.2.2 hold verbatim with $\mathrm{H}{\ast,S}(d,n)$ in place of $\mathrm{H}{\ast}(d,n)$. Over $\mathbb{N}$ and $\mathbb{Z}$, sum-of-squares aggregation
+$P_{1}=\cdots=P_{m}=0\;\Longleftrightarrow\;\sum_{i=1}^{m}P_{i}(\vec{x})^{2}=0$ (27)
 reduces systems to one equation trivially, but doubles the degree bound ($d\mapsto 2d$). This general collapse is therefore not suitable for preserving a cubic degree bound.
-Proposition 1.7 (Aggregation for emitted systems).
-Our first contribution will be to show that valid systems emitted by the compiler of Theorem 4.31 admit an effective map producing a single equation of degree $\leq 3$ with the same solution set (up to harmless slack variables for band capacities).
+Proposition 1.2.4 (Aggregation for emitted systems).
+Our first contribution will be to show that valid systems emitted by the compiler of Theorem 4.9.4 admit an effective map producing a single equation of degree $\leq 3$ with the same solution set (up to harmless slack variables for band capacities).
 The construction rewrites each constraint as an equality of two nonnegative polynomials and merges all equalities into a single cubic Diophantine equation using Fibonacci-banded Zeckendorf isolation (Section 4.8); no sum-of-squares blow-up is used.
 1.3 Ansatz
-Ansatz 1.8 (Target).
-Our target is to prove that the predicate
-$\mathsf{CubicEqSat}(e):\equiv\exists\vec{x}\;P_{e}(\vec{x})=0,$
-where $e$ codes a single polynomial of total degree $\leq 3$ with its variable count as part of the code, is $\Sigma^{0}_{1}$-complete.
+Ansatz 1.3.1 (Target).
+We fix the following as the canonical decision predicate throughout:
+$\mathsf{CubicEqSat}(e):\equiv\exists\vec{x}\;P_{e}(\vec{x})=0,$ (28)
+where $e$ codes a single polynomial of total degree $\leq 3$ with its variable count as part of the code.
+All completeness and undecidability statements in this paper are formulated for $\mathsf{CubicEqSat}$ in the $R=0$ form; alternative surfaces (such as the pair surface $P=Q$ used in the mechanized kernel) are connected by Lemma 2.2.3.
+Our target is to prove that $\mathsf{CubicEqSat}$ is $\Sigma^{0}_{1}$-complete.
 Equivalently, satisfiability of single cubic equations over $\mathbb{N}$ with instance-dependent arity is undecidable.
@@
   1. Layer 1.
-    Compilation. We construct a uniform primitive recursive compiler
-    $\varphi\longmapsto\mathcal{S}_{\varphi}$ (1.28)
+    We construct a uniform primitive recursive compiler
+    $\varphi\longmapsto\mathcal{S}_{\varphi}$ (29)
     mapping $\mathsf{RA}$-sentence codes to finite systems of Diophantine equations over $\mathbb{N}$, each of degree $\leq 3$, such that
-    $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\;\Longleftrightarrow\;\mathsf{CubicSat}(\ulcorner\mathcal{S}_{\varphi}\urcorner),$ (1.29)
-    where $\mathsf{CubicSat}$ denotes satisfiability of a single degree-$\leq 3$ equation coinciding with “theoremhood” inside a Hilbert–Ackermann System.
+    $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\;\Longleftrightarrow\;\mathsf{CubicSat}(\ulcorner\mathcal{S}_{\varphi}\urcorner),$ (30)
+    where $\mathsf{CubicSat}$ denotes satisfiability of a single degree-$\leq 3$ equation coinciding with “theoremhood” inside a Hilbert–Ackermann System [Hilbert and Ackermann, 1938].
   2. Layer 2.
-    Constraining.
     The degree budget is controlled locally: finite proof traces are arithmetized via Carryless Pairing in the Fibonacci basis (Section 3), syntactic and inference checks compile to quadratic constraints (Section 4), and the only cubic terms arise when a linear selector activates a quadratic obligation.
@@
   3. Layer 3.
-    Metatheory.
-By the classic Diagonal argument, any total correct decider for $\mathsf{CubicSat}$ would compose with the compiler to decide $\mathsf{Thm}$, (Section 5) leading to contradiction.
-If $\mathsf{RA}$ could internally certify correctness of such a decider on the compiler image, the provability bridge would force $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, contradicting Gödel’s Second Incompleteness Theorem (Gödel’s 2nd henceforth) [Gödel, 1931; Kleene, 1952; Feferman, 1960; Boolos et al., 2012] (Section 5.3).
-Ansatz 1.9 (Metalogic).
-The development harmonizes two formal methodologies: $\mathsf{RA}$ as our theory, while proof ereasure inside Calculus of Inductive Constructions of Coquand and Huet [1988] acts as a second metalogical verification tool.
-Layer $1$ and $2$ of the argument are fully constructive: all maps are primitive recursive, equivalences are witnessed by explicit transformations, and the entire pipeline is mechanized in Rocq without axioms beyond weak aritmetics.
-The construction remains constructive as far as it can: classical reasoning enters only at Layer $3$, where the Diagonal Lemma produces a fixed-point sentence whose existence relies on the final discharge via Gödel’s 2nd.
-This boundary is not a limitation but a structural feature: the undecidability of cubic satisfiability is inherently in need of a classical impossibility result, and the point at which constructivity is exhausted marks precisely the logical depth of the claim.
-Interestingly, the pipeline does not appeal to MRDP [Matiyasevich, 1993]: the Diophantine representation of proof evaluation is constructed directly, with explicit degree control at every step, rather than inherited from the historic theorem.
-Proposition 1.10 (Numeralwise representability).
+    By the classic Diagonal argument, any total correct decider for $\mathsf{CubicSat}$ would compose with the compiler to decide $\mathsf{Thm}$ (Section 5), leading to contradiction.
+If $\mathsf{RA}$ could internally certify correctness of such a decider on the compiler image, the provability bridge would imply
+    $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA}),$ (31)
+    contradicting Gödel’s Second Incompleteness Theorem (Gödel’s 2nd henceforth) [Gödel, 1931; Kleene, 1952; Feferman, 1960; Boolos et al., 2012] (Section 5.3).
+Ansatz 1.3.2 (Automated type checking as metalogic).
+The development uses $\mathsf{RA}$ as object theory and proof erasure specified by the Calculus of Inductive Constructions [Coquand and Huet, 1988] as metalogical verification.
+Layers $1$ and $2$ are fully constructive: all maps are primitive recursive, all equivalences explicitly witnessed, and the pipeline is mechanized in Rocq without axioms beyond weak arithmetics.
+Classical reasoning enters only at Layer $3$, where the Diagonal Lemma produces a fixed-point sentence discharged via Gödel’s 2nd.
+The pipeline does not appeal to MRDP [Matiyasevich, 1993]: the Diophantine representation of proof evaluation is constructed directly, with explicit degree control at every step.
+Proposition 1.3.3 (Numeralwise representability).
 For every primitive recursive function $f:\mathbb{N}^{k}\to\mathbb{N}$ and relation $R\subseteq\mathbb{N}^{k}$ there is an $\mathcal{L}_{\mathsf{RA}}$-formula $\varphi_{f}(\vec{x},y)$ such that for all numerals $\overline{\vec{n}},\overline{m}$:
-$f(\vec{n})=m$ $\;\Longrightarrow\;\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m}),$ (1.30)
-$\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m})\wedge\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m^{\prime}})$ $\;\Longrightarrow\;m=m^{\prime}.$ (1.31)
+$f(\vec{n})=m\;\Longrightarrow\;\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m}),$ (32)
+$\mathsf{RA}\vdash\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m})\wedge\varphi_{f}(\overline{n}_{1},\ldots,\overline{n}_{k},\,\overline{m^{\prime}})\;\Longrightarrow\;m=m^{\prime}.$ (33)
 A similar clause holds for relations $R$ with a witnessing formula $\varphi_{R}(\vec{x})$. This numeralwise representability suffices to arithmetize the compiler and checker.
-Proposition 1.11 (Instancewise checker–constraint equivalence).
+Proposition 1.3.4 (Instancewise checker–constraint equivalence).
 The map $\varphi\mapsto\mathcal{S}_{\varphi}$ is primitive recursive. Moreover, for each specific pair of numerals $(\bar{p},\bar{t})$,
-$\mathrm{check}(\bar{p},\bar{t})\;\Rightarrow\;\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x}),$ (1.32)
+$\mathrm{check}(\bar{p},\bar{t})\;\Rightarrow\;\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x}),$ (34)
 and conversely,
-$\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x})\;\Rightarrow\;\mathrm{check}(\bar{p},\bar{t})\text{ holds in }\mathbb{N}.$ (1.33)
-In particular, for each standard proof code $p$ of $\varphi$, $\mathsf{RA}$ proves the existence of a satisfying assignment for the emitted constraint system.
+$\mathsf{RA}\vdash\exists\vec{x}\;\mathtt{all\_zero}(\mathcal{S}_{\bar{t}},\bar{p},\vec{x})\;\Rightarrow\;\mathrm{check}(\bar{p},\bar{t})\text{ holds in }\mathbb{N}.$ (35)
+In particular, for each standard proof code $p$ of $\varphi$, $\mathsf{RA}$ proves the existence of a satisfying assignment for the emitted constraint system (Table 1).
 By construction.
@@
 We use these facts only for numeralwise representability: for each concrete tuple of numerals $\overline{\vec{n}}$, $\mathsf{RA}$ proves the corresponding instance of $\varphi_{f}$.
-We do not assert that $\mathsf{RA}$ proves $\forall\vec{x}\,\exists!y\,\varphi_{f}(\vec{x},y)$ for arbitrary primitive recursive $f$.
-∎
+We do not assert that $\mathsf{RA}$ proves
+$\forall\vec{x}\,\exists!y\,\varphi_{f}(\vec{x},y)$ (36)
+for arbitrary primitive recursive $f$. ∎
 2 Constructive Environment
-2.1 Mechanization
-Ansatz 2.1.
+2.1 Mechanization in Rocq
+Mechanization 2.1.1 (Proof environment T002).
 This section establishes the computational infrastructure, coding of syntax, the proof predicate with derivability conditions, and some prerequisites of our fixed point.
-The pipeline described in the Ansatz 1.9 is implemented in Gallina, the specification language of
-By Curry–Howard [Sørensen and Urzyczyn, 2006], the implicational fragment with $\bot$ admits a “simply-typed” $\lambda$-definable interpretation: each axiom (K, S, EFQ) realizes as a closed term and Modus ponens corresponds to application.
+The pipeline described in the Ansatz 1.3.2 is implemented in Gallina, the specification language of
+By Curry–Howard Correspondence [Sørensen and Urzyczyn, 2006], the implicational fragment with $\bot$ admits a “simply-typed” $\lambda$-definable interpretation: each axiom (K, S, EFQ) realizes as a closed term and modus ponens corresponds to application.
 The development formalizes the Hilbert-style proof evaluator in $\mathtt{Kernel\_API}$, $\mathtt{R13}$ (Section 4.1), compiles proof traces to cubic Diophantine constraints (Section 4), and extracts a certified boolean kernel,
-$\mathtt{cubic\_accepts}:\mathbb{N}\longrightarrow\mathbb{B},$ (2.34)
+$\mathtt{cubic\_accepts}:\mathbb{N}\longrightarrow\mathbb{B},$ (37)
 whose acceptance set coincides with the acceptance predicate of the proof evaluator. The computational content extracts to a standalone OCaml artifact.
@@
 2.2 Repository Structure and Trust Base
-Ansatz 2.2 (Compilation layers).
+Mechanization 2.2.1 (Reduction).
 The reduction $\varphi\mapsto\mathcal{S}_{\varphi}$ is organized in four formally separated layers, each corresponding to a segment of the $\mathtt{R\_SigmaReduction}$ module chain:
@@
   2. Module 2.
-    Verification and constraint compilation ($\mathtt{R04}$$\mathtt{R11}$).
-Hilbert formula syntax and the executable proof checker ($\mathtt{R04}$$\mathtt{R05}$), formula and sequence coding via Carryless Pairing ($\mathtt{R06}$), and constraint assembly ($\mathtt{R08}$$\mathtt{R11}$) mapping proof-checking conditions to polynomial equations.
+    Verification and constraint compilation ($\mathtt{R04}$$\mathtt{R06}$, $\mathtt{R08}$$\mathtt{R11}$).
+Hilbert-formula syntax and the executable proof checker ($\mathtt{R04}$$\mathtt{R05}$), formula and sequence coding via Carryless Pairing ($\mathtt{R06}$), and constraint assembly ($\mathtt{R08}$$\mathtt{R11}$) mapping proof-checking conditions to polynomial equations.
 Axiom and target constraints have degree $\leq 2$; cubic terms arise only when degree-1 selector variables activate quadratic obligations, keeping every emitted constraint at degree $\leq 3$.
@@
   4. Module 4.
-    Reduction endpoint and proof theory ($\mathtt{R14}$$\mathtt{R19}$ and $\mathtt{P00}$$\mathtt{P06}$).
+    Reduction endpoint and proof theory ($\mathtt{R14}$$\mathtt{R19}$, with auxiliary $\mathtt{R20}$$\mathtt{R22}$, and $\mathtt{P00}$$\mathtt{P06}$).
 The reduction core and code bridge ($\mathtt{R14}$$\mathtt{R15}$), the reduction nucleus $\mathtt{compiler\_primitive}$ with coded evaluator semantics ($\mathtt{R16}$), OCaml extraction ($\mathtt{R17}$), and stable API boundaries ($\mathtt{R18}$$\mathtt{R19}$).
+Auxiliary modules cover aggregated evaluation and channel/pairing support ($\mathtt{R20}$$\mathtt{R22}$).
 The proof-theory layer ($\mathtt{P00}$$\mathtt{P06}$) mechanizes concrete provability ($\mathtt{P00}$), HBL derivability conditions ($\mathtt{P01}$), the fixed-point construction ($\mathtt{P02}$), the premise-free coded undecidability theorem ($\mathtt{P03}$), the RA-certification corollary ($\mathtt{P04}$), and the toggle-witness contradiction core ($\mathtt{P05}$$\mathtt{P06}$).
-Ansatz 2.3.
+Mechanization 2.2.2 (Extraction).
 Rocq extraction ($\mathtt{R17}$) emits two self-contained OCaml modules, each free of standard-library dependencies:
@@
     cubic_checker.ml exports:
-    $\mathtt{cubic\_accepts}:\mathsf{Proof}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\},$ (2.35)
+    $\mathtt{cubic\_accepts}:\mathsf{Proof}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\},$ (38)
     the proof-checking kernel.
-Internally the module re-derives its own bool, nat, list, and form types and contains the full verification pipeline: structural formula equality (form_eqb), axiom-schema recognizers (is_K, is_S, is_EFQ), bounded Modus ponens witness search (mp_witness), and line-by-line validation (check_lines).
+Internally the module re-derives its own bool, nat, list, and form types and contains the full verification pipeline: structural formula equality (form_eqb), logical axiom-schema recognizers (is_K, is_S, is_EFQ), non-logical $\mathsf{RA}$ axiom-schema recognizers (successor axioms, $\Delta_{0}$-induction instances, $\mathrm{B}\Sigma_{1}$-collection instances), bounded modus ponens witness search (mp_witness), and line-by-line validation (check_lines).
 The entry point $\mathtt{cubic\_accepts}$ is definitionally equal to $\mathtt{check}$; the kernel API ($\mathtt{R13}$) re-exports the checker under a reduction-facing alias.
   2. (ii)
-    cubic_compile.ml exports a (legacy) polynomial,
-    $\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial})$ (2.36)
+    cubic_compiler.ml exports a (legacy) polynomial,
+    $\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial})$ (39)
     the single equation,
-    $\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{Polynomial}\times\mathsf{Polynomial},$ (2.37)
+    $\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{Polynomial}\times\mathsf{Polynomial},$ (40)
     together with the symbolic expression type,
-    $\mathtt{Expr}(\mathtt{Const}\mid\mathtt{Var}\mid\mathtt{Add}\mid\mathtt{Sub}\mid\mathtt{Mult})$ (2.38)
+    $\mathtt{Expr}(\mathtt{Const}\mid\mathtt{Var}\mid\mathtt{Add}\mid\mathtt{Sub}\mid\mathtt{Mult})$ (41)
     and the structural-recursion function,
-    $\mathtt{degree}:\mathtt{Expr}\to\mathbb{N}$ (2.39)
-    Given a proof and target, $\mathtt{emit\_cubic\_system}$ produces polynomial constraints, while the the endpoint $\mathtt{emit\_single\_cubic}$ applies two-channel aggregation and returns a single cubic equality surface $(P,Q)$ representing $P=Q$.
-Mapping $\mathtt{degree}\circ\mathtt{poly\_expr}$ over the emitted expressions empirically verifies the degree-$3$ bound of Proposition 4.28.
-The module additionally includes the constraint types ($\mathtt{C\_Axiom}$, $\mathtt{C\_MP}$, $\mathtt{C\_Justification}$, $\mathtt{C\_Target}$), the canonical assignment builder, and the Carryless Pairing and carryless sequence-indexing infrastructure required by the compilation pipeline.
+    $\mathtt{degree}:\mathtt{Expr}\to\mathbb{N}$ (42)
+    Given a proof and target, $\mathtt{emit\_cubic\_system}$ produces polynomial constraints, while the endpoint $\mathtt{emit\_single\_cubic}$ applies two-channel aggregation and returns a single cubic equality surface $(P,Q)$ representing $P=Q$.
+Mapping $\mathtt{degree}\circ\mathtt{poly\_expr}$ over the emitted expressions empirically verifies the degree-$3$ bound of Proposition 4.8.10.
+The module additionally includes the constraint types ($\mathtt{C\_Axiom}$, $\mathtt{C\_MP}$, $\mathtt{C\_Justification}$, $\mathtt{C\_Target}$), the canonical assignment builder, and the Carryless Pairing and sequence-indexing infrastructure required by the compilation pipeline.
+    Surface convention.
+The kernel API exposes the pair surface $(P,Q)$ to avoid introducing subtraction into the internal AST.
+For the canonical Diophantine coding used in all completeness statements (Equation 28), the compiler output is collapsed to a single polynomial equation via PQ=0P-Q=0 (degree unchanged by Lemma 2.2.3).
+The two surfaces are definitionally intertranslatable; see Lemma 2.2.3.
 The correctness of the mechanized development depends on at least three components external to the formalization: (i) the Rocq kernel (the core type checker); (ii) the extraction mechanism; (iii) OCaml compilation.
 No axioms, parameters, or $\mathtt{Admitted}$ obligations are introduced: $\mathtt{PrintAssumptions}$ on the endpoint theorems of both $\mathtt{SigmaReduction}$ and $\mathtt{ProofTheory}$ yields an empty context.
-Intuition.
-A clarification on the diagonal construction is necessary, since it appears at two distinct levels.
-The Rocq development contains a code-level diagonal ($\mathtt{P02}$$\mathtt{P03}$): it operates entirely on natural-number codes, producing a fixed-point code $g$ satisfying $g=\mathtt{apply\_template}\;T\;g$ by arithmetic on encodings, and derives $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$)—the impossibility of a coded decider—constructively, without classical axioms.
-This is a theorem about codes: it quantifies over natural numbers representing programs, not over Turing Machines as external objects.
-The classical meta-theoretic diagonal of Section 5.2 (Theorem 5.2) is a separate argument: it applies the Diagonal Lemma to arbitrary Turing Machines with case analysis via the Law of the Excluded Middle, and is not Rocq-verified.
-Likewise, Gödel’s 2nd as applied to $\mathsf{RA}$ (Corollary 5.7) relies on the arithmetization of Section 3.4 and standard classical logic, and is presented as an ordinary mathematical argument.
-A machine-readable cross-reference map of manuscript claims to Rocq theorems is provided in $\mathtt{TechnicalSummary.yaml}$.
-Remark.
+Lemma 2.2.3 (Interface equivalence).
+There are primitive recursive, degree-preserving translations
+$\mathsf{toZero}(P,Q)$ $:=P-Q,$ (43)
+$\mathsf{toPair}(R)$ $:=(R,\,0),$ (44)
+such that for all assignments $\rho$ over $\mathbb{N}$:
+$P(\rho)=Q(\rho)\;\Longleftrightarrow\;\mathsf{toZero}(P,Q)(\rho)=0,$ (45)
+$R(\rho)=0\;\Longleftrightarrow\;\mathsf{toPair}(R)(\rho)=(R(\rho),\,0).$ (46)
+In particular, both encodings yield the same satisfiability predicate up to a primitive recursive many-one equivalence:
+$\deg(P-Q)\leq\max(\deg P,\deg Q)$ (47)
+whenever $P\neq Q$, and
+$\deg R=\max(\deg R,0).$ (48)
+The Rocq artifact pins the pair surface $(P,Q)$ because it is convenient for normalization and avoids subtraction in the internal AST; the proof-of-concept and all manuscript theorems use the $R=0$ surface (Equation 28).
+Immediate. $P=Q$ iff PQ=0P-Q=0, and $R=0$ iff $(R,0)$ is a pair with equal components. Both translations are compositional, primitive recursive in the polynomial codes, and preserve total degree. ∎
+Mechanization 2.2.4 (Fixed points).
+The diagonal construction appears at two distinct levels:
+  1. (i)
+    Code-level ($\mathtt{P02}$$\mathtt{P03}$).
+Operates on natural-number codes, producing a fixed-point code $g$ with $g=\mathtt{apply\_template}\;T\;g$, and derives $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) constructively, without classical axioms.
+  2. (ii)
+    Meta-theoretic (Section 5.2, Theorem 5.2.1).
+Applies the Diagonal Lemma to arbitrary Turing Machines via the Law of the Excluded Middle; not Rocq-verified.
+Gödel’s 2nd as applied to $\mathsf{RA}$ (Corollary 5.3.5) likewise uses standard classical logic.
+A cross-reference map of manuscript claims to Rocq theorems is provided in $\mathtt{TechnicalSummary.yaml}$.
+Remark 2.2.5.
 The reduction used for $\Sigma^{0}_{1}$-hardness does not fix a proof-length bound.
-The compiler $f(u)$ treats the entire proof witness as a single carryless list code $p$ (together with its per-line selectors), and emits one cubic equation whose free variables include $p$, the target code $u$, and all selector/auxiliary variables.
+The compiler $f(u)$ treats the entire proof witness as a single carryless-list code $p$ (together with its per-line selectors), and emits one cubic equation whose free variables include $p$, the target code $u$, and all selector/auxiliary variables.
 The Diophantine instance is therefore
-$\exists\,p,\,\vec{x}\;\Phi(u,p,\vec{x})=0$
+$\exists\,p,\,\vec{x}\;\Phi(u,p,\vec{x})=0$ (49)
 with $\deg\Phi\leq 3$ and a fixed arity determined solely by the structure of the checker/aggregator—not by a chosen proof length.
@@
 This yields, for each $u$,
-$\operatorname{Prov}_{\mathsf{RA}}(u)\;\Longleftrightarrow\;\exists\,p,\vec{x}\;\Phi(u,p,\vec{x})=0\;\Longleftrightarrow\;\mathsf{CubicEqSat}(f(u)),$
+$\operatorname{Prov}_{\mathsf{RA}}(u)\;\Longleftrightarrow\;\exists\,p,\vec{x}\;\Phi(u,p,\vec{x})=0\;\Longleftrightarrow\;\mathsf{CubicEqSat}(f(u)),$ (50)
 where $f$ is primitive recursive and $\mathsf{CubicEqSat}$ ranges over a single cubic equation with fixed variable budget.
 The banded aggregation uses dynamically sized Fibonacci bands: the width parameter $K$ is computed from the size of the emitted constraint system (number of constraints / channels), not from a fixed global constant.
-Each channel is paired with a capacity constraint of the form $B_{K}-1-A_{i}=\sum_{j=1}^{4}s_{ij}^{2}$, where $B_{K}=F_{K+1}$ and $s_{ij}$ are slack variables.
-Thus any assignment with $A_{i}\geq B_{K}$ is rejected.
-Since every $A_{i},B_{i}$ is a polynomial in the proof/selector variables of degree $\leq 2$, the capacity bound is computable from the instance size, and the added slack variables preserve equivalence: a satisfying assignment to the original system extends to one satisfying the capacity equations, and any violation of the bound is ruled out by the slack equality.
+Each channel is paired with a capacity constraint of the form
+$B_{K}-1-A_{i}=\sum_{j=1}^{4}s_{ij}^{2},$ (51)
+where
+$B_{K}=F_{K+1}\quad\text{and}\quad s_{ij}$ (52)
+are slack variables, hence any satisfying assignment automatically satisfies
+$0\leq A_{i}<B_{K}\quad\text{and}\quad 0\leq B_{i}<B_{K},$ (53)
+an assignment with $A_{i}\geq B_{K}$ is rejected because the slack equation has no nonnegative solution.
+Since every $A_{i},B_{i}$ is a polynomial in the proof/selector variables of degree $2$, the capacity bound is computable from the instance size, and the added slack variables preserve equivalence: a satisfying assignment to the original system extends to one satisfying the capacity equations, and any violation of the bound is ruled out by the slack equality.
+No external representation bound is needed; the base is determined syntactically and the digit bounds are enforced internally (Lemma 4.8.5).
 3 Arithmetic Infrastructure
 3.1 Fibonacci Preliminaries
-Intuition.
+Remark 3.1.1.
 The encoding of finite proof traces as Diophantine witnesses requires an injective, primitive recursive pairing function over $\mathbb{N}$.
@@
 This return to Fibonacci arithmetic is not incidental; it follows a historical line of development in work on Hilbert’s Tenth Problem by Chudnovsky, Jones, Matiyasevich, and others, see Jones [1975]; Matiyasevich [1993].
-The resulting “carryless” coding discipline and band separation refine this tradition.
+The resulting “carryless” discipline and band separation refine this tradition.
 The additive Fibonacci recurrence yields a canonical representation with explicit inversion, while avoiding interference between adjacent components.
 This separation permits controlled aggregation of constraints, and the Diophantine degree cost remains quadratic throughout the encoding layer.
-See Rosko [2026] for a self-contained treatment of the construction and its formal properties.
-Definition 3.1 (Fibonacci canonicalization).
+Definition 3.1.2 (Fibonacci canonicalization).
 The Fibonacci sequence is defined here recursively by
-$F_{0}=0,\quad F_{1}=1,\quad F_{n+2}=F_{n+1}+F_{n}\quad\text{for all }n\geq 0.$ (3.40)
-Lemma 3.2 (Fibonacci monotonicity).
+$F_{0}=0,\quad F_{1}=1,\quad F_{n+2}=F_{n+1}+F_{n}\quad\text{for all }n\geq 0.$ (54)
+Lemma 3.1.3 (Fibonacci monotonicity).
 For all $n\geq 1$,
-$F_{n+1}\geq F_{n},\quad F_{n+2}>F_{n}.$ (3.41)
-Induction coincides with the recursive definition. ∎
-Theorem 3.3 (Zeckendorf’s Theorem).
+$F_{n+1}\geq F_{n},\quad F_{n+2}>F_{n}.$ (55)
+Immediate. Induction coincides with the recursive definition. ∎
+Theorem 3.1.4 (Zeckendorf’s Theorem).
 Every natural number $n>0$ admits a unique representation
-$n=\sum_{k\in S}F_{k},\quad S\subseteq\{2,3,4,\ldots\}$ (3.42)
+$n=\sum_{k\in S}F_{k},\quad S\subseteq\{2,3,4,\ldots\}$ (56)
 where (i) $S$ is a finite set of indices and (ii) no two consecutive integers belong to $S$. We call $S$ the Zeckendorf support of $n$ and write $Z(n):=S$.
-By Zeckendorf [1972].
+By induction [Zeckendorf, 1972].
 The non-adjacency constraint is preserved by the monotonic gap property $F_{k+2}>F_{k}$.
@@
 The Fibonacci identity
-$\sum_{i=2}^{k_{\max}-2}F_{i}=F_{k_{\max}}-1$ (3.43)
+$\sum_{i=2}^{k_{\max}-2}F_{i}=F_{k_{\max}}-2$ (57)
 implies that any combination of non-consecutive indices below $k_{\max}$ sums to strictly less than $F_{k_{\max}}$, yielding a contradiction. ∎
 3.2 Carryless Pairing
-Intuition.
+Remark 3.2.1.
 The Zeckendorf Representation assigns to each $n$ a canonical support set $Z(n)\subseteq\mathbb{N}$.
@@
 No arithmetic carries occur and composition is achieved by set-theoretic union of supports.
-For a presentation with consult either Rosko [2026] or the repository:
+For a presentation, consult either Rosko [2026] or the repository:
 Repository: github.com/…/Proofcase/…/T001/
-Theorem 3.4 (Carryless Pairing).
-Define $Z(0):=\varnothing$. For $x,y\in\mathbb{N}$, set
-$E(x)$ :={ $:=\{\,2k:k\in Z(x)\,\},$ (3.44)
-$r(x)$ $:=\min\{\,m\geq 2:F_{m}>x\,\},$ (3.45)
-$B(x)$ $:=2r(x),$ (3.46)
-$O(x,y)$ $:=\{\,B(x)+(2j-1):j\in Z(y)\,\}.$ (3.47)
-Define the pairing
-$\mathrm{pair}(x,y)\;:=\;F_{2}\;+\;\sum_{i\in E(x)\cup O(x,y)}F_{i}.$ (3.48)
+Theorem 3.2.2 (Carryless Pairing).
+Over $Z(0):=\varnothing$ and $x,y\in\mathbb{N}$, set
+$E(x)$ :={ $:=\{\,2k:k\in Z(x)\,\},$ (58)
+$r(x)$ $:=\min\{\,m\geq 2:F_{m}>x\,\},$ (59)
+$B(x)$ $:=2r(x),$ (60)
+$O(x,y)$ $:=\{\,B(x)+(2j-1):j\in Z(y)\,\}.$ (61)
+Define a pairing function,
+$\mathrm{pair}(x,y)\;:=\;F_{2}\;+\;\!\!\!\sum_{i\in E(x)\cup O(x,y)}\!\!\!F_{i}.$ (62)
 Then $\mathrm{pair}(x,y)\geq 1$ for all $x,y$, the sets $E(x)$ and $O(x,y)$ are disjoint with
-$\max E(x)<B(x)\leq\min O(x,y)-3,$ (3.49)
-and $\mathrm{pair}$ is injective. Moreover there are primitive recursive projections $\mathrm{hd},\mathrm{tl}$ such that
-$u>0\;\to\;\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u,\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0,$ (3.50)
-and $\mathrm{unpair}(u)=(\mathrm{hd}(u),\mathrm{tl}(u))$ is the inverse of $\mathrm{pair}$ on $\mathbb{N}^{2}$.
-By induction over uniquness, construction of $B(x)=2r(x)$, and rank definition.
-Band separation ensures the union $E(x)\cup O(x,y)$ is a valid Zeckendorf support.
-By uniqueness (Theorem 3.3), the support is recoverable; the even and odd bands are then separated and decoded independently.
-∎
-Definition 3.5 (List encoding).
+$\max E(x)<B(x)\leq\min O(x,y)-3,$ (63)
+and $\mathrm{pair}(\cdot)$ is injective. Moreover there are primitive recursive projections $\mathrm{hd},\mathrm{tl}$ such that
+$u>0\;\to\;\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u,\quad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0,$ (64)
+and
+$\mathrm{unpair}(u)=(\mathrm{hd}(u),\mathrm{tl}(u))$ (65)
+is the inverse of $\mathrm{pair}$ on $\mathbb{N}^{2}$.
+By construction. Band separation, $B(x)=2r(x),$ ensures the union
+$E(x)\cup O(x,y)$ (66)
+is a valid Zeckendorf support. By uniqueness (Theorem 3.1.4), the support is recoverable; the even and odd bands are then separated and decoded independently. ∎
+Definition 3.2.3 (List encoding).
 Define recursively:
-$\mathrm{code}_{\mathrm{list}}([])$ $:=0,$ (3.51)
-$\mathrm{code}_{\mathrm{list}}(x_{1},\ldots,x_{n})$ $:=\mathrm{pair}(x_{1},\mathrm{code}_{\mathrm{list}}(x_{2},\ldots,x_{n})).$ (3.52)
-The function $\mathrm{code}_{\mathrm{list}}$ inherits injectivity from Theorem 3.4.
-Remark.
+$\mathrm{code}_{\mathrm{list}}([])$ $:=0,$ (67)
+$\mathrm{code}_{\mathrm{list}}(x_{1},\ldots,x_{n})$ $:=\mathrm{pair}(x_{1},\mathrm{code}_{\mathrm{list}}(x_{2},\ldots,x_{n})).$ (68)
+The function $\mathrm{code}_{\mathrm{list}}$ inherits injectivity from Theorem 3.2.2.
+Remark 3.2.4.
 The construction is termed “carryless” because composite codes are formed by set union of disjoint supports rather than additive mixing: decoding is local and requires no interdependent constraints.
@@
 3.3 Carryless Sequence Indexing
-Intuition.
-Finite proof traces are encoded directly by nested Carryless Pairing.
-The same pairing infrastructure used for formula coding also provides line access: the $i$-th entry is obtained by iterating the tail projection $i$ times and reading the head.
+Remark 3.3.1.
+Finite proof traces are encoded directly by Carryless Pairing.
+The same infrastructure used for formula coding also provides line access: the $i$-th entry is obtained by iterating the tail projection $i$ times and reading the head.
 This removes any dependence on modular sequence decoding while keeping extraction primitive recursive and degree-local.
-Definition 3.6 (Head and tail projections).
-Let $\mathrm{pair}:\mathbb{N}^{2}\to\mathbb{N}$ be the carryless pairing of Theorem 3.4. Define primitive recursive projections
-$\mathrm{hd},\mathrm{tl}:\mathbb{N}\to\mathbb{N}$ (3.53)
+Definition 3.3.2 (Head and tail projections).
+Let $\mathrm{pair}:\mathbb{N}^{2}\to\mathbb{N}$ be the carryless pairing of Theorem 3.2.2. Define primitive recursive projections
+$\mathrm{hd},\mathrm{tl}:\mathbb{N}\to\mathbb{N}$ (69)
 by
-$\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u\;\;(u>0),\qquad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0.$ (3.54)
-Definition 3.7 (Indexed extractor).
+$\mathrm{pair}(\mathrm{hd}(u),\mathrm{tl}(u))=u\;\;(u>0),\quad\mathrm{hd}(0)=0,\;\mathrm{tl}(0)=0.$ (70)
+Definition 3.3.3 (Indexed extractor).
 Define by primitive recursion:
-$\mathrm{tail}^{0}(s)$ $:=s,$ (3.55)
-$\mathrm{tail}^{i+1}(s)$ $:=\mathrm{tl}(\mathrm{tail}^{i}(s)),$ (3.56)
-$\mathrm{line}(s,i)$ $:=\mathrm{hd}(\mathrm{tail}^{i}(s)).$ (3.57)
-Theorem 3.8 (Carryless sequence representation).
+$\mathrm{tail}^{0}(s)$ $:=s,$ (71)
+$\mathrm{tail}^{i+1}(s)$ $:=\mathrm{tl}(\mathrm{tail}^{i}(s)),$ (72)
+$\mathrm{line}(s,i)$ $:=\mathrm{hd}(\mathrm{tail}^{i}(s)).$ (73)
+Theorem 3.3.4 (Carryless sequence representation).
 For every finite sequence $a=(a_{0},\ldots,a_{n-1})$, there exists $s\in\mathbb{N}$ such that
-$\mathrm{line}(s,i)=a_{i}\quad\text{for all }i<n.$ (3.58)
-Take $s:=\mathrm{code}_{\mathrm{list}}(a_{0},\ldots,a_{n-1})$ from Definition 3.5. The claim follows by induction on $i$, using invertibility of $\mathrm{pair}$ (Theorem 3.4). ∎
-Definition 3.9 (Bounded extraction witnesses).
+$\mathrm{line}(s,i)=a_{i}\quad\text{for all }i<n.$ (74)
+By induction over $s:=\mathrm{code}_{\mathrm{list}}(a_{0},\ldots,a_{n-1})$ from Definition 3.2.3. The claim follows by induction on $i$, using invertibility of $\mathrm{pair}$ (Theorem 3.2.2). ∎
+Definition 3.3.5 (Bounded extraction witnesses).
 For fixed index $i$, introduce witness variables
-$t_{i,0},\ldots,t_{i,i+1},\quad h_{i,0},\ldots,h_{i,i}.$ (3.59)
+$t_{i,0},\ldots,t_{i,i+1},\quad h_{i,0},\ldots,h_{i,i}.$ (75)
 Define the chain by
-$t_{i,0}$ $=s,$ (3.60)
-$t_{i,r}$ $=\mathrm{pair}(h_{i,r},t_{i,r+1})\quad(0\leq r\leq i).$ (3.61)
+$t_{i,0}$ $=s,$ (76)
+$t_{i,r}$ $=\mathrm{pair}(h_{i,r},t_{i,r+1})\quad(0\leq r\leq i).$ (77)
 Then $h_{i,i}=\mathrm{line}(s,i)$ and $t_{i,i+1}=\mathrm{tail}^{\,i+1}(s)$.
-Intuition.
+Remark 3.3.6.
 In the sequel, proof sequences are represented by a single carryless list code $s$.
@@
 3.4 Coding of Syntax
-Intuition.
+Remark 3.4.1.
 […]
-Definition 3.10 (Syntactic and proof predicates).
+Definition 3.4.2 (Syntactic and proof predicates).
 Using Gödel Numbering, we define the following predicates on $\mathbb{N}$:
 […]
-$\operatorname{Sent}(u)$ $\;:\equiv\;\operatorname{Form}(u)\;\wedge\;u\text{ codes a sentence},$ (3.63)
-$\operatorname{Proof}_{\mathsf{RA}}(p,u)$ $\;:\equiv\;p\text{ codes a valid $\mathsf{RA}$-proof coded by }u,$ (3.64)
-$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;:\equiv\;\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u).$ (3.65)
-Observe how $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ holds when $p$ codes a finite sequence such that $\varphi_{n-1}$ is the sentence coded by $u$, and for each $i<n$, either $\varphi_{i}$ is an instance of an axiom schema of $\mathsf{RA}$, or there exist $j,k<i$ with $\varphi_{k}\equiv\varphi_{j}\to\varphi_{i}$ by Modus ponens.
-Lemma 3.11 (Primitive recursiveness).
-The predicates $\operatorname{Form}$, $\operatorname{Sent}$, and $\operatorname{Proof}_{\mathsf{RA}}$ are primitive recursive, consequently, $\operatorname{Prov}_{\mathsf{RA}}$ is $\Sigma_{1}$.
-By definition.
+$\operatorname{Sent}(u)$ $\;:\equiv\;\operatorname{Form}(u)\;\wedge\;u\text{ codes a sentence},$ (79)
+[…]
+$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;:\equiv\;\exists p\;\operatorname{Proof}_{\mathsf{RA}}(p,u).$ (81)
+Observe how $\operatorname{Proof}_{\mathsf{RA}}(p,u)$ holds when $p$ codes a finite sequence such that $\varphi_{n-1}$ is the sentence coded by $u$, and for each $i<n$, either $\varphi_{i}$ is an instance of an axiom schema of $\mathsf{RA}$, or there exist $j,k<i$ with
+$\varphi_{k}\equiv\varphi_{j}\to\varphi_{i}$ (82)
+by modus ponens.
+Lemma 3.4.3 (Primitive recursiveness).
+The predicates $\operatorname{Form}$, $\operatorname{Sent}$, and $\operatorname{Proof}_{\mathsf{RA}}$ are primitive recursive; consequently, $\operatorname{Prov}_{\mathsf{RA}}$ is $\Sigma_{1}$.
+Immediate.
 Each predicate is decided by bounded inspection of the input code.
@@
 ∎
-Definition 3.12 (Bridge translation).
-Let $\operatorname{tr}:\mathcal{L}_{\mathsf{RA}}\to\mathcal{L}_{\bot,\to}$ be the primitive recursive encoding used by the kernel checker $\mathrm{check}_{\to}(p,u)$ on the implicational language.
-The bridge is captured by two numeralwise implications:
-$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;\Longrightarrow\;\mathsf{RA}\vdash\exists p\,\mathrm{check}_{\to}(p,\operatorname{tr}(\bar{u})),$ (3.66)
+Definition 3.4.4 (Bridge translation).
+Let $\operatorname{tr}:\mathcal{L}_{\mathsf{RA}}\to\mathsf{Form}$ be the primitive recursive encoding of $\mathcal{L}_{\mathsf{RA}}$-sentences as elements of $\mathsf{Form}$ (Definition 4.1.2).
+Since the kernel checker $\mathrm{check}(p,u)$ validates proofs in the Hilbert system of Definition 4.1.3—which includes both the logical schemas K, S, EFQ and the non-logical axioms of $\mathsf{RA}$—its acceptance set coincides with $\mathsf{RA}$-provability on $\operatorname{tr}$-images.
+Concretely, the bridge is captured by two numeralwise implications:
+$\operatorname{Prov}_{\mathsf{RA}}(u)$ $\;\Longrightarrow\;\mathsf{RA}\vdash\exists p\,\mathrm{check}(p,\operatorname{tr}(\bar{u})),$ (83)
 […]
-That is, provable sentences have kernel proofs (formalizable in $\mathsf{RA}$), and any kernel proof witnesses truth of the corresponding $\mathsf{RA}$ sentence in the standard model.
 No global biconditional inside $\mathsf{RA}$ is assumed.
 3.5 Fixed-Point Construction
-Theorem 3.13 (Derivability Conditions).
+Theorem 3.5.1 (Derivability Conditions).
 The provability predicate $\operatorname{Prov}_{\mathsf{RA}}$ satisfies the Hilbert–Bernays–Löb derivability conditions. For all sentences $\varphi,\psi$:
-$\textsf{D1}:\;$ $\text{If}\;\mathsf{RA}\vdash\varphi,\;\text{then}\;\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner).$ (3.68)
-$\textsf{D2}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\urcorner\bigr).$ (3.69)
-$\textsf{D3}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\to\psi\urcorner)\to\bigl(\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\psi\urcorner)\bigr).$ (3.70)
-These conditions are required for Gödel’s 2nd as applied in Corollary 5.7.
-$\mathsf{D1}$ is immediate: a proof of $\varphi$ in $\mathsf{RA}$ is a concrete finite sequence whose code witnesses $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)$.
-(D2) and (D3) are standard consequences of the primitive recursiveness of $\operatorname{Proof}_{\mathsf{RA}}$ (Lemma 3.11) and the representability of primitive recursive functions in $\mathsf{RA}$ (Proposition 1.10): $\mathsf{RA}$ can internally simulate concatenation of proof sequences and verify that the result is again a valid proof.
+$\textsf{D1}:\;$ $\text{If}\;\mathsf{RA}\vdash\varphi,\;\text{then}\;\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner).$ (85)
+$\textsf{D2}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\urcorner\bigr).$ (86)
+$\textsf{D3}:\;$ $\mathsf{RA}\vdash\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\to\psi\urcorner)\to\bigl(\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)\to\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\psi\urcorner)\bigr).$ (87)
+These conditions are required for Gödel’s 2nd as applied in Corollary 5.3.5.
+Immediate.
+$\mathsf{D1}$: a proof of $\varphi$ in $\mathsf{RA}$ is a concrete finite sequence whose code witnesses $\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\varphi\urcorner)$.
+(D2) and (D3) are standard consequences of the primitive recursiveness of $\operatorname{Proof}_{\mathsf{RA}}$ (Lemma 3.4.3) and the representability of primitive recursive functions in $\mathsf{RA}$ (Proposition 1.3.3): $\mathsf{RA}$ can internally simulate concatenation of proof sequences and verify that the result is again a valid proof.
 See Boolos et al.
@@
 ∎
-Theorem 3.14 (Diagonal Lemma).
+Theorem 3.5.2 (Diagonal Lemma).
 For every $\mathcal{L}_{\mathsf{RA}}$-formula $\varphi(x)$ with exactly one free variable $x$, there exists a sentence $u$ such that
-$\mathsf{RA}\;\vdash\;u\;\leftrightarrow\;\varphi(\ulcorner u\urcorner).$ (3.71)
+$\mathsf{RA}\;\vdash\;u\;\leftrightarrow\;\varphi(\ulcorner u\urcorner).$ (88)
+By construction.
 Define the substitution function $\mathrm{sub}(e,n)$ to be the code of the sentence obtained by substituting the numeral $\overline{n}=S^{n}(0)$ for the free variable in the formula coded by $e$.
-The function $\mathrm{sub}$ is primitive recursive, it acts by bounded structural replacement in the formula tree, and is therefore representable in $\mathsf{RA}$ by Proposition 1.10.
+The function $\mathrm{sub}$ is primitive recursive, it acts by bounded structural replacement in the formula tree, and is therefore representable in $\mathsf{RA}$ by Proposition 1.3.3.
 Let
-$\psi(x):\equiv\varphi(\mathrm{sub}(x,x))\quad\text{with}\quad e:=\ulcorner\psi\urcorner$ (3.72)
+$\psi(x):\equiv\varphi\big(\mathrm{sub}(x,x)\big)\quad\text{with}\quad e:=\ulcorner\psi\urcorner$ (89)
 Define $u$ to be the sentence with code $\mathrm{sub}(e,e)$. Then
-$u\;\equiv\;\psi(\overline{e})\;\equiv\;\varphi\bigl(\mathrm{sub}(\overline{e},\overline{e})\bigr)\;\equiv\;\varphi(\ulcorner u\urcorner),$ (3.73)
+$u\;\equiv\;\psi(\overline{e})\;\equiv\;\varphi\big(\mathrm{sub}(\overline{e},\overline{e})\big)\;\equiv\;\varphi(\ulcorner u\urcorner),$ (90)
 and $\mathsf{RA}$ proves this equivalence because $\mathrm{sub}$ is representable. See Rogers [1987]; Boolos et al. [2012]. ∎
-4 From Proof Checking to Polynomial Constraints
+4 From Proof Checking to Constraints
 4.1 The Hilbert–Ackermann Kernel
-Intuition.
-We fix a minimal proof system for propositional logic in the style of Hilbert and Ackermann [1938], with implication and falsity as primitive connectives.
-A proof is a finite sequence of formulas, each either an axiom instance or a consequence of Modus ponens from earlier lines, concluding at the target formula.
-A total boolean checker decides correctness of any proposed proof; soundness (Theorem 4.5) confirms it is faithful to derivability.
-Being primitive recursive, the checker arithmetizes directly via the encoding mechanisms of the preceding sections.
-We then reflect each of its verification conditions into polynomial equalities over $\mathbb{N}$, obtaining a finite constraint system whose solvability is equivalent to checker acceptance (Theorem 4.19) with every emitted equation of degree at most $3$ (Proposition 4.20).
-Definition 4.1 (Formula syntax).
+Remark 4.1.1.
+We fix an automated proof system for $\mathsf{RA}$ in the style of Hilbert and Ackermann [1938].
+Formulas are built from $\mathcal{L}_{\mathsf{RA}}$ atoms, implication ($\to$), and falsity ($\bot$).
+A proof is a finite sequence of formulas, each either a logical axiom instance ($\mathsf{K,S,EFQ}$), an instance of $\mathsf{RA}$ (Definition 1.1.2), or a consequence of modus ponens from earlier lines, concluding at the target.
+A total boolean checker decides correctness; being primitive recursive, it arithmetizes directly via the preceding encoding mechanisms.
+We then reflect each verification condition into polynomial equalities over $\mathbb{N}$, obtaining a constraint system whose solvability is equivalent to checker acceptance (Theorem 4.7.3) with every emitted equation of degree at most $3$ (Proposition 4.7.4).
+Definition 4.1.2 (Formula syntax).
 The set $\mathsf{Form}$ of formulas is defined inductively by:
-$\mathsf{Form}\;::=\;\bot\mid(\varphi\to\psi),$ (4.74)
-where $\bot$ denotes falsity and $\varphi\to\psi$ denotes implication.
-The mechanized kernel ($\mathtt{R04}$) uses exactly these two constructors; propositional variables are not needed, since the axiom schemas already quantify over all formulas.
-Definition 4.2 (Axiom schemas).
-The proof system admits three axiom schemas [Troelstra and Schwichtenberg, 2000].
-The schemas K and S are the logical counterparts of the combinators $\mathbf{K}$ and $\mathbf{S}$.
-For all formulas $\varphi,\psi,\chi\in\mathsf{Form}$:
+$\mathsf{Form}\;::=\;\bot\mid(\varphi\to\psi),$ (91)
+Atomic $\mathcal{L}_{\mathsf{RA}}$-formulas (equalities and their negations) and bounded quantifiers are encoded as nested $\bot/{\to}$ trees via the Gödel Numbering of Section 3.4; the two constructors suffice to represent the full $\mathcal{L}_{\mathsf{RA}}$-formula language.
+The mechanized kernel ($\mathtt{R04}$) uses exactly these two constructors.
+Definition 4.1.3 (Axiom schemas).
+The proof system admits logical and non-logical axiom schemas. The logical schemas K, S, and EFQ [Troelstra and Schwichtenberg, 2000] hold for all formulas $\varphi,\psi,\chi\in\mathsf{Form}$:
   1. (i)
     Schema K:
-    $\varphi\to(\psi\to\varphi).$ (4.75)
+    $\varphi\to(\psi\to\varphi).$ (92)
   2. (ii)
     Schema S:
-    $\big(\varphi\to(\psi\to\chi))\to((\varphi\to\psi)\to(\varphi\to\chi)\big).$ (4.76)
+    $\big(\varphi\to(\psi\to\chi)\big)\to\big((\varphi\to\psi)\to(\varphi\to\chi)\big).$ (93)
   3. (iii)
     Schema EFQ (“ex falso quodlibet”):
-    $\bot\to\varphi.$ (4.77)
-Definition 4.3 (Proof sequence).
+    $\bot\to\varphi.$ (94)
+  4. (iv)
+    Non-logical axioms of $\mathsf{RA}$.
+Each axiom of $\mathsf{RA}$ (Definition 1.1.2)—the successor axioms (2)–(6), instances of the $\Delta_{0}$-induction schema, and instances of the $\mathrm{B}\Sigma_{1}$-collection schema—is accepted as a valid proof line when its $\mathsf{Form}$-encoding matches the corresponding pattern.
+Recognition of each schema instance is primitive recursive.
+Definition 4.1.4 (Proof sequence).
 A proof of a formula $\varphi$ is a finite sequence
-$p=(\varphi_{0},\ldots,\varphi_{n-1})$ (4.78)
+$p=(\varphi_{0},\;\ldots,\;\varphi_{n-1})$ (95)
 such that:
@@
     1. (a)
-      Axiom: $\varphi_{i}$ is an instance of one of the axiom schemas (Definition 4.2).
+      Axiom: $\varphi_{i}$ is an instance of one of the axiom schemas (Definition 4.1.3).
     2. (b)
       Modus ponens: There exist indices $j,k<i$ such that
-      $\varphi_{k}=(\varphi_{j}\to\varphi_{i}).$ (4.79)
+      $\varphi_{k}=(\varphi_{j}\to\varphi_{i}).$ (96)
   2. (ii)
     The final line matches the target:
-    $\varphi_{n-1}=\varphi.$ (4.80)
-Definition 4.4 (Proof checker).
+    $\varphi_{n-1}=\varphi.$ (97)
+Definition 4.1.5 (Proof checker).
 Define a total boolean function
-$\mathrm{check}:\mathsf{ProofSeq}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\}$ (4.81)
-that returns $\mathrm{true}$ if and only if $p$ satisfies the conditions of Definition 4.3 with target $\varphi$.
+$\mathrm{check}:\mathsf{ProofSeq}\times\mathsf{Form}\to\{\mathrm{true},\mathrm{false}\}$ (98)
+that returns $\mathrm{true}$ if and only if $p$ satisfies the conditions of Definition 4.1.4 with target $\varphi$.
 The search for modus ponens witnesses at each line $i$ is bounded by $i$, so $\mathrm{check}$ is total and primitive recursive.
-Theorem 4.5 (Kernel soundness).
+Theorem 4.1.6 (Kernel soundness).
 If $\mathrm{check}(p,\varphi)=\mathrm{true}$, then $\varphi$ is derivable in the Hilbert–Ackermann System, i.e.,
-$\vdash\varphi.$ (4.82)
+$\vdash\varphi.$ (99)
 By induction on the length of the proof sequence $p$.
@@
 4.2 Arithmetization
-Proposition 4.6 (Decidability structure).
+Proposition 4.2.1 (Decidability structure).
 The provability evaluation
-$\mathrm{Provable}(\varphi)\equiv\exists p\;\mathrm{check}(p,\varphi)=\mathrm{true}$ (4.83)
+$\mathrm{Provable}(\varphi)\equiv\exists p\;\mathrm{check}(p,\varphi)=\mathrm{true}$ (100)
 is $\Sigma^{0}_{1}$. When the proof system is instantiated for arithmetic, $\mathrm{Provable}$ is $\Sigma^{0}_{1}$-complete [Rogers, 1987].
-Intuition.
+Remark 4.2.2.
 Formulas are coded as natural numbers via Carryless Pairing (Section 3), and proof sequences are represented by carryless list codes with bounded extraction witnesses (Section 3.3).
@@
 4.3 Arithmetic Representation of Proof Data
-Definition 4.7 (Proof witness assignment).
+Definition 4.3.1 (Proof witness assignment).
 An assignment for a proof of length $n$ targeting formula code $t$ is a tuple
-$a=(s,n,t,j_{0},k_{0},\ldots,j_{n-1},k_{n-1},a_{0},m_{0},\ldots,a_{n-1},m_{n-1},\ldots)$ (4.84)
-where:
-  1. (i)
-    $s\in\mathbb{N}$ is the carryless list code for the proof sequence,
-  2. (ii)
-    $n\in\mathbb{N}$ is the proof length,
-  3. (iii)
-    $t\in\mathbb{N}$ is the target formula code,
-  4. (iv)
-    $j_{i},k_{i}\in\mathbb{N}$ are Modus ponens witness indices for line $i$,
-  5. (v)
-    $a_{i},m_{i}\in\{0,1\}$ are justification selector booleans (axiom vs. Modus ponens),
-  6. (vi)
-    additional auxiliary variables for bounds, line-extraction chains, and intermediate computations.
+$a=(s,n,t,j_{0},k_{0},\ldots,j_{n-1},k_{n-1},a_{0},m_{0},\ldots,a_{n-1},m_{n-1},\ldots)$ (101)
+where (i) $s\in\mathbb{N}$ is the carryless list code for the proof sequence, (ii) $n\in\mathbb{N}$ is the proof length, (iii) $t\in\mathbb{N}$ is the target formula code, (iv) $j_{i},k_{i}\in\mathbb{N}$ are modus ponens witness indices for line $i$, (v) $a_{i},m_{i}\in\{0,1\}$ are justification selector booleans (axiom vs.
+modus ponens), (vi) additional auxiliary variables for bounds, line-extraction chains, and intermediate computations.
 The assignment bundles all numerical witnesses required to verify proof correctness arithmetically.
 4.4 Polynomial Reflection of Basic Tests
-Lemma 4.8 (Equality constraint, degree $2$).
+Lemma 4.4.1 (Equality constraint, degree $2$).
 For $x,y\in\mathbb{N}$, the equality $x=y$ holds if and only if
-$(x-y)^{2}=0.$ (4.85)
-Over $\mathbb{N}$, $(x-y)^{2}=0$ implies xy=0x-y=0, hence $x=y$. Conversely, $x=y$ implies $(x-y)^{2}=0$ immediately. ∎
-Lemma 4.9 (Boolean constraint, degree $2$).
+$(x-y)^{2}=0.$ (102)
+By correctness. Over $\mathbb{N}$, $(x-y)^{2}=0$ implies xy=0x-y=0, hence $x=y$. Conversely, $x=y$ implies $(x-y)^{2}=0$. ∎
+Lemma 4.4.2 (Boolean constraint, degree $2$).
 A variable $b\in\mathbb{N}$ satisfies $b\in\{0,1\}$ if and only if
-b(1b)=0.b(1-b)=0. (4.86)
-If $b=0$ or $b=1$, then b(1b)=0b(1-b)=0. Conversely, b(1b)=0b(1-b)=0 forces either $b=0$ or $b=1$ over $\mathbb{N}$. ∎
-Lemma 4.10 (Conjunction constraint, degree $2$).
+b(1b)=0.b(1-b)=0. (103)
+By correctness. If $b=0$ or $b=1$, then b(1b)=0b(1-b)=0. Conversely, b(1b)=0b(1-b)=0 forces either $b=0$ or $b=1$ over $\mathbb{N}$. ∎
+Lemma 4.4.3 (Conjunction constraint, degree $2$).
 For boolean variables $b_{1},b_{2}\in\{0,1\}$, the conjunction $b=b_{1}\land b_{2}$ is reflected by
-$b=b_{1}b_{2}.$ (4.87)
-Definition 4.11 (Axiom matching constraint family).
+$b=b_{1}b_{2}.$ (104)
+Definition 4.4.4 (Axiom matching constraint family).
 For each line index $i<n$, define a finite family $P_{\mathrm{axiom}}(a,i)$ of polynomial equalities whose simultaneous satisfaction indicates that the formula code $\mathrm{line}(s,i)$ matches one of the axiom schemas.
-The construction proceeds by pattern matching on the structural encoding of formulas: (i) extract subformula codes via line extraction and Carryless “unpairing”; (ii) enforce structural equality constraints using Lemma 4.8; (iii) check that the pattern matches $\mathsf{K}$, $\mathsf{S}$, or $\mathsf{EFQ}$ by verifying the implication nesting structure.
-Proposition 4.12 (Axiom constraint degree).
+The construction proceeds by pattern matching on the structural encoding of formulas: (i) extract subformula codes via line extraction and Carryless-unpairing; (ii) enforce structural equality constraints using Lemma 4.4.1; (iii) check that the pattern matches $\mathsf{K}$, $\mathsf{S}$, $\mathsf{EFQ}$, or an $\mathsf{RA}$ non-logical axiom (successor, $\Delta_{0}$-induction, $\mathrm{B}\Sigma_{1}$) by verifying the implication nesting structure and the appropriate schema parameters.
+Proposition 4.4.5 (Axiom constraint degree).
 Every member of $P_{\mathrm{axiom}}(a,i)$ has degree at most $2$.
-Pattern matching and structural equality tests involve only arithmetic extraction (bounded carryless line extraction, unpairing) and quadratic equality constraints (Lemma 4.8).
+By construction.
+Pattern matching and structural equality tests involve only arithmetic extraction (bounded Carryless-line extraction, unpairing) and quadratic equality constraints (Lemma 4.4.1).
 No cubic terms are introduced.
 ∎
-Definition 4.13 (Modus ponens constraint family).
+Definition 4.4.6 (Modus ponens constraint family).
 For each line index $i<n$, let $\mathcal{P}_{\mathrm{MP}}(a,i)$ be the finite family of polynomial equalities implementing the modus ponens check.
 The family asserts the existence of witness indices $j_{i},k_{i}<i$ (included in the assignment $a$) such that
-$\mathrm{line}(s,k_{i})=\mathrm{enc}_{\mathrm{form}}(\mathrm{line}(s,j_{i})\to\mathrm{line}(s,i)).$ (4.88)
-It comprises: (i) Bound constraints enforcing $j_{i}<i$ and $k_{i}<i$ via four-squares representation; (ii) Structural equality constraints expressing that the implication code at index $k_{i}$ equals the pairing of codes at $j_{i}$ and $i$.
+$\mathrm{line}(s,k_{i})=\mathrm{enc}_{\mathrm{form}}(\mathrm{line}(s,j_{i})\to\mathrm{line}(s,i)).$ (105)
+It comprises: (i) Bound constraints enforcing $j_{i}<i$ and $k_{i}<i$ via Four-squares, (ii) Structural equality constraints expressing that the implication code at index $k_{i}$ equals the pairing of codes at $j_{i}$ and $i$.
 Each member of $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree $\leq 2$.
-Proposition 4.14 (Modus ponens constraint degree).
+Proposition 4.4.7 (Modus ponens constraint degree).
 Every polynomial in $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree at most $2$.
-Bound constraints use quadratic four-square encodings. Structural equality uses quadratic equality tests and the degree-$2$ pairing encoding. No cubic terms appear within any member of the family. ∎
+Immediate.
+Bound constraints use quadratic Four-square encodings.
+Structural equality uses quadratic equality tests and the degree-$2$ pairing encoding.
+No cubic terms appear within any member of the family.
+∎
 4.5 Justification Constraint
-Definition 4.15 (Justification disjunction).
+Definition 4.5.1 (Justification disjunction).
 For each line $i<n$, the justification condition requires that line $i$ is verified either as an axiom or via modus ponens.
 This disjunction is encoded using boolean selector variables $a_{i},m_{i}\in\{0,1\}$ (part of the assignment $a$):
-$a_{i}(1-a_{i})$ $=0,$ (4.89)
-$m_{i}(1-m_{i})$ $=0,$ (4.90)
-$a_{i}+m_{i}-1$ $=0,$ (4.91)
-$a_{i}\cdot Q$ $=0\quad\text{for each }Q\in P_{\mathrm{axiom}}(a,i),$ (4.92)
-$m_{i}\cdot Q$ $=0\quad\text{for each }Q\in\mathcal{P}_{\mathrm{MP}}(a,i).$ (4.93)
-Constraints (4.89)–(4.90) enforce boolean values. Constraint (4.91) enforces exclusive selection. Constraints (4.92)–(4.93) activate the corresponding verification polynomial.
-Proposition 4.16 (Justification constraint degree).
+$a_{i}(1-a_{i})$ $=0,$ (106)
+$m_{i}(1-m_{i})$ $=0,$ (107)
+$a_{i}+m_{i}-1$ $=0,$ (108)
+$a_{i}\cdot Q$ $=0\quad\text{for each }Q\in P_{\mathrm{axiom}}(a,i),$ (109)
+$m_{i}\cdot Q$ $=0\quad\text{for each }Q\in\mathcal{P}_{\mathrm{MP}}(a,i).$ (110)
+Constraints (106)–(107) enforce boolean values. Constraint (108) enforces exclusive selection. Constraints (109)–(110) activate the corresponding verification polynomial.
+Proposition 4.5.2 (Justification constraint degree).
 Every emitted justification constraint has degree at most $3$.
-Constraints (4.89)–(4.91) are degree $\leq 2$.
+By exhaustion.
+Constraints (106)–(108) are degree $\leq 2$.
 Each axiom constraint family member has degree $\leq 2$; activation $a_{i}\cdot Q=0$ raises degree to at most $3$.
 Each MP constraint is drawn from the family $\mathcal{P}_{\mathrm{MP}}(a,i)$ of degree $\leq 2$ polynomials; activation $m_{i}\cdot Q=0$ again yields degree $\leq 3$.
-No product of a selector with a degree-33 polynomial is emitted.
+No product of a selector with a degree-$3$ polynomial is emitted.
 ∎
 4.6 Target Constraint
-Definition 4.17 (Target matching polynomial).
+Definition 4.6.1 (Target matching polynomial).
 The final line of the proof must match the target formula code $t$:
-$(\mathrm{line}(s,n-1)-t)^{2}=0.$ (4.94)
+$(\mathrm{line}(s,n-1)-t)^{2}=0.$ (111)
 This constraint has degree $2$.
-4.7 The Complete Semantic Constraints
-Definition 4.18 (Proof constraint system).
+4.7 Semantic Constraints
+Definition 4.7.1 (Proof constraint system).
 For a proof of length $n$ targeting formula code $t$, let $\mathcal{C}_{n,t}$ be the finite set of polynomial equalities consisting of:
@@
   5. (v)
-    The target constraint (Definition 4.17).
+    The target constraint (Definition 4.6.1).
 No unguarded $P_{\mathrm{axiom}}(a,i)=0$ or $P_{\mathrm{MP}}(a,i)=0$ constraints are included; axiom- and MP-specific auxiliary constraints appear only multiplied by their selectors.
-Remark.
+Remark 4.7.2.
 The system $\mathcal{C}_{n,t}$ is the pre-aggregation constraint list emitted by $\mathtt{emit\_cubic\_system}$.
-It consists of finitely many cubic equations, each of degree at most $3$.
+It consists of finitely many equations, each of degree at most $3$.
 The subsequent aggregation layer (Section 4.8) collapses this finite conjunction to a single cubic equation without increasing degree.
-Theorem 4.19 (Constraint equivalence).
+Theorem 4.7.3 (Constraint equivalence).
 For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$,
-$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists a\in\mathbb{N}^{k}\;\bigwedge_{P\in\mathcal{C}_{n,t}}P(a)=0.$ (4.95)
-By directions.
-  1. Direction 1.
+$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists a\in\mathbb{N}^{k}\;\bigwedge_{P\in\mathcal{C}_{n,t}}P(a)=0.$ (112)
+By correctness.
 Forward.
 Assume $\mathrm{check}(p,\varphi)=\mathrm{true}$.
-Construct an assignment $a$ by (i) Setting $s$ to the carryless list code of $p$ (Theorem 3.8), (ii) instantiating extraction-chain witnesses so that $\mathrm{line}(s,i)$ equals the $i$-th formula code, (iii) for each line $i$: if justified by axiom, set $a_{i}=1,m_{i}=0$; if justified by Modus ponens with witnesses $j,k$, set $a_{i}=0,m_{i}=1,j_{i}=j,k_{i}=k$, and (iv) setting all remaining auxiliary variables to satisfy bound and equality constraints.
+Construct an assignment $a$ by (i) Setting $s$ to the Carryless-list code of $p$ (Theorem 3.3.4), (ii) instantiating extraction-chain witnesses so that $\mathrm{line}(s,i)$ equals the $i$-th formula code, (iii) for each line $i$: if justified by axiom, set $a_{i}=1,m_{i}=0$; if justified by modus ponens with witnesses $j,k$, set $a_{i}=0,m_{i}=1,j_{i}=j,k_{i}=k$, and (iv) setting all remaining auxiliary variables to satisfy bound and equality constraints.
 By construction, all constraints in $\mathcal{C}_{n,t}$ vanish under $a$.
-  2. Direction 2.
 Reverse.
 Assume $a$ satisfies all constraints in $\mathcal{C}_{n,t}$.
-Extract the proof sequence via carryless line decoding: $\varphi_{i}=\mathrm{dec}_{\mathrm{form}}(\mathrm{line}(s,i))$ for $i<n$.
+Extract the proof sequence via Carryless-line extraction: $\varphi_{i}=\mathrm{dec}_{\mathrm{form}}(\mathrm{line}(s,i))$ for $i<n$.
 The justification constraints ensure each line is either an axiom instance or follows by modus ponens.
@@
 Hence $\mathrm{check}(p,\varphi)=\mathrm{true}$.
+∎
 Constraint Expanded equation $\boldsymbol{\delta}$ Mechanism
-Pre-aggregation cubic system $\mathcal{C}_{n,t}$ (no selector activation)
 Line extraction $t_{i,r}-\mathrm{pair}(h_{i,r},t_{i,r+1})=0$ 2 Carryless unpair
@@
 MP structural $M_{i}(\vec{x}):=f_{k_{i}}-\mathrm{pair}(f_{j_{i}},f_{i})=0$ 2 Implication encoding
-Axiom $A_{i}(\vec{x})\cdot M_{i}(\vec{x})=0$ 3 Quadratic $\times$ linear disj.
+Boolean selectors $a_{i}(1-a_{i})=0,\;m_{i}(1-m_{i})=0$ 2 Forces $a_{i},m_{i}\in\{0,1\}$
+Exclusivity $a_{i}+m_{i}-1=0$ 1 Exactly one active
+Axiom activation $a_{i}\cdot A_{i}(\vec{x})=0$ 3 Selector $\times$ quadratic
+MP activation $m_{i}\cdot M_{i}(\vec{x})=0$ 3 Selector $\times$ quadratic
 Target check $f_{n-1}-g_{\mathrm{T}}=0$ 1 Linear equality
-Table 1: Degree bookkeeping for the pre-aggregation cubic constraint system $\mathcal{C}_{n,t}$ after eliminating selector-based activation.
-Axiom and Modus ponens are combined into a single cubic justification equation $A_{i}\cdot M_{i}=0$, where $A_{i}$ is a quadratic axiom-matching obligation and $M_{i}$ is the (linear-in-the-encoding) MP structural equality.
-No boolean selectors and no exactly-one constraints are used.
+Table 1: Degree bookkeeping for the pre-aggregation cubic constraint system $\mathcal{C}_{n,t}$.
+Justification is enforced by selector-activated constraints $a_{i}\cdot A_{i}=0$ and $m_{i}\cdot M_{i}=0$, with exactly-one enforced by $a_{i}+m_{i}=1$.
+Each activation multiplies a degree-$1$ selector by a degree-$\leq 2$ obligation, yielding degree $\leq 3$.
 The subsequent mixed-radix aggregation layer collapses this finite conjunction to a single cubic equation without increasing degree.
-Linear equalities are written as xy=0x-y=0 throughout; squaring is not needed for degree accounting and would only raise degrees.
-Proposition 4.20 (Global degree bound).
+Proposition 4.7.4 (Global degree bound).
 Every polynomial in the constraint system $\mathcal{C}_{n,t}$ has degree at most $3$.
-The mechanization certifies the bound syntactically.
-Each constraint is emitted as an $\mathtt{Expr}$ whose degree is computed by structural recursion in $\mathtt{R00\_\_Degree\_Framework}$.
-Axiom constraints are degree $\leq 2$, MP family members are degree $\leq 2$, and activation multiplies these by linear selectors, keeping every emitted equation at degree $\leq 3$.
-Linear equalities (xy=0x-y=0) are included in both channels to maintain consistent aggregation.
-The assembly theorem
-$\mathtt{polynomial\_system\_forall\_degree\_le\_3}$ (4.96)
-proves that every emitted polynomial has syntactic degree $\leq 3$, independently of any assignment. Table 1 gives the expanded concrete bookkeeping that explains why degree $3$ suffices. ∎
-Lemma 4.21 (AST degree correspondence).
+By exhaustion.
+Each polynomial in $\mathcal{C}_{n,t}$ is obtained from one of the following constructions.
+Linear extraction and readout constraints such as $t_{i,0}-s=0$ and $f_{i}-h_{i,i}=0$ have degree $1$.
+Boolean well-formedness constraints b(1b)=0b(1-b)=0 have degree $2$ (Lemma 4.4.2).
+Equality tests $(x-y)^{2}=0$ have degree $2$ (Lemma 4.4.1).
+Zeckendorf digit and non-adjacency constraints are of the form $d-d^{2}=0,$ and $d_{\kappa}d_{\kappa+1}=0,$ hence degree $2$.
+By Proposition 4.4.5 and 4.4.7, every polynomial in $P_{\mathrm{axiom}}(a,i)$ and $\mathcal{P}_{\mathrm{MP}}(a,i)$ has degree at most $2$.
+Activated obligations are emitted in the form $a_{i}\cdot Q=0,\quad m_{i}\cdot Q=0,$ where $a_{i},m_{i}$ are selector variables and $Q$ is one of the above degree-$\leq 2$ polynomials.
+Since $a_{i}$ and $m_{i}$ are variables of degree $1$, each activated constraint has total degree at most $1+2=3$.
+No constraint multiplies a selector with a polynomial of degree $3$.
+See Table 1.
+The target condition
+$(\mathrm{line}(s,n-1)-t)^{2}=0$ (113)
+has degree $2$ (Definition 4.6.1). These cases exhaust all emitted constraints in $\mathcal{C}_{n,t}$. ∎
+Lemma 4.7.5 (AST degree correspondence).
 For every emitted expression $e$, $\mathtt{degree}(e)$ computed on the AST equals the total polynomial degree of its denotation in the ring $\mathbb{N}[\vec{x}]$.
-By structural induction on $e$: leaves $\mathtt{Const}$ and $\mathtt{Var}$ have degree 0 and $1$ respectively; constructors $\mathtt{Add},\mathtt{Sub}$ take the maximum of subdegrees; $\mathtt{Mult}$ adds subdegrees.
+By induction on the structure of $e$: leaves $\mathtt{Const}$ and $\mathtt{Var}$ have degree 0 and $1$ respectively; constructors $\mathtt{Add},\mathtt{Sub}$ take the maximum of subdegrees; $\mathtt{Mult}$ adds subdegrees.
 This matches the algebraic definition of total degree.
 ∎
-Intuition.
+Remark 4.7.6.
 The reduction from proof checking to cubic solvability is uniform: given a target formula $\varphi$ with code $t$, the compiled single equation $\Phi(t,p,\vec{x})=0$ existentially quantifies over an unbounded proof code $p$; there is no computable proof-length bound.
@@
 4.8 Mixed-Radix Aggregation
-Intuition.
+Remark 4.8.1.
 We collapse the finite list of cubic constraints without squaring by using standard base-$B$ expansion uniqueness.
@@
 Choosing $B:=F_{K}$ (or $F_{K}+1$ for a weak inequality) keeps the Fibonacci aspect but the argument is ordinary mixed radix.
-Lemma 4.22 (Two-channel decomposition).
+Lemma 4.8.2 (Two-channel decomposition).
 Every polynomial $P(\vec{x})\in\mathbb{Z}[\vec{x}]$ is syntactically decomposed as P=ABP=A-B, where $A$ collects the positive-coefficient monomials and $B$ collects the absolute values of the negative-coefficient monomials.
 Over $\mathbb{N}$, both $A(\vec{x}),B(\vec{x})\geq 0$ on every assignment, and
-$P(\vec{x})=0\quad\Longleftrightarrow\quad A(\vec{x})=B(\vec{x}).$ (4.97)
+$P(\vec{x})=0\quad\Longleftrightarrow\quad A(\vec{x})=B(\vec{x}).$ (114)
 Degree does not increase: $\deg A,\deg B\leq\deg P$.
-Definition 4.23 (Capacity constraints).
+Definition 4.8.3 (Capacity constraints).
 For each equality $A_{i}(\vec{x})=B_{i}(\vec{x})$ arising from a constraint $P_{i}=0$, fix an integer base $B\geq 2$ and enforce the digit bounds
-$0\leq A_{i}(\vec{x})<B,\qquad 0\leq B_{i}(\vec{x})<B.$ (4.98)
+$0\leq A_{i}(\vec{x})<B,\quad 0\leq B_{i}(\vec{x})<B.$ (115)
 Encode the inequalities via Four-square slack variables (Lagrange):
-$B-1-A_{i}(\vec{x})=u_{i,1}^{2}+u_{i,2}^{2}+u_{i,3}^{2}+u_{i,4}^{2},$ (4.99)
+$B-1-A_{i}(\vec{x})=u_{i,1}^{2}+u_{i,2}^{2}+u_{i,3}^{2}+u_{i,4}^{2},$ (116)
 and similarly for $B_{i}$.
@@
 In the implementation we set $B:=F_{K}$ (or $F_{K}+1$ for non-strict bounds), but any fixed $B\geq 2$ suffices.
-Lemma 4.24 (Mixed-radix aggregation).
-Let $(A_{i}=B_{i})_{i=1}^{m}$ be equalities of nonnegative polynomials satisfying $0\leq A_{i},B_{i}<B$ for a fixed base $B\geq 2$. Then
-$\sum_{i=1}^{m}B^{i}\,A_{i}=\sum_{i=1}^{m}B^{i}\,B_{i}\quad\Longleftrightarrow\quad\forall i,\;A_{i}=B_{i}.$ (4.100)
-Under the digit bounds, every natural number has a unique representation $\sum_{i=1}^{m}B^{i}d_{i}$ with $d_{i}\in\{0,\ldots,B-1\}$.
-Apply uniqueness with digits $d_{i}:=A_{i}$ and $d_{i}:=B_{i}$ on the two sides.
+Lemma 4.8.4 (Mixed-radix aggregation).
+Let $(A_{i}=B_{i})_{i=1}^{m}$ be equalities of nonnegative polynomials satisfying
+$0\leq A_{i},B_{i}<B$ (117)
+for a fixed base $B\geq 2$. Then
+$\sum_{i=1}^{m}B^{i}\,A_{i}=\sum_{i=1}^{m}B^{i}\,B_{i}\quad\Longleftrightarrow\quad\forall i,\;A_{i}=B_{i}.$ (118)
+By correctness. Under the digit bounds, every natural number has a unique representation
+$\sum_{i=1}^{m}B^{i}d_{i},\quad d_{i}\in\{0,\ldots,B-1\},$ (119)
+so uniqueness with digits $d_{i}:=A_{i}$ and $d_{i}:=B_{i}$ is applied on the two sides. ∎
+Lemma 4.8.5 (Self-bounding aggregation correctness).
+Given a finite list of cubic constraints
+$(L_{i}(\vec{x})=0)_{i=1}^{m}$ (120)
+over $\mathbb{N}$, (Lemma 4.8.2) decompose each
+$L_{i}=A_{i}-B_{i}\quad\text{with}\quad A_{i},B_{i}\geq 0.$ (121)
+Let $B\geq 2$ be a base determined syntactically from the number of channels $m$ (e.g.
+$B:=F_{K+1}$ where $K$ is computed from the constraint count), and include the capacity equalities $C_{1},\ldots,C_{r}$ of Definition 4.8.3 for every channel value $A_{i},B_{i}$, enforcing
+$B-1-A_{i}(\vec{x})=\textstyle\sum_{j=1}^{4}u_{i,j}^{2},\quad B-1-B_{i}(\vec{x})=\textstyle\sum_{j=1}^{4}v_{i,j}^{2},$ (122)
+so that any satisfying assignment automatically satisfies the digit bounds
+$0\leq A_{i},B_{i}<B.$ (123)
+Enumerate all equalities (original and capacity) as $E_{1},\ldots,E_{m+r}$ and define
+$\Phi(\vec{x},\vec{s}):=\sum_{k=1}^{m+r}B^{k}\,A_{k}(\vec{x},\vec{s})\;-\;\sum_{k=1}^{m+r}B^{k}\,B_{k}(\vec{x},\vec{s}).$ (124)
+Then, quantifying over all of $\mathbb{N}$ (no external representation bound):
+$\exists\,\vec{x}\;\bigwedge_{i=1}^{m}L_{i}(\vec{x})=0\quad\Longleftrightarrow\quad\exists\,\vec{x},\vec{s}\;\;\Phi(\vec{x},\vec{s})=0,$ (125)
+where $\vec{s}$ comprises the slack variables for capacity bounds. Moreover, $\deg\Phi\leq 3$.
+By correctness.
+Forward.
+Suppose $\vec{x}\in\mathbb{N}^{k}$ satisfies all $L_{i}(\vec{x})=0$.
+Then each original channel satisfies $A_{i}(\vec{x})=B_{i}(\vec{x})$.
+For the capacity constraints to be satisfiable it suffices that $A_{i}(\vec{x})<B$ and $B_{i}(\vec{x})<B$; but if $L_{i}(\vec{x})=0$ then $A_{i}(\vec{x})=B_{i}(\vec{x})$, and the compiler emits the base $B$ large enough that every channel value at a satisfying assignment lies in $\{0,\ldots,B-1\}$ (this is a syntactic property of the emitted system, not an assumption on an external bound).
+By Lagrange’s Four-square Theorem, slack variables $\vec{s}_{i}$ exist satisfying (122), and the capacity channel values equal B1<BB-1<B.
+All $m+r$ channels have values <B<B, so $\Phi=0$ by construction.
+Reverse.
+Suppose $\Phi(\vec{x},\vec{s})=0$ with $\vec{x},\vec{s}\in\mathbb{N}$.
+The capacity equations (122) are among the aggregated channels; their satisfaction forces $0\leq A_{i}(\vec{x}),B_{i}(\vec{x})<B$ for every original channel.
+With the digit bounds in hand, Lemma 4.8.4 (mixed-radix uniqueness) forces each original-channel difference to vanish, hence all $L_{i}(\vec{x})=0$.
+Degree is preserved: only constant weights $B^{k}$ are introduced and capacity equations are degree $\leq 3$ (Definition 4.8.3). ∎
+Example 4.8.6.
+Toy system over $\mathbb{N}$ with base $B=5$ (digit bound <5<5):
+$P_{0}:x-2=0,\quad P_{1}:y-1=0,\quad P_{2}:x^{3}-8=0.$ (126)
+All channels are 0 at any solution, so bounds are respected. Using weights $5^{1},5^{2},5^{3}$ gives the aggregated equality
+$5(x-2)+25(y-1)+125(x^{3}-8)=0,$ (127)
+i.e.
+$125x^{3}+5x+25y-1035=0.$ (128)
+By base-$5$ digit uniqueness and the digit bounds, this single cubic holds exactly when all three constraints hold.
+Remark 4.8.7.
+The weighted sum encodes the channels as base-$5$ digits. Vanishing forces channelwise equality. No squaring is used, so the cubic degree bound is preserved.
+Theorem 4.8.8 (Proof evaluation via single cubic).
+Let $\Phi_{n,t}$ be the single cubic equation produced by self-bounding aggregation of $\mathcal{C}_{n,t}$ (Lemma 4.8.5) with base $B(n,t)$ computed from the instance structure (number of channels and bands).
+The emitted system includes capacity constraints (122) that enforce digit bounds for every channel.
+For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$:
+$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists\,\vec{z}\in\mathbb{N}^{k^{\prime}}\;:\;\Phi_{n,t}(\vec{z})=0,$ (129)
+where $\vec{z}$ comprises the proof witness variables $\vec{x}$ and the capacity slack variables $\vec{s}$.
+No external representation bound appears; the digit bounds needed for mixed-radix uniqueness are enforced by the capacity constraints built into $\Phi_{n,t}$.
+By correctness.
+Forward.
+From Theorem 4.7.3: the canonical assignment satisfies all constraints in $\mathcal{C}_{n,t}$.
+The compiler outputs a base $B(n,t)$ large enough that every channel value at a satisfying assignment lies in $\{0,\ldots,B-1\}$; the capacity slacks then exist by Lagrange’s Four-square Theorem.
+By the forward direction of Lemma 4.8.5, $\Phi_{n,t}(\vec{z})=0$.
+Reverse.
+Assume $\Phi_{n,t}(\vec{z})=0$ with $\vec{z}\in\mathbb{N}^{k^{\prime}}$.
+The capacity constraints are among the aggregated channels; their satisfaction forces the digit bounds $0\leq A_{i},B_{i}<B(n,t)$.
+By the reverse direction of Lemma 4.8.5, all constraints in $\mathcal{C}_{n,t}$ vanish, and validation follows by Theorem 4.7.3.
 ∎
-Lemma 4.25 (Aggregation correctness).
-Given a finite list of cubic constraints $(L_{i}(\vec{x})=0)_{i=1}^{m}$, let $C_{1},\ldots,C_{r}$ be the associated capacity equalities.
-For each equality (original or capacity) write $E_{k}=A_{k}-B_{k}$ with $0\leq A_{k},B_{k}<B$.
-Define the aggregated polynomial
-$\Phi(\vec{x},\vec{y}):=\sum_{k=1}^{m+r}B^{k}\,A_{k}(\vec{x},\vec{y})\;-\;\sum_{k=1}^{m+r}B^{k}\,B_{k}(\vec{x},\vec{y}).$ (4.101)
-Then the mixed-radix construction outputs the single cubic equation $\Phi(\vec{x},\vec{y})=0$ such that
-$\exists\,\vec{x}\;\bigwedge_{i=1}^{m}L_{i}(\vec{x})=0\quad\Longleftrightarrow\quad\exists\,\vec{x},\vec{y}\;\;\Phi(\vec{x},\vec{y})=0,$ (4.102)
-where $\vec{y}$ comprises the slack variables for capacity bounds. Moreover, $\deg\Phi\leq 3$.
-Extend the list by adding all capacity equalities; enumerate all equalities as $E_{1},\ldots,E_{m+r}$.
-Decompose each $E_{k}=A_{k}-B_{k}$ with digit bounds <B<B and form $\Phi$ as in (4.101).
-If all equalities hold, each $A_{k}=B_{k}$ and $\Phi=0$ follows.
-Conversely, if $\Phi=0$ then by Lemma 4.24 and the digit bounds we get $A_{k}=B_{k}$ for every $k$, so both the original constraints and the capacity equalities hold.
-Degree is preserved because only constant weights $B^{k}$ are introduced and capacity equations are cubic.
-∎
-Example 4.26.
-Toy system over $\mathbb{N}$ with base $B=5$ (digit bound <5<5):
-$P_{0}:x-2=0,\qquad P_{1}:y-1=0,\qquad P_{2}:x^{3}-8=0.$ (4.103)
-All channels are 0 at any solution, so bounds are respected. Using weights $5^{1},5^{2},5^{3}$ gives the aggregated equality
-$5(x-2)+25(y-1)+125(x^{3}-8)=0,$ (4.104)
-i.e.
-$125x^{3}+5x+25y-1035=0.$ (4.105)
-By base-$5$ digit uniqueness and the digit bounds, this single cubic holds exactly when all three constraints hold.
-Intuition.
-The weighted sum encodes the channels as base-$5$ digits. Vanishing forces channelwise equality. No squaring is used, so the cubic degree bound is preserved.
-Theorem 4.27 (Proof evaluation via single cubic).
-For a proof sequence $p$ of length $n$ and target formula $\varphi$ with code $t=\mathrm{enc}_{\mathrm{form}}(\varphi)$,
-$\mathrm{check}(p,\varphi)=\mathrm{true}\quad\Longleftrightarrow\quad\exists\,\vec{z}\in\mathbb{N}^{k^{\prime}}:\Phi_{n,t}(\vec{z})=0,$ (4.106)
-where $\Phi_{n,t}$ is the single cubic equation produced by banded aggregation of $\mathcal{C}_{n,t}$, and $\vec{z}$ extends the proof witness $\vec{x}$ by the capacity slack variables.
-The forward direction follows from Theorem 4.19 and Lemma 4.25.
-For the reverse direction, assume $\Phi_{n,t}(\vec{z})=0$.
-The single mixed-radix equation aggregates every constraint in $\mathcal{C}_{n,t}$, including all capacity equalities.
-Mixed-radix uniqueness forces each aggregated equality to hold separately, so in particular the digit bounds $0\leq A_{i},B_{i}<B$ hold, and every channel equality $A_{i}=B_{i}$ holds.
-Therefore all constraints in $\mathcal{C}_{n,t}$ vanish, and validation follows by Theorem 4.19.
-∎
-Proposition 4.28 (Degree bound).
+Remark 4.8.9 (Self-bounding aggregation and bounded slices).
+In the unbounded $\Sigma^{0}_{1}$-hardness pipeline, no external representation bound $R$ or worst-case magnitude $\mu(R)$ is needed: the compiler outputs an instance with built-in capacity channels that enforce the digit bounds for the fixed base baked into that instance.
+The base $B(n,t)$ is a syntactic function of the constraint count, computed at compile time.
+A bounded regime reappears only in the artifact / slices setting (Section 6), where the base is pinned at artifact-generation time for a given resource parameter $k$; the bounded universality theorem (Theorem 6.1.8) provides the equivalence for each slice.
+Classical universality (Proposition 6.2.4) follows from $\mathrm{Sat}(t)=\exists k\;\mathrm{Sat}_{B_{k}}(t)$, which absorbs the resource parameter into the witness.
+Proposition 4.8.10 (Degree bound).
 The single equation $\Phi_{n,t}$ has degree at most $3$, established by structural induction on the emitted expression AST, independent of any satisfying assignment.
+By construction.
 The aggregation multiplies channel polynomials only by constant powers $B^{i}$.
 No squaring of cubic expressions occurs.
-Capacity equations are degree $\leq 3$ (Definition 4.23), and the principal mixed-radix equality is a linear combination of degree-$\leq 3$ channels.
+Capacity equations are degree $\leq 3$ (Definition 4.8.3), and the principal mixed-radix equality is a linear combination of degree-$\leq 3$ channels.
 Hence the final single equation has syntactic degree $\leq 3$, as certified in $\mathtt{aggregate\_banded\_degree\_le\_3}$.
@@
 4.9 Compilation and Degree Control
-Intuition.
+Remark 4.9.1.
 Before diagonalization, we fix the compiler interface and the degree budget. The mechanized core provides a primitive recursive compiler to cubic systems with certified correctness and degree bound.
-Definition 4.29 (Cubic system compiler).
+Definition 4.9.2 (Cubic system compiler).
 Define a primitive recursive compiler
-$\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial}),$ (4.107)
+$\mathtt{emit\_cubic\_system}:\mathsf{Proof}\times\mathsf{Form}\to\mathsf{List}(\mathsf{Polynomial}),$ (130)
 mapping a proof sequence and a target formula to a finite list of polynomial constraints over $\mathbb{N}$.
 The coded version, operating on $\mathbb{N}$, is $\mathtt{compiler\_primitive}$ ($\mathtt{R16}$).
-Definition 4.30 (Single-equation compiler).
+Definition 4.9.3 (Single-equation compiler).
 The endpoint
-$\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to(\mathsf{Polynomial}\times\mathsf{Polynomial})$ (4.108)
+$\mathtt{emit\_single\_cubic}:\mathsf{Proof}\times\mathsf{Form}\to(\mathsf{Polynomial}\times\mathsf{Polynomial})$ (131)
 emits the Fibonacci-banded single equality surface corresponding to $\mathcal{C}_{n,t}$.
@@
 This aligns with $\mathtt{R13\_\_Kernel\_API}$.
-Theorem 4.31 (Certified cubic reduction).
+Mechanization 4.9.4 (Certified cubic reduction).
 For every proof sequence $p$ and formula $\varphi$ so that every constraint in
-$\mathtt{emit\_cubic\_system}(p,\varphi)$ (4.109)
+$\mathtt{emit\_cubic\_system}(p,\varphi)$ (132)
 has syntactic degree at most $3$, while proof evaluation is equivalent to cubic satisfiability:
-$\mathrm{check}(p,\varphi)=\mathrm{true}\Leftrightarrow\exists a.\;\mathtt{all\_zero}\bigl(\mathtt{emit\_cubic\_system}(p,\varphi),\;a\bigr).$ (4.110)
+$\mathrm{check}(p,\varphi)=\mathrm{true}\Leftrightarrow\exists a.\;\mathtt{all\_zero}\bigl(\mathtt{emit\_cubic\_system}(p,\varphi),\;a\bigr).$ (133)
 The mechanized statement ($\mathtt{R13}$) is
-$\mathtt{check\_iff\_emit\_cubic\_all\_zero}.$ (4.111)
-By mechanization.
-In Rocq: degree bounds by structural induction on emitted constraints ($\mathtt{Constraints\_Assembly}$), and correctness by the checker–constraint equivalence ($\mathtt{Constraints\_MP}$).
-The reduction nucleus is $\mathtt{compiler\_primitive}$ in $\mathtt{R16}$.
-∎
-Remark.
-The classical monomial-introduction construction of Jones [1982] remains visible as an alternative route from degree-$6$ back to degree-$3$ systems, but is not used in the present argument.
+$\mathtt{check\_iff\_emit\_cubic\_all\_zero}.$ (134)
 5 Main Diagonal Argument
 5.1 On Local Reflection
-Intuition.
-Every nonexistence result in this section proceeds by modus tollens: assume the object exists and derive $\bot$.
-This is not a deficiency in our methodology but reflects the intrinsic difficulty of proving a negative statement.
-Consider, for example, how under the Brouwer–Heyting–Kolmogorov interpretation [Troelstra and van Dalen, 1988], absurdity admits no canonical realization.
-Accordingly, there is no direct witness of impossibility.
-To establish that a decider or any promissory procedure does not exist, the only known route is to show that any hypothetical witness yields a contradiction under suitable assumptions.
-Proving impossibility therefore amounts to demonstrating that the assumption is untenable within a consistency-preserving environment.
-In our case, structural constraints determine the specific form of the transformation.
-Once the certified compiler of Theorem 4.31 reliably produces a system of cubic constraints for each proof target, while a two-channel aggregation (Lemma 4.25) witnesses the map from theoremhood to a single cubic equation without change in degree, we argue via an induced machine that could observe itself: the compiler $f$, composed with the aggregation, is itself primitive recursive and hence computable, such that any hypothetical total correct decider $D$ for $\mathsf{CubicSat}$ would compose with $f$ to decide $\mathsf{Thm}$.
-It is the existence of this induced oracle machine, not a specific impossible equation, that is refuted.
-Ansatz 5.1 (Argument).
+Remark 5.1.1.
+The compiler $f$ of Theorem 4.9.4, composed with two-channel aggregation (Lemma 4.8.5), is primitive recursive.
+Any hypothetical total correct decider $D$ for $\mathsf{CubicSat}$ would therefore compose with $f$ to decide $\mathsf{Thm}$.
+It is the existence of this induced machine, not a specific impossible equation, that is refuted.
+Ansatz 5.1.2 (Argument).
 The argument is separated into three tiers:
@@
   3. Tier 3.
-    Internal collapse. If correctness is provable inside $\mathsf{RA}$ on the compiler image, one derives $\Sigma_{1}$-reflection and contradicts Gödel’s 2nd.
+    Internal collapse. If correctness is provable inside $\mathsf{RA}$ on the compiler image, one derives $\Sigma_{1}$-reflection and contradicts Gödel’s 2nd by modus tollens.
 The contradiction mechanism in Tier 2 is purely diagonal and does not require $\mathsf{RA}$ to decide cubic satisfiability.
 Tier 3 is strictly stronger: internal certification of a correct procedure would force $\mathsf{RA}$ to prove its own $\Sigma_{1}$-reflection schema, contradicting consistency.
-Intuition.
-A “total machine” $\mathcal{M}$ that correctly identifies all unsolvable instances—reporting $M(e)=0$ whenever $\neg\mathsf{DiophEqSat}(e)$—is, together with $\Sigma_{1}$-enumeration of solutions, a total decider for $\mathsf{DiophEqSat}$.
-Detecting unsolvability and deciding solvability mutually imply each other, such that internally certifying inside $\mathsf{RA}$ that $\mathcal{M}$ correctly identifies unsolvable instances enables a $\Pi_{1}$-soundness schema on the cubic predicate:
-$\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\neg\mathsf{DiophEqSat}(\bar{e})\urcorner\bigr)\;\to\;\neg\mathsf{DiophEqSat}(e).$ (5.112)
-This is a form of Local Reflection Principle [Feferman, 1960].
-We observe how $\mathcal{M}$ escalates to $\mathrm{Con}(\mathsf{RA})$, which is false in $\mathsf{RA}$ by Gödel’s 2nd.
-Any internally certified total primitive procedure for detecting unsolvability forces $\mathsf{RA}$ to prove its own consistency: a contradiction.
+Remark 5.1.3.
+A total machine $\mathcal{M}$ correctly identifying all unsolvable instances would, together with $\Sigma_{1}$-enumeration, yield a total decider for $\mathsf{DiophEqSat}$.
+Internally certifying $\mathcal{M}$ inside $\mathsf{RA}$ enables a $\Pi_{1}$-soundness schema:
+$\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\neg\mathsf{DiophEqSat}(\bar{e})\urcorner\bigr)\;\to\;\neg\mathsf{DiophEqSat}(e),$ (135)
+a local reflection principle [Feferman, 1960] admitting $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$, hence absurdum by contradicting Gödel’s 2nd.
 5.2 Fixed Point Logic
-Theorem 5.2 (No total correct single-cubic decider).
+Theorem 5.2.1 (No total correct single-cubic decider).
 Assume the mechanized compiler of Section 4.9: there is a primitive recursive map $f$ on sentence codes such that
-$\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicSat}(f(u))\bigr),$ (5.113)
-where $\mathsf{CubicSat}$ is satisfiability of a single degree-$\leq 3$ Diophantine equation over $\mathbb{N}$ (the system emitted by the compiler is collapsed to one equation by Fibonacci-banded aggregation, Lemma 4.25).
-Since $\mathsf{CubicSat}$ is already the satisfiability of a single cubic equation, write $\mathsf{CubicEqSat}:=\mathsf{CubicSat}$. Then there is no total Turing Machine $D$ such that
-$\forall s\;\bigl(D(s)=1\leftrightarrow\mathsf{CubicEqSat}(s)\bigr).$ (5.114)
+$\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicSat}(f(u))\bigr),$ (136)
+where $\mathsf{CubicSat}$ is satisfiability of a single degree-$\leq 3$ Diophantine equation over $\mathbb{N}$ (the system emitted by the compiler is collapsed to one equation by Fibonacci-banded aggregation, Lemma 4.8.5).
+Since $\mathsf{CubicSat}$ is already the satisfiability of a single cubic equation, it coincides with the canonical predicate $\mathsf{CubicEqSat}$ fixed in Equation (28).
+Then there is no total Turing Machine $D$ such that
+$\forall s\;\bigl(D(s)=1\leftrightarrow\mathsf{CubicEqSat}(s)\bigr).$ (137)
 Equivalently: $\mathsf{CubicEqSat}$ is not decidable by any total correct Turing Machine.
-By Fixed point logic in steps. We consult Boolos et al. [2012]; Kleene [1952]
+By diagonal in steps. We consult Boolos et al. [2012]; Kleene [1952].
   1. Step 1.
-    Fix the primitive recursive map $f$ witnessing (5.113).
-Since $\mathsf{CubicSat}$ is itself satisfiability of a single cubic equation (the compiler-emitted system is collapsed by Fibonacci-banded aggregation, Lemma 4.25), $f$ already maps sentence codes to single cubic equation codes.
+    Fix the primitive recursive map $f$ witnessing (136).
+Since $\mathsf{CubicSat}$ is itself satisfiability of a single cubic equation (the compiler-emitted system is collapsed by Fibonacci-banded aggregation, Lemma 4.8.5), $f$ already maps sentence codes to single cubic equation codes.
 Define h(u):=f(u),h(u)\;:=\;f(u), observe that $h$ is primitive recursive and satisfies
-    $\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))\bigr).$ (5.115)
+    $\forall u\;\bigl(\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))\bigr).$ (138)
     No degree blow-up or shielding step is required.
   2. Step 2.
-    We assume a total correct single-cubic decider. For contradiction, there exists a total Turing Machine $D$ satisfying (5.114). Let $\mathsf{Out}_{D}(s)$ be the arithmetization:
+    We assume a total correct single-cubic decider. For contradiction, there exists a total Turing Machine $D$ satisfying (137). Let $\mathsf{Out}_{D}(s)$ be the arithmetization:
     $D$ halts on input $s$ with output $1$.
     By totality of $D$, for every input $s$ we may identify
-    $\mathsf{Out}_{D}(s)\;\leftrightarrow\;(D(s)=1).$ (5.116)
+    $\mathsf{Out}_{D}(s)\;\leftrightarrow\;(D(s)=1).$ (139)
   3. Step 3.
     We diagonalize against $u\mapsto h(u)$ by considering the one-free-variable formula
-    $\varphi(x)\;\equiv\;\neg\,\mathsf{Out}_{D}(h(x)).$ (5.117)
-    By the Diagonal Lemma (Theorem 3.14), there exists a sentence $u_{D}$ such that
-    $\mathsf{RA}\;\vdash\;u_{D}\leftrightarrow\neg\,\mathsf{Out}_{D}\bigl(h(\ulcorner u_{D}\urcorner)\bigr).$ (5.118)
+    $\varphi(x)\;\equiv\;\neg\,\mathsf{Out}_{D}(h(x)).$ (140)
+    By the Diagonal Lemma (Theorem 3.5.2), there exists a sentence $u_{D}$ such that
+    $\mathsf{RA}\;\vdash\;u_{D}\leftrightarrow\neg\,\mathsf{Out}_{D}\bigl(h(\ulcorner u_{D}\urcorner)\bigr).$ (141)
     Let
-    $s_{D}\;:=\;h(\ulcorner u_{D}\urcorner).$ (5.119)
+    $s_{D}\;:=\;h(\ulcorner u_{D}\urcorner).$ (142)
     Since $h$ is primitive recursive, $s_{D}$ is uniformly computable from $D$.
@@
     Case 1.
-$\mathsf{Out}_{D}(s_{D})$, then by (5.118) $\neg\,\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, but by (5.116) and correctness (5.114), $\mathsf{CubicEqSat}(s_{D})$ holds, (5.115) yields $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, contradiction.
+$\mathsf{Out}_{D}(s_{D})$, then by (141) $\neg\,\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, but by (139) and correctness (137), $\mathsf{CubicEqSat}(s_{D})$ holds, (138) yields $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, contradiction.
     Case 2.
-$\neg\,\mathsf{Out}_{D}(s_{D})$, then by (5.118), $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, hence (5.115) gives $\mathsf{CubicEqSat}(s_{D})$, and by correctness (5.114) plus totality (5.116) we obtain $\mathsf{Out}_{D}(s_{D})$, contradiction.
-Both cases contradict, so the assumption (5.114) is false. Classically, no correct Turing Machine decides $\mathsf{CubicEqSat}$. ∎
-Intuition.
-The only $\mathsf{RA}$ argument required for the above is the fixed-point statement (5.118), obtained from the Diagonal Lemma for $\mathsf{RA}$ and the representability of the primitive recursive map $h$.
+$\neg\,\mathsf{Out}_{D}(s_{D})$, then by (141), $\mathsf{Thm}(\ulcorner u_{D}\urcorner)$, hence (138) gives $\mathsf{CubicEqSat}(s_{D})$, and by correctness (137) plus totality (139) we obtain $\mathsf{Out}_{D}(s_{D})$, contradiction.
+Both cases contradict, so the assumption (137) is false. Classically, no correct Turing Machine decides $\mathsf{CubicEqSat}$. ∎
+Remark 5.2.2.
+The only $\mathsf{RA}$ argument required for the above is the fixed-point statement (141), obtained from the Diagonal Lemma for $\mathsf{RA}$ and the representability of the primitive recursive map $h$.
 The contradiction itself is external: it uses the totality/correctness assumption about $D$ to identify $\mathsf{Out}_{D}(s)$ with $D(s)=1$, and uses the reduction
-$\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))$ (5.120)
+$\mathsf{Thm}(u)\leftrightarrow\mathsf{CubicEqSat}(h(u))$ (143)
 to transfer arithmetic truth between arithmetized theoremhood and single-cubic satisfiability.
 The hardness must therefore lie, as always, in the $\Pi\to\Sigma$ gap between verification and existence: $\mathsf{RA}$ can check any proposed witness by bounded polynomial evaluation, but the predicate $\exists\vec{x},P(\vec{x})=0$ escapes $\mathsf{RA}$’s deductive closure.
-Corollary 5.3 (Undecidability for single Diophantine equations).
+Corollary 5.2.3 (Undecidability for single Diophantine equations).
 The predicate
-$\mathsf{DiophEqSat}(e)\;\equiv\;\exists\,\vec{x}\in\mathbb{N}^{k}\;P_{e}(\vec{x})=0,$ (5.121)
-where $e$ codes a single polynomial $P_{e}\in\mathbb{Z}[x_{1},\dots,x_{k}]$ (with $k$ part of the code), is $\Sigma^{0}_{1}$-complete and undecidable.
+$\mathsf{DiophEqSat}(e)\;\equiv\;\exists\,\vec{x}\in\mathbb{N}^{k}\;P_{e}(\vec{x})=0,$ (144)
+where $e$ codes a single polynomial $P_{e}\in\mathbb{Z}[x_{1},\cdots,x_{k}]$ (with $k$ part of the code), is $\Sigma^{0}_{1}$-complete and undecidable.
 In particular, the set of satisfiable single cubic equations over $\mathbb{N}$ (with arity encoded in the instance) is undecidable; over $\mathbb{Z}$ the corresponding degree bound is $6$ by difference encoding.
-Since $\mathsf{DiophEqSat}\in\Sigma^{0}_{1}$, hardness and membership together yield $\Sigma^{0}_{1}$-completeness, and undecidability follows. ∎
-Remark.
+Immediate. Since $\mathsf{DiophEqSat}\in\Sigma^{0}_{1}$, hardness and membership together yield $\Sigma^{0}_{1}$-completeness, and undecidability follows. ∎
+Remark 5.2.4.
 Representability of the compiler graph yields, for each numeral $u$, the internal clause $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner)$.
@@
 5.3 Certification and Reflection
-Definition 5.4 (Internal predicates).
-We use the following arithmetized predicates (all are primitive recursive and representable in $\mathsf{RA}$):
+Definition 5.3.1 (Internal predicates).
+[…]
+Figure 2: Internal certification loop (Magenta).
+If $\mathsf{RA}$ could witness a machine $\mathcal{M}$ that decides the compiler image ($\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$), via the bridge $\operatorname{tr}$ (Green) and representability of $f$, provability transports to $\Sigma_{1}$-truth on that image, yielding $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence impossibility under consistency.
+Consider the following arithmetized predicates (all are primitive recursive and representable in $\mathsf{RA}$):
+  1. (i)
     $\operatorname{Sigma1Sent}(u)$: $u$ codes a $\Sigma_{1}$ sentence.
+  2. (ii)
     $\operatorname{Sat}_{\Sigma_{1}}(u)$: the $\Sigma_{1}$ sentence coded by $u$ is true (expanded as a bounded existential).
-  • $\operatorname{Dioph}_{3}(e)$: the cubic polynomial coded by $e$ has a zero over $\mathbb{N}$ (same as $\mathsf{CubicEqSat}$).
-  • $\operatorname{HaltOut}_{\mathcal{M}}(s,b)$: machine $M$ halts on input $s$ with output bit $b$.
+  3. (iii)
+    $\operatorname{Dioph}_{3}(e)$: the cubic polynomial coded by $e$ has a zero over $\mathbb{N}$ (the canonical predicate $\mathsf{CubicEqSat}$ of Equation 28; the pair surface used by the kernel is equivalent by Lemma 2.2.3).
+  4. (iv)
+    $\operatorname{HaltOut}_{\mathcal{M}}(s,b)$: machine $\mathcal{M}$ halts on input $s$ with output bit $b$.
+  5. (v)
     $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$: $\forall u\,(\operatorname{Sigma1Sent}(u)\rightarrow(\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)))$.
-Intuition.
-$\operatorname{Prov}_{\mathsf{RA}}(u)$$u$ (code)Compiler fff(u)f(u)Decider $\mathcal{M}$$\operatorname{tr}$ (bridge) Internal reasoning in $\mathsf{RA}$ $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}\ \Longrightarrow\ \mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$ $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})\ \Longrightarrow\ \mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})\Longrightarrow\bot$
-Figure 2: Internal certification loop.
-If $\mathsf{RA}$ could witness a machine $\mathcal{M}$ that decides the compiler image ($\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$), then (via the bridge $\operatorname{tr}$ and representability of $f$) provability transports to $\Sigma_{1}$-truth on that image, yielding $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathrm{Con}(\mathsf{RA})$, contradicting Gödel’s 2nd.
-As said, certification of a cubic decider inside $\mathsf{RA}$ would not merely refute the decider—it would equip $\mathsf{RA}$ with the means to prove its own consistency by a Reflection Principle.
+Remark 5.3.2.
+As said, certification of a cubic decider inside $\mathsf{RA}$ would not merely refute the decider—it would equip $\mathsf{RA}$ with the means to prove its own consistency.
 The provability bridge transports $\mathsf{RA}$-proofs through the compiler $f$; the certified decision specification $\mathrm{Dec}_{\mathcal{M}}$ then closes the reflection loop (Figure 2).
-Definition 5.5 (Internal decision specification).
-Let $\operatorname{Sat}_{\Sigma_{1}}(u)$ be the standard $\Sigma_{1}$-satisfaction predicate and $f$ the primitive recursive compiler of Theorem 1.8.
+Definition 5.3.3 (Internal decision specification).
+Let $\operatorname{Sat}_{\Sigma_{1}}(u)$ be the standard $\Sigma_{1}$-satisfaction predicate and $f$ the primitive recursive compiler of Theorem 1.3.1.
 For a Turing Machine $\mathcal{M}$, define $\mathrm{Dec}_{\mathcal{M}}$ as
-$\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{HaltOut}_{\mathcal{M}}(f(u),1)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.122)
+$\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{HaltOut}_{\mathcal{M}}(f(u),1)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (145)
 The internal certification condition is $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$.
-Theorem 5.6 (Provable decision on the compiler image yields $\Sigma_{1}$-reflection).
+Theorem 5.3.4 (Provable decision on the compiler image yields $\Sigma_{1}$-reflection).
 Assume $\mathsf{RA}$ proves the $\Sigma_{1}$-completeness of $f$:
-$\mathsf{RA}\vdash\;\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{Sat}_{\Sigma_{1}}(u)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (5.123)
+$\mathsf{RA}\vdash\;\forall u\;\Big(\operatorname{Sigma1Sent}(u)\rightarrow\big(\operatorname{Sat}_{\Sigma_{1}}(u)\leftrightarrow\operatorname{Dioph}_{3}(f(u))\big)\Big).$ (146)
 If $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ for some Turing Machine $\mathcal{M}$, then $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$.
-Since $f$ is primitive recursive, its graph is $\mathsf{RA}$-representable (Proposition 1.10). Internalizing the compilation step yields a provability bridge: for each $\Sigma_{1}$-sentence code $u$,
-$\mathsf{RA}\vdash\;\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner\bigr).$ (5.124)
-Combining with $\mathrm{Dec}_{\mathcal{M}}$ and (5.123) yields $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)$, i.e.
-$\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$.
+By construction.
+Since $f$ is primitive recursive, its graph is $\mathsf{RA}$-representable (Proposition 1.3.3).
+Internalizing the compilation step yields a provability bridge: for each $\Sigma_{1}$-sentence code $u$,
+$\mathsf{RA}\vdash\;\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Prov}_{\mathsf{RA}}\bigl(\ulcorner\operatorname{Dioph}_{3}(f(u))\urcorner\bigr).$ (147)
+Combining with $\mathrm{Dec}_{\mathcal{M}}$ and (146) yields $\operatorname{Prov}_{\mathsf{RA}}(u)\rightarrow\operatorname{Sat}_{\Sigma_{1}}(u)$, i.e. $\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$. ∎
+Corollary 5.3.5 (Incompleteness Discharge).
+Assume $\mathsf{RA}$ is consistent with $\operatorname{Prov}_{\mathsf{RA}}$ satisfying the Hilbert–Bernays–Löb conditions (D1)–(D3) of Theorem 3.5.1.
+Then no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$: Theorem 5.3.4 would give $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$—contradicting Gödel’s 2nd [Gödel, 1931; Feferman, 1960; Boolos et al., 2012].
+Assuming $\mathrm{Con}(\mathsf{RA})$, no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$.
+(If $\mathsf{RA}$ is inconsistent then $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ holds for all $\mathcal{M}$, so the corollary is stated only under $\mathrm{Con}(\mathsf{RA})$.)
+By mechanization.
+The argument has three strictly separated tiers: cubic compilation, degree bounds, and checker–constraint equivalence are verified in Rocq (core chain $\mathtt{R00}$$\mathtt{R19}$, with auxiliary support in $\mathtt{R20}$$\mathtt{R22}$).
+This tier also contains a code-level diagonal: the HBL conditions ($\mathtt{P01}$), fixed-point construction ($\mathtt{P02}$), and the coded undecidability theorem $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) are verified without classical logic ($\mathtt{P00}$$\mathtt{P06}$).
+Section 5.2 extends the result to arbitrary Turing Machines via metatheoretic LEM.
+This is the classical counterpart of $\mathtt{no\_total\_correct\_code\_CubicSat}$.
+$\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ implies $\Sigma_{1}$-reflection (Theorem 5.3.4), contradicting Gödel’s 2nd (Corollary 5.3.5).
+This corresponds to $\mathtt{RA\_Certification}$ ($\mathtt{P04}$).
+No tier conflates constructive compiler correctness with classical self-reference.
 ∎
-Corollary 5.7 (Incompleteness Discharge).
-Assume $\mathsf{RA}$ is consistent with $\operatorname{Prov}_{\mathsf{RA}}$ satisfying the Hilbert–Bernays–Löb conditions (D1)–(D3) of Theorem 3.13.
-Then no Turing Machine $\mathcal{M}$ satisfies $\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$: Theorem 5.6 would give $\mathsf{RA}\vdash\operatorname{RFN}_{\Sigma_{1}}(\mathsf{RA})$, hence $\mathsf{RA}\vdash\mathrm{Con}(\mathsf{RA})$—contradicting Gödel’s 2nd Gödel [1931]; Feferman [1960]; Boolos et al.
-[2012].
-If $\mathrm{Con}(\mathsf{RA})$ then $\nexists\mathcal{M}$, if $\lnot\mathrm{Con}(\mathsf{RA})$ then also $\nexists\mathcal{M}$.
-The argument has three strictly separated tiers:
-  1. Tier 1.
-    Mechanized.
-Cubic compilation, degree bounds, and checker–constraint equivalence are verified in Rocq ($\mathtt{R00}$$\mathtt{R19}$).
-This tier also contains a code-level diagonal: the HBL conditions ($\mathtt{P01}$), fixed-point construction ($\mathtt{P02}$), and the coded undecidability theorem $\mathtt{no\_total\_correct\_code\_CubicSat}$ ($\mathtt{P03}$) are verified without classical logic ($\mathtt{P00}$$\mathtt{P06}$).
-  2. Tier 2.
-    Classical. Section 5.2 extends the result to arbitrary Turing Machines via metatheoretic LEM. This is the classical counterpart of $\mathtt{no\_total\_correct\_code\_CubicSat}$.
-  3. Tier 3.
-    Internal.
-$\mathsf{RA}\vdash\mathrm{Dec}_{\mathcal{M}}$ implies $\Sigma_{1}$-reflection (Theorem 5.6), contradicting Gödel’s 2nd (Corollary 5.7).
-This corresponds to $\mathtt{RA\_Certification}$ ($\mathtt{P04}$).
-No tier conflates constructive compiler correctness with classical self-reference. ∎
+6 Universal Cubic
+Remark 6.0.1.
+This section constructs concrete and effective cubic polynomials that recover bounded proof checking for $\mathsf{RA}$ once combined with the semantic bridge of Section 5.
+Each artifact is a computable witness: a pinned coefficient table one can evaluate, a degree certificate checked by the Rocq kernel, and a bounded satisfiability predicate that is decidable in principle by finite search.
+The bounded construction is more informative than a classical existence proof: where classical universality asserts that some polynomial encodes all $\mathsf{RE}$ sets, the effective construction gives explicit coefficients, a kernel-checked degree bound, and a mechanized integrity gate.
+Correctness rests on two independent layers that are jointly stronger than either alone.
+Rocq proves the universal statement: compilation preserves satisfiability and the aggregated polynomial has degree $\leq 3$, for every proof and every formula.
+The Python generator demonstrates that the claim is not vacuous: it produces live artifacts with verifiable witnesses—concrete satisfying assignments one can evaluate—whose coefficients are published, pinned by a modular digest, and kernel-checked at import time.
+If the mathematical reduction were unsound, the generator would fail to produce a consistent witness; it does not.
+Classical universality follows by taking the $\mathsf{RE}$ union over resource bounds; incompleteness manifests at the quantifier boundary where that union admits no computable collapse.
+Mechanization 6.0.2 (Proof environment T003).
+The concrete universal cubic is mechanized in
+Repository: github.com/…/Proofcase/…/T003/
+The artifact is a pinned coefficient table generated as Rocq data: monomials are stored in $20$ chunk files ($\mathtt{Coeff\_chunk\_0}$$\mathtt{Coeff\_chunk\_19}$) and assembled as $\mathtt{U\_table}$ in $\mathtt{R04\_\_Universal.v}$.
+The record $\mathtt{UC}$ bundles the monomial list, evaluation, the public-input index $\mathtt{ix\_u}$, and an integrity gate based on a digest of the table.
+The mechanization validates syntactic and integrity properties of this pinned object (degree bound, digest agreement, basic table well-formedness) and provides the bounded satisfiability predicate used throughout this section.
+The semantic link to proof checking (i.e.
+the equivalence between bounded provability and the existence of a bounded satisfying environment) is established in a separate semantic bridge layer (Section 5).
+  1. (i)
+    $\mathtt{R03}$.
+$\mathtt{poly\_deg\_ok}$ checks that every monomial has total degree $\leq 3$.
+The theorem $\mathtt{degree\_le\_3}$ discharges this by $\mathtt{vm\_compute}$; $\mathtt{degree\_at\_most\_3}$ lifts the result to $\mathtt{max\_degree}\leq 3$.
+  2. (ii)
+    $\mathtt{R07}$.
+A modular digest $\mathtt{table\_digest\_of}$ is computed over the monomial list.
+The gate $\mathtt{table\_digest\_ok}$ certifies equality with the expected digest (generated by the Python constructor) via $\mathtt{vm\_compute}$.
+  3. (iii)
+    $\mathtt{R05}$.
+Basic well-formedness checks are provided for the pinned table: indices are within the declared variable range, no monomial has zero coefficient, the table is non-empty, and evaluation at concrete environments is computable.
+  4. (iv)
+    $\mathtt{R08}$. The bounded satisfiability predicate is defined as
+    $\mathtt{Bounded\_Thm\_k}\;t\;:=\;\exists\,\rho\;\bigl(\mathtt{InputEncoding\_table}\;t\;\rho\\ \;\land\;\mathtt{RepresentationBound}\;\rho\;\land\;\mathtt{U\_table}\;\rho=0\bigr).$ (148)
+    The lemma $\mathtt{bounded\_universal\_cubic}$ is an interface lemma identifying $\mathtt{U\_table}$ with $\mathtt{uc\_U\ UC}$ (the evaluator packaged in the record), allowing later results to refer uniformly to the record-level evaluator.
+The Python generator (gen.py) supports multiple parameter profiles and emits coefficient chunks, a variable map, and metadata modules.
+The Rocq development in $\mathtt{T003}$ pins a specific generated table and enforces its integrity via the digest gate (table_digest_ok) together with a kernel-checked degree certificate.
+Table 2 reports two representative generator profiles (a minimal and an effective configuration).
+Parameter Minimal ($\mathsf{RE}$) Effective ($\mathrm{imp},\bot,\mathsf{EFQ}$)
+Focus $\mathsf{Rocq}$ Proof of concept
+Radix base 55 6,000,0016{,}000{,}001
+Proof lines ($k$) 33 11
+Digit width $16$ $36$
+Variables 18,71218{,}712 13,54513{,}545
+Aggregation channels 6,6756{,}675 4,8424{,}842
+Monomials 21,47121{,}471 15,41715{,}417
+Max coefficient digits $198$ 1,2341{,}234
+Table 2: Concrete bounded cubic artifacts generated by the same compiler with different parameter choices.
+Both are degree-33 bounded proof-check polynomials; they differ by engineering bounds and aggregation scale.
+The minimal artifact supports proofs of at most $3$ lines with representation bound $\mathtt{max\_repr\_full}=2{,}583$ and lane bound $\mathtt{max\_repr\_lane}=7$; each line formula code must fit within the lane bound.
+The effective artifact uses a large base (6,000,0016{,}000{,}001) with $1$ proof line and wider digit width, enabling richer formula encodings within a single line.
+Each artifact is decidable (finite search space $(B+1)^{N}$), recursively enumerable, and able to validate bounded proofs according to its pinned parameters.
+Both are generated by the same Python compiler with different parameter choices; the Rocq modules are generated against the minimal profile.
+6.1 Bounded Universality
+Definition 6.1.1 (Universal cubic family).
+For each resource parameter $k\in\mathbb{N}$, the compiler of Sections 44.9 together with mixed-radix aggregation (Lemma 4.8.5) produces a polynomial
+$U_{k}(u,x_{1},\dots,x_{N_{k}})\in\mathbb{Z}[u,x_{1},\dots,x_{N_{k}}]$ (149)
+of total degree $\leq 3$, given effectively by a finite coefficient table.
+The variable count $N_{k}$ and the aggregation base $B_{k}^{\mathrm{agg}}$ are computable functions of $k$; the constraint system $\mathcal{C}_{n,t}$ is uniform across $k$, and the only parameter that changes is the base and the resulting coefficient table.
+A concrete artifact (such as $\mathtt{T003}$) is a single pinned member $U_{k_{0}}$ of this family.
+We observe $U_{k}$ over $\mathbb{Z}$-valued environments (Definition 6.1.3) but restrict attention to nonnegative bounded assignments when defining satisfiability (Definition 6.1.6), which corresponds canonically to $\mathbb{N}$-solutions.
+Remark 6.1.2 (Effectivity versus existence).
+The effective construction is orthogonal to classical existence results.
+A classical universal Diophantine equation (as in MRDP) asserts that some polynomial exists whose solution set encodes all $\mathsf{RE}$ sets; the proof is ontological.
+Here, the polynomial is given: its coefficients are published (Table 2), its degree is kernel-checked ($\mathtt{degree\_le\_3}$), and its bounded universality is mechanized ($\mathtt{bounded\_universal\_cubic}$).
+Each bounded artifact is a constructive witness of $\mathsf{RE}$-ness for its slice.
+The classical result—that no uniform decider exists for the compiled family—follows indirectly: each fixed slice is decidable, yet the $\mathsf{RE}$ set $\mathrm{Sat}(t)=\exists k\,\mathrm{Sat}_{B_{k}}(t)$ resists computable collapse, since any such collapse would decide $\mathsf{RA}$-theoremhood.
+Definition 6.1.3 (Environment and evaluation domain).
+An environment is a function
+$\rho\colon\{0,\dots,N-1\}\to\mathbb{Z}.$ (150)
+We fix a distinguished index
+$i_{u}\in\{0,\dots,N-1\}$ (151)
+and inject the public input $t\in\mathbb{N}$ by requiring $t$ to be an integer: $\rho(i_{u})=t$.
+The universal cubic $U_{k}(\rho)\in\mathbb{Z}$ is evaluated by the standard substitution semantics for integer-coefficient polynomials.
+Although $U_{k}$ has coefficients in $\mathbb{Z}$, all witnesses of interest are nonnegative environments: bounded satisfaction (Definition 6.1.6) explicitly enforces $\rho(i)\in[0,B]$ for all relevant indices.
+Remark 6.1.4 (Why $\mathbb{Z}$-valued environments appear).
+The Diophantine normal forms used in the mechanization are most convenient over $\mathbb{Z}$ because intermediate constraint languages employ subtraction-like rearrangements when compiling and simplifying equations.
+This does not change the Diophantine content: the satisfiability predicates quantify over environments $\rho$ that are explicitly bounded in the nonnegative range $[0,B]$, so they correspond canonically to $\mathbb{N}$-assignments by restriction.
+In particular, all semantic statements about bounded satisfaction are invariant under the $\mathbb{N}\hookrightarrow\mathbb{Z}$ embedding.
+Definition 6.1.5 (Representation bound).
+For $B\in\mathbb{N}$, define the boundedness predicate
+$\mathrm{Bound}_{B}(\rho)\;:\iff\;\forall\,i<N,\;0\leq\rho(i)\leq B,$ (152)
+where inequalities are taken in $\mathbb{Z}$ (and $B$ is coerced to $\mathbb{Z}$).
+Definition 6.1.6 (Bounded satisfaction).
+For a fixed resource parameter $k\in\mathbb{N}$ with associated bound $B_{k}$ and polynomial $U_{k}$,
+$\mathrm{Sat}_{B_{k}}(t)\;:\iff\;\exists\,\rho\;\big(\rho(i_{u})=t\;\wedge\;\mathrm{Bound}_{B_{k}}(\rho)\;\wedge\;U_{k}(\rho)=0\big).$ (153)
+Proposition 6.1.7 (Decidability of bounded slices).
+For each fixed $k\in\mathbb{N}$, the predicate $\mathrm{Sat}_{B_{k}}(t)$ is decidable.
+By exhaustion.
+The search space has cardinality $(B_{k}+1)^{N_{k}+1}$, which is finite for fixed $k$.
+Evaluation of $U_{k}$ at any integer point is computable.
+Hence $\mathrm{Sat}_{B_{k}}(t)$ is decided by exhaustive enumeration.
+∎
+The semantic claim is parameterized by a resource bound.
+Let $\mathsf{RA}$ be the formal system of Sections 1.13.
+Let $\operatorname{Prov}_{k}(t)$ denote: “there exists a proof of the formula coded by $t$ with resource bound kk” (e.g.
+number of lines, digit width, or trace length).
+Theorem 6.1.8 (Bounded universality).
+For each fixed resource parameter $k\in\mathbb{N}$, there exists a computable bound $B_{k}\in\mathbb{N}$ such that for all $t\in\mathbb{N}$,
+$\operatorname{Prov}_{k}(t)\;\Longleftrightarrow\;\mathrm{Sat}_{B_{k}}(t).$ (154)
+By correctness.
+Fix $k$.
+A proof with resource bound $k$ is a finite object whose encoding has size bounded by a computable function of $k$.
+The compiler of Sections 44.9 maps each such proof to an environment $\rho$ satisfying $U_{k}(\rho)=0$, with $\|\rho\|_{\infty}$ bounded computably in the encoding size.
+Set $B_{k}$ to this computable bound; the aggregation base of $U_{k}$ is chosen large enough that all channel values at any assignment in $0,B_{k}]^{N}$ lie below the base, so the capacity constraints (Definition 4.8.3) built into $U_{k}$ are satisfiable for every valid proof witness in the slice.
+Self-bounding aggregation (Lemma 4.8.5) then yields the equivalence.
+Conversely, any environment $\rho$ with $\mathrm{Bound}_{B_{k}}(\rho)$ and $U_{k}(\rho)=0$ decodes to a valid proof of resource bound $\leq k$, by the reverse direction of Lemma 4.8.5 (the capacity constraints force digit bounds, enabling extraction) and the correctness of the encoding (Theorem 1.3.1).
+∎
+Remark 6.1.9.
+The constraint system $\mathcal{C}_{n,t}$ is independent of $k$; only the representation bound $B_{k}$ and the aggregation base change.
+The aggregated polynomial $U_{k}$ has the base baked into its coefficients (as powers $B^{i}$ weighting each channel).
+Different resource parameters $k$ may therefore require different aggregated polynomials.
+For a fixed artifact (such as $\mathtt{T003}$), $k$ and the base are pinned at generation time; the bounded universality theorem applies to that specific slice.
+6.2 From Bounded to Classical Universality
+Remark 6.2.1.
+From an effective realization, classical results follow.
+Lemma 6.2.2 (Existential bound extraction).
+Define the unbounded satisfiability predicate:
+$\mathrm{Sat}(t)\;:\iff\;\exists\,k\in\mathbb{N}\;\;\mathrm{Sat}_{B_{k}}(t),$ (155)
+where $B_{k}$ and the corresponding aggregated polynomial $U_{k}$ are those of Theorem 6.1.8 and Remark 6.1.9. Equivalently,
+$\mathrm{Sat}(t)\;\iff\;\exists\,k\in\mathbb{N}\;\exists\,\rho\;\big(\rho(i_{u})=t\;\wedge\;\mathrm{Bound}_{B_{k}}(\rho)\;\wedge\;U_{k}(\rho)=0\big).$ (156)
+Immediate. The two formulations are equivalent by unfolding $\mathrm{Sat}_{B_{k}}$. ∎
+Remark 6.2.3.
+The above is elementary but decisive.
+It isolates the quantifier shift: unbounded provability is existential over resource bounds plus bounded search.
+Each slice $\mathrm{Sat}_{B_{k}}$ involves a single cubic equation $U_{k}$ with fixed base; the existential over $k$ absorbs the dependence of the base on the proof size.
+Proposition 6.2.4 (Classical universality).
+Let $\operatorname{Prov}(t)$ denote ordinary (unbounded) provability in $\mathsf{RA}$. Under the correctness of the encoding (Theorem 1.3.1),
+$\operatorname{Prov}(t)\;\iff\;\mathrm{Sat}(t).$ (157)
+By correctness.
+Forward. Any proof has finite size, hence fits within some resource bound $k$. By Theorem 6.1.8, $\operatorname{Prov}_{k}(t)\implies\mathrm{Sat}_{B_{k}}(t)\implies\mathrm{Sat}(t)$.
+Reverse.
+$\mathrm{Sat}(t)$ means $\exists k\;\mathrm{Sat}_{B_{k}}(t)$.
+The reverse direction of bounded universality (Theorem 6.1.8) gives $\operatorname{Prov}_{k}(t)$, hence $\operatorname{Prov}(t)$.
+∎
+6.3 Decidability versus Semi-Decidability
+Proposition 6.3.1.
+Each slice $\mathrm{Sat}_{B_{k}}(t)$ is decidable (Proposition 6.1.7).
+Corollary 6.3.2.
+$\mathrm{Sat}(t)$ is recursively enumerable.
+By construction. Enumerate pairs $(k,\rho)$ with $\mathrm{Bound}_{B_{k}}(\rho)$ and test $U_{k}(\rho)=0$. ∎
+Proposition 6.3.3 (Computable bounding implies decidability).
+The critical structural observation is that the RE predicate $\mathrm{Sat}(t)$ cannot collapse to any single decidable slice.
+If there existed a computable function $f\colon\mathbb{N}\to\mathbb{N}$ such that
+$\mathrm{Sat}(t)\;\iff\;\mathrm{Sat}_{B_{f(t)}}(t),$ (158)
+then $\mathrm{Sat}(t)$ would be decidable.
+Immediate. Given $t$, compute $f(t)$, then decide $\mathrm{Sat}_{B_{f(t)}}(t)$ by Proposition 6.1.7. ∎
+Theorem 6.3.4 (Undecidability consequence).
+If $\mathsf{RA}$ is sufficiently strong and consistent, then no computable bounding function $f$ as in Proposition 6.3.3 exists.
+By reductio.
+Proposition 6.2.4, $\mathrm{Sat}(t)\iff\operatorname{Prov}(t)$.
+The set of theorems of a consistent, sufficiently strong theory is $\mathsf{RE}$ but not decidable (Church’s theorem Church [1936]; Gödel’s first incompleteness theorem Gödel [1931]).
+If $f$ existed, $\mathrm{Sat}(t)$ would be decidable by Proposition 6.3.3, contradicting undecidability of $\operatorname{Prov}(t)$.
+∎
+6.4 Conceptual Synthesis
+Remark 6.4.1.
+The results of this section are summarized as follows.
+  1. (i)
+    Effective construction produces concrete cubic artifacts: finite coefficient tables (Table 2) with kernel-checked degree certificates and digest gates.
+Each artifact is a computable object, not an existence assertion.
+  2. (ii)
+    Bounded universality yields decidable Diophantine characterizations: for each $k\in\mathbb{N}$, the predicate $\mathrm{Sat}_{B_{k}}$ is decidable by finite search (Proposition 6.1.7).
+Each bounded artifact is a constructive witness of $\mathsf{RE}$-ness for its slice.
+  3. (iii)
+    Classical universality is their $\mathsf{RE}$ union:
+    $\mathrm{Sat}\;=\;\bigcup_{k\in\mathbb{N}}\mathrm{Sat}_{B_{k}}.$ (159)
+    The bounded family indirectly recovers the classical result: any $\mathsf{RE}$ set that arises as such a union is classically recognizable, and the compiler witnesses that $\mathsf{RA}$-theoremhood is among them.
+  4. (iv)
+    Incompleteness is the non-existence of a computable collapse from the union to a single decidable slice.
+No computable selector $k=f(t)$ exists (Theorem 6.3.4).
+This is the incompleteness phenomenon made computationally visible through the Diophantine lens.
+Mechanization 6.4.2.
+The contribution is an explicit, effectively generated cubic compiler tied to $\mathsf{RA}$ with mechanized bounded encoding.
+The $\mathtt{T003}$ artifact publishes both parameterizations (minimal and effective) as machine-readable JSON coefficient tables alongside rendered HTML equation views.
+The Python generator $\mathtt{gen.py}$ is the source of truth for encoding semantics; Rocq consumes the generated table and $\mathtt{VarMap}$.
+The universality equivalence is proved in a separate semantic bridge layer (Section 5) without normalizing the coefficient table.
+[…]
+Figure 3: External generation and kernel-level validation of the cubic Diophantine artifact.
+The Python generator produces a coefficient table (chunked) together with an expected digest; Rocq checks digest agreement and the degree-$3$ certificate, then provides evaluation and the bounded satisfiability predicate.
+The semantic equivalence to bounded proof checking is established in a separate bridge layer (Section 5).
 7 Cubic Thresholds
-6.1 On the Universal Cubic
-Intuition.
-The preceding construction already furnishes a uniform primitive recursive compiler from arithmetic sentences to single cubic equations, suggesting that universality at degree $3$ is not obstructed.
-Achieving a fixed universal cubic would require internalizing this compiler into one static cubic template while preserving strict degree control via band isolation and guard activation.
-Conjecture 6.1 (Universal Cubic Conjecture).
-There must be a finite equation,
-$U(\vec{x},n)\in\mathbb{Z}[\vec{x},n]$ (6.125)
-of total degree $\leq 3$ such that for every recursively enumerable set $W\subseteq\mathbb{N}$ there exists a parameter $n_{W}\in\mathbb{N}$ with
-$k\in W\;\Longleftrightarrow\;\exists\vec{y}\in\mathbb{N}^{m}\;U(k,\vec{y},n_{W})=0.$ (6.126)
 7.1 Cubic Upper Bound
-Intuition.
+Remark 7.1.1.
 The result obtained below is a quick and dense worst-case closure bound.
-It is not intrinsic to degree $3$, and one could certainly lower it by introducing only the multiplicative wiring actually used by the chosen quartic universal polynomial, as shown in Corollary 6.4.
-Lemma 6.2 (Quartic-to-cubic closure).
+It is not intrinsic to degree $3$, and one could certainly lower it by introducing only the multiplicative wiring actually used by the chosen quartic universal polynomial, as shown in Corollary 7.1.4.
+Lemma 7.1.2 (Quartic-to-cubic).
 For every $v\in\mathbb{N}$ there exists a uniform many–one reduction from quartic to cubic Diophantine satisfiability given by
-$\mathrm{H}{\mathbb{N}}(4,v)\;\leq_{m}\;\mathrm{H}{\mathbb{N}}(3,\,v+\binom{v+1}{2}\,)$. (6.127)
+$\mathrm{H}{\mathbb{N}}(4,v)\;\leq_{m}\;\mathrm{H}{\mathbb{N}}(3,\,v+\binom{v+1}{2}\,)$. (160)
+By construction.
 Introduce auxiliary variables $p_{ij}(1\leq i\leq j\leq v)$ subject to $p_{ij}=x_{i}x_{j}.$ Each such constraint has degree $2$ and every quartic monomial $x_{a}x_{b}x_{c}x_{d}$ occurring in $P$ is rewritten as $x_{a}x_{b}x_{c}x_{d}\;\longmapsto\;p_{ab}\,p_{cd},$ applying a fixed pairing rule (e.g., $a\leq b$, $c\leq d$, $ab\leq cd$ lexicographically), let $\widetilde{P}(x,p)$ denote the resulting expression.
@@
 Solvability is preserved for $\exists x\in\mathbb{N}^{v}\;P(x)=0$ if and only if,
-$\exists(x,p)\in\mathbb{N}^{v+\binom{v+1}{2}}\;\Bigl(\widetilde{P}(x,p)=0\;\wedge\;\forall\,i\leq j:\;p_{ij}=x_{i}x_{j}\Bigr).$ (6.128)
-Corollary 6.3 (Trivial cubic threshold).
+$\exists(x,p)\in\mathbb{N}^{v+\binom{v+1}{2}}\;\Bigl(\widetilde{P}(x,p)=0\;\wedge\;\forall\,i\leq j:\;p_{ij}=x_{i}x_{j}\Bigr).$ (161)
+Corollary 7.1.3 (Trivial cubic threshold).
 By Jones [1980], we have a “quartic universal” as a starting point:
-$\mathrm{H}{\mathbb{N}}(4,58)={\rm U}.$ (6.129)
-Hence, by Lemma 1.5 and Corollary 5.3 it can be said,
-$\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}.$ (6.130)
-Follows from Lemma 6.2. The number of auxiliary variables is
-$\binom{59}{2}=\frac{59\cdot 58}{2}=1711.$ (6.131)
-The total variable count is therefore $58+1711=1769$. Since $\mathrm{H}{\mathbb{N}}(4,58)={\rm U}$, Lemma 6.2 yields $\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}$. ∎
-Corollary 6.4 (Sparsity).
+$\mathrm{H}{\mathbb{N}}(4,58)={\rm U}.$ (162)
+Hence, by Lemma 1.2.2 and Corollary 5.2.3,
+$\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}.$ (163)
+Immediate from Lemma 7.1.2. The number of auxiliary variables is
+$\binom{59}{2}=\frac{59\cdot 58}{2}=1711.$ (164)
+The total variable count is therefore $58+1711=1769$. Since $\mathrm{H}{\mathbb{N}}(4,58)={\rm U}$, Lemma 7.1.2 yields $\mathrm{H}{\mathbb{N}}(3,1769)={\rm U}$. ∎
+Corollary 7.1.4 (Sparsity).
 Let $P$ be a fixed quartic universal polynomial over $58$ variables. Define $m_{2}(P)$ as:
@@
 Then,
-$\mathrm{H}{\mathbb{N}}(3,58+m_{2}(P))={\rm U}.$ (6.132)
+$\mathrm{H}{\mathbb{N}}(3,58+m_{2}(P))={\rm U}.$ (165)
 In particular,
-$m_{2}(P)\leq\binom{59}{2}$, (6.133)
+$m_{2}(P)\leq\binom{59}{2}$, (166)
 so this always improves or matches the dense bound $1769$.
-Repeat the construction of Lemma 6.2, but introduce an auxiliary variable $p_{ij}$ only for pairs $(i,j)$ that actually appear in the chosen quartic-monomial decomposition of $P$.
+By construction.
+Repeat the construction of Lemma 7.1.2, but introduce an auxiliary variable $p_{ij}$ only for pairs $(i,j)$ that actually appear in the chosen quartic-monomial decomposition of $P$.
 No other product variables are used in the rewritten system, so the number of added variables is exactly $m_{2}(P)$.
@@
 7.2 Non-Enumerability of the Region
-Theorem 6.5 (Region Impredicativity).
+Theorem 7.2.1 (Region Impredicativity).
 Define
-$\mathcal{U}={(d,n)\mid\mathrm{H}{\mathbb{N}}(d,n)=\mathsf{U}}.$ (6.134)
+$\mathcal{U}=\{(d,n)\mid\mathrm{H}{\mathbb{N}}(d,n)=\mathsf{U}\}.$ (167)
 Fix a standard Gödel Numbering $e\mapsto\Phi_{e}$ of Turing Machines and a uniform coding of polynomial instances of type $(d,n)$. Let $\mathrm{Sol}_{d,n}(x)$ mean
 Instance $x$ has a solution in $\mathbb{N}$
-Then, (i) the Membership $(d,n)\in\mathcal{U}$ is a $\Pi^{0}_{2}$ predicate, (ii) for each $d$ there exists a minimal threshold
-$n^{*}(d)=\min{n\mid(d,n)\in\mathcal{U}}\in\mathbb{N}\cup{\infty},$ (6.135)
+Then, (i) the Membership $(d,n)\in\mathcal{U}$ is a $\Pi^{0}_{\geq 2}$ predicate, (ii) for each $d$ there exists a minimal threshold
+$n^{*}(d)=\min\{n\mid(d,n)\in\mathcal{U}\}\in\mathbb{N}\cup\{\infty\},$ (168)
 such that $n^{*}$ is not computable, (iii) hence $\mathcal{U}$ is not decidable.
-By Reductio ad absurdum in steps. We consult Kleene [1952],
+By reductio in steps. We consult Kleene [1952].
   1. Step 1.
     Fix Undecidability of $\mathrm{H}{\mathbb{N}}(d,n)$ as the statement where $\mathrm{Tot}_{d,n}(e)$ asserts totality of $\Phi_{e}$
-    $\forall e;\neg\Big(\mathrm{Tot}*{d,n}(e)\wedge\forall x,\big(\Phi_{e}(x)=1\leftrightarrow\mathrm{Sol}*{d,n}(x)\big)\Big),$ (6.136)
-    on instances of type $(d,n)$. Totality is $\Pi^{0}_{2}$, and correctness is arithmetic; hence the whole statement is $\Pi^{0}_{2}$ in $(d,n)$. This proves (i).
+    $\forall e\;\neg\Big(\mathrm{Tot}_{d,n}(e)\wedge\forall x\,\big(\Phi_{e}(x)=1\leftrightarrow\mathrm{Sol}_{d,n}(x)\big)\Big),$ (169)
+    on instances of type $(d,n)$. Totality is $\Pi^{0}_{\geq 2}$, and correctness is arithmetic; hence the whole statement is $\Pi^{0}_{\geq 2}$ in $(d,n)$. This proves (i).
   2. Step 2.
     For fixed $d$, if $\mathrm{H}{\mathbb{N}}(d,n)$ is undecidable, then so is $\mathrm{H}{\mathbb{N}}(d,n^{\prime})$ for all $n^{\prime}\geq n$.
-Hence each degree-slice of $\mathcal{U}$ is upward closed in (n).
-Therefore the minimal threshold $n^{*}(d)$ is well-defined in $\mathbb{N}\cup{\infty}$.
+Hence each degree-slice of $\mathcal{U}$ is upward closed in $n$.
+Therefore the minimal threshold $n^{*}(d)$ is well-defined in $\mathbb{N}\cup\{\infty\}$.
   3. Step 3.
@@
 Hence the entire undecidability region would admit a computable boundary, in that $d\longmapsto n^{*}(d)$ encodes the minimal variable count at which equations admit arbitrary $\Sigma^{0}_{1}$ predicates.
-Determining $n^{*}(d)$ would admit whether a fixed universal simulation can be compressed to $(d,n)$.
-Contradiction.
+Determining $n^{*}(d)$ would admit inconsistency.
 More concretely: from an index $e$, one could effectively construct a degree $d_{e}$ encoding of $\Phi_{e}$ whose minimal variable count reflects whether $\Phi_{e}$ is total.
@@
 This establishes (ii), (iii) follows immediately.
-Intuition.
+Remark 7.2.2.
 The present results formalize the intuition that the boundary of undecidability for Diophantine equations is inherently impredicative.
@@
 The guard gadget introduces the minimal impredicative coupling—asserting existence of a positive witness without explicit construction—and it is this coupling that both raises the degree from $2$ to $3$ and enables the encoding of self-referential constraints.
-This has a structural consequence.
-Any formal specification of the “first” degree $k$ at which undecidability arises, if carried out inside a theory strong enough to internalize the reduction of Theorem 1.8, would interact nontrivially with the same reflection mechanism.
-The act of precisely defining the threshold is, by the reflection route of Theorem 5.6 and Corollary 5.7, the act of crossing it.
-Whether $\mathcal{U}$ is itself recursively enumerable remains ope.
-What the present construction does establish is that the boundary resists specification from within any theory strong enough to formalize the reduction.
-The question,
+What the construction shows is that the boundary resists specification from within any theory strong enough to formalize the reduction. The question,
 At what point does the undecidability of undecidability begin?
@@
 One cannot “see” the cubic threshold without building the cubic encoding that establishes it.
-This motivates the following conjecture, understood less as a claim than as a structural bet, the act of precisely locating the undecidability boundary already require crossing it.
+This motivates the following conjecture, understood less as a claim than as a structural bet, the act of precisely locating the undecidability boundary already requires crossing it.
 We close with a bet that we cannot win, only lose by falsification.
-Conjecture 6.6 (Diophantine Gap Conjecture).
+Conjecture 7.2.3 (Diophantine Gap Conjecture).
 Let $\operatorname{T}$ be a consistent recursively axiomatized extension of weak arithmetic capable of representing primitive recursive functions and its own provability predicate.
-We suspect that $\operatorname{T}$ will not emit a $\rm D$} nor $\rm U$} for $\varphi$ in some arithmetized sentence expressing $\mathrm{H}{\mathbb{N}}(3,2)=\varphi$.
+We suspect that $\operatorname{T}$ will prove neither a $\rm D$} nor a $\rm U$} for $\varphi$ in some arithmetized sentence expressing $\mathrm{H}{\mathbb{N}}(3,2)=\varphi$.
 Acknowledgments